### v1

# L’approximation arithmétique de Fujita sur courbes adéliques

Abstract
In this paper, we will prove an analogue of Fujita’s approximation theorem under the framework of Arakelov theory over adelic curves, which proves a conjecture of Huayi Chen and Atsushi Moriwaki.
{altabstract}
Dans cet article, on démontrera une analogie du théorème de l’approximation de Fujita sous le cadre de la théorie d’Arakelov sur courbes adéliques, qui démontre une conjecture de Huayi Chen et Atsushi Moriwaki.
1 Introduction
Fujita approximation is an approximative version of Zariski decomposition of pseudo-effective divisors [25]. More precisely, let $X$ be a projective variety over a field $k$ and $L$ be a line bundle on $X$. The volume of $L$ is defined as
$\operatorname{vol}(L)=\limsup_{D\rightarrow\infty}\frac{\dim_{k}H^{0}(X,L^{\otimes D})}{D^{\dim(X)}/(\dim(X))!}.$
We say that $L$ is a big line bundle if $\operatorname{vol}(L)>0$. Fujita’s approximation theorem states that, for all $\epsilon>0$, there exist a $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(L^{\otimes p})$ can be decomposed the tensor product of an ample line bundle $A$ and an effective line bundle $M$, which satisfy
$p^{-\dim(X)}\operatorname{vol}(A)\geqslant\operatorname{vol}(L)-\epsilon.$
In fact, this result asserts that a big line bundle can be arbitrarily closely approximated by ample line bundles.
This approximated decomposition theorem is proved by T. Fujita [10] for the case of characteristic 00, and by S. Takagi [21] for an arbitrary characteristic. We refer the readers to [15, §11.4] for a survey on this topic.
1.1 Arithmetic analogue over number fields
It is quite nature to find an analogue of Fujita’s approximation theorem under the framework of Arakelov geometry, since we can define arithmetic positivities and arithmetic volume of a normed line bundle over arithmetic varieties.
1.1.1
The arithmetic analogue of volume function and arithmetic bigness under the framework of Arakelov geometry are introduced in [17, 18] by A. Moriwaki. Soon after these, voluminous works appeared on this topic, such as [19, 4, 24, 12, 20, 13].
Let $K$ be a number field, $\mathscr{X}$ be an arithmetic variety over $\operatorname{Spec}\mathcal{O}_{K}$ of dimension $d$, and $\overline{\mathscr{L}}$ be a Hermitian line bundle on $\mathscr{X}$. We define
$\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes D})=\log\#\{s\in H^{0}(\mathscr{X},\mathscr{L}^{\otimes D})\mid\|s\|_{\sup,\sigma}\leqslant 1,\;\forall\sigma\in M_{K,\infty}\},$
and
$\widehat{\operatorname{vol}}(\overline{\mathscr{L}})=\limsup_{n\to\infty}\frac{\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes n})}{n^{\dim(\mathscr{X})}/\dim(\mathscr{X})!}.$
By the works mentioned above, the arithmetic volume function $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ has many similar behaviors to those of the geometric setting $\operatorname{vol}(\raisebox{1.72218pt}{.})$. So it is expected that an arithmetically big line bundle $\overline{\mathscr{L}}$, which satisfies $\widehat{\operatorname{vol}}(\overline{\mathscr{L}})>0$, also has the Fujita approximation. In fact, this used to be a conjecture of A. Moriwaki.
1.1.2
In [2, 23], H. Chen and X. Yuan proved the arithmetic Fujita approximation over number field at almost the same time. Although they both applied [16, §3], but their methods are essential different. X. Yuan [23] takes pluri-subharmonic envelope to deal with the infinite part, which is well-known in complex analysis.
H. Chen applied his $\mathbb{R}$-filtration method developed in [3]. More precisely, by considering the successive minima on $E_{n}=H^{0}(\mathscr{X},\mathscr{L}^{\otimes n})$, we can make that $\bigoplus\limits_{n\geqslant 0}E_{n}$ is a filtrated graded linear series. Then by combining some properties of approximable graded linear series in [16], we apply some convergent properties of $\mathbb{R}$-filtrations to construct an ample line bundle which approximates $\overline{\mathscr{L}}$ well.
1.2 Arakelov geometry over adelic curves
In [7], H. Chen and A. Moriwaki developed a kind of Arakelov theory over so-called adelic curves. Based on this, they built the arithmetic intersection theory in [8], and studied the arithmetic positivities in [9]. Under this setting, we can study the arithmetic geometry over number fields, function fields and other bases with adelic structures by a united method.
1.2.1
Briefly speaking, an adelic curve $S=(K,(\mathcal{A},\Omega,\nu),\phi)$ is an arbitrary field $K$ which is equipped with a parameter set $\Omega$ of all its absolute values. We require the set $\Omega$ parameterized by $\phi$ has a $\sigma$-algebra structure $\mathcal{A}$ and a measure $\nu$.
Under this setting, we can generalize many classic results and topics on Arakelov geometry, including the arithmetic volume function, arithmetic positivity, and Hilbert-Samuel formula. Since the adelic structure becomes more complicated, the obstructions original from measurability, integrability, and non-discreteness make the developing process more technique. In fact, some classic properties are not valid any longer in this framework.
1.2.2
Inspired by [2, 23], it is quite natural to ask whether the Fujita approximation is still valid over adelic curves. Over a general adelic curve, we are lack of the Minkowskian property, so we cannot apply the filtrated method by successive minina. At the same time, some original constructions do not work any longer without the help of integral models.
In [9, §8.5], H. Chen and A. Moriwaki proved a similar result to the arithmetic Fujita approximation over adelic curves, which is called the relative Fujita approximation. More precisely, let an adelic line bundle $\overline{L}$ where $L$ is big on $X$. Then for any $t<\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$, there exists a $p\in\mathbb{N}$ sufficiently large and a birational morphism $\varphi:X^{\prime}\rightarrow X$, such that $(\varphi^{*}L)^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{N}$ is relatively ample and $\overline{M}$ is effective, and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})>pt$.
Compared with this relative version, the classic arithmetic Fujita approximation seems to take an "average" at every minima of minimal slope, and approximates all of them simultaneously. In fact, this is a conjecture proposed by Chen-Moriwaki in [9, Remark 9.2.7].
1.3 Main result
In this article, we prove the following result on the arithmetic Fujita approximation over adelic curves under the framework [7, 9]. More general than the conjecture proposed in [9, Remark 9.2.7], we prove the case of adelic $\mathbb{R}$-line bundle.
[…]
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, which satisfies either $\mathcal{A}$ is discrete or has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be an arithmetically big adelic $\mathbb{R}$-line bundle on $X$ over $S$. Then for any $\epsilon>0$, there exist a $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(\overline{L}^{\otimes p})=\overline{A}\otimes\overline{M}$, where $\overline{A}$ is arithmetically ample and $\overline{M}$ is arithmetically effective. Moreover, the inequality
$p^{-(\dim(X)+1)}\widehat{\operatorname{vol}}(\overline{A})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon$
is verified.
1.3.1
In [9], they develop a method which construct a graded linear series over the adelic curve of trivially valued field from the general setting by its Harder-Narasimhan filtration. In addition, we can modify the original norms by its spectral norm, which makes the graded linear series to be filtrated. Under this construction, the new graded linear series has some asymptotic invariants same as the original one. Then some ideas in [2, 16] can be absorbed in this work.
Once we prove the case of integral and $\mathbb{Q}$-line bundles, we can approximate an $\mathbb{R}$-line bundle by a $\mathbb{Q}$-line bundle in an effective direction, and then we can prove the case of $\mathbb{R}$-line bundle.
1.3.2
In the application of $\mathbb{R}$-filtration method, we need to take some limits of some graded linear series and measures, and most operations only work when the objects are measurable and integrable. If either $\mathcal{A}$ is discrete or has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$, there will be no obstructions on the measurability or integrability, and see [7, §4] for more details of the necessity. Of course, in [2, 23], there is no such problem since the above condition is satisfied automatically.
1.4 Organization of article
In §2, we introduce the notions of Arakelov theory over adelic curves and $\mathbb{R}$-filtration. In §3, we consider graded linear series over a general adelic curve and construct those over the adelic curve of trivially valued field filtrated by the Harder-Narasimhan filtration. In §4, we prove the property of Fujita approximation of certain graded linear series. In §5, we prove the arithmetic Fujita approximation of adelic line bundles.
Acknowledgements
I would like to thank Prof. Atsushi Moriwaki for introducing this subject to me. In addition, I would like to thank Prof. Huayi Chen for many useful discussions and suggestions, and Dr. Wenbin Luo for the advices on some key techniques. Parts of this work was accomplished during my visit to Institute for Theoretical Sciences, Westlake University, and I would like to thank their generous reception. Chunhui Liu was supported by National Science Foundation of China No. 12201153.
2 Arakelov theory over adelic curves
In this section, we will introduce the foundations of Arakelov over adelic curves and $\mathbb{R}$-filtration. The main references are [7, 9].
2.1 Foundations of Arakelov geometry over adelic curves
In this part, we recall some useful knowledge on the Arakelov geometry over adelic curves, see [7] for a systematical introduction to this subject.
2.1.1 Adelic curve
Let $K$ be a field. We denote by $M_{K}$ the set of all absolute values on $K$.
[…]
We say $(\Omega,\mathcal{A},\nu)$ an adelic structure on $K$ equipped with a mapping
$\begin{array}[]{rrcl}\phi:&\Omega&\longrightarrow&M_{K}\\ &\omega&\mapsto&|\raisebox{1.72218pt}{.}|_{\omega}\end{array}$
which satisfies the properties below:
  1. (1)
    $\mathcal{A}$ is a $\sigma$-algebra on $\Omega$, and $\nu$ is a measure on $(\Omega,\mathcal{A})$;
  2. (2)
    for all $a\in K^{\times}$, the function
    $\begin{array}[]{rcl}\Omega&\longrightarrow&\mathbb{R}\\ w&\mapsto&\log|a|_{w}\end{array}$
    is $\mathcal{A}$-measurable, integrable with respect to $\nu$.
In addition, the data $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ is called an adelic curve.
We denote by $\Omega_{f}$ the subset of $\Omega$ whose images in $M_{K}$ with respect to $\phi$ is non-archimedean, and by $\Omega_{\infty}$ the sbuset of $\Omega$ whose image in $M_{K}$ is archimedean. We have $\Omega=\Omega_{f}\sqcup\Omega_{\infty}$ by definition directly.
[…]
We say that an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ is proper if for all $a\in K^{\times}$, we always have
$\int_{S}\log|a|_{w}\nu(\mathrm{d}w)=0,$
and the above equality is called the product formula of $S$.
We refer the readers to [7, §3] for more details on this notion.
[…]
Let $K$ be a number field or a function field. Then the usual adelic ring $\mathbb{A}_{K}$ gives an adelic structure, and constructs examples of proper adelic curves.
Let $K=\mathbb{Q}(\varphi_{1},\ldots,\varphi_{d})$ be a finitely transcendental extension over $\mathbb{Q}$. In [17], A. Moriwaki constructed a special adelic structure over $K$, and defined a height function with respect to it. This height function has been applied to study many arithmetic geometry subjects.
[…]
Let $K$ be an arbitrary field, and |.$|\raisebox{1.72218pt}{.}|_{0}$ be the trivial absolute value over $K$, which means
$|a|_{0}=\begin{cases}1,\text{ if }a\neq 0,\\ 0,\text{ if }a=0.\end{cases}$
We note $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ the adelic curve which counts the absolute value |.$|\raisebox{1.72218pt}{.}|_{0}$ only once. This example will be applied later.
We refer the readers to [7, §3.2] for more examples of adelic curve.
2.1.2 Adelic vector bundle
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, and $E$ be a vector space over $K$. For a family of norms (.$(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ on $E_{w}=E\otimes_{K}K_{w}$ over $K_{w}$, if the norm family (.$(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ on $E$ and its dual norm family (.$(\|\raisebox{1.72218pt}{.}\|^{\vee}_{w})_{w\in\Omega}$ on $E^{\vee}$ are both $\mathcal{A}$-measurable on $\Omega$, then we say $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ is an adelic vector bundle over $S$. If $\dim(E)=1$, we say that $\overline{E}$ is an adelic line bundle over $S$.
We say that an adelic vector bundle $\overline{F}=(F,(\|\raisebox{1.72218pt}{.}\|^{\prime}_{w})_{w\in\Omega})$ is an adelic sub-bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $F$ is a $K$-vector subspace of $E$, and .$\|\raisebox{1.72218pt}{.}\|^{\prime}_{w}$ is the restriction of .$\|\raisebox{1.72218pt}{.}\|_{w}$ on $F$ for each $w\in\Omega$.
An adelic bundle $\overline{G}=(G,(\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w})_{w\in\Omega})$ is said to be an adelic quotient bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $G$ is a quotient vector space of $E$, and .$\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w}$ is the quotient norm .$\|\raisebox{1.72218pt}{.}\|_{w}$ with respect to the canonical homomorphism $E\rightarrow G$ for each $w\in\Omega$. If $G$ is the quotient space of $E$ modulo $F$, and equip with the quotient metrics from $\overline{F}$, then we denote $\overline{G}=\overline{E}/\overline{F}=\overline{E/F}$.
For an adelic vector bundle $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, we define $\det(\overline{E})=(\det(E),(\|\raisebox{1.72218pt}{.}\|_{\det,w})_{w\in\Omega})$ as follows: the vector $K$-space $\det(E)=\bigwedge^{\dim(E)}(E)$ is the determinant space of $E$, and .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$ is the determinant norm on $\det(E_{w})$ of .$\|\raisebox{1.72218pt}{.}\|_{w}$ for each $w\in\Omega$. We refer the readers to [7, §1.2.3] for the details of the construction of .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$.
[…]
In general, for an adelic vector bundle $\overline{E}$ over $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, we cannot make sure that all its subspace equipped with the subspace norms is still an adelic vector bundle, and its quotient spaces admit the same situation. However, if $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$, we will get adelic vector bundles from all the above constructions. See [7, Proposition 4.1.24] for more details.
2.1.3 Some Arakelov invariants
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, and $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ be an adelic vector bundle over $S$. For the determinant bundle $\det(\overline{E})=(\det(E),(\|\raisebox{1.72218pt}{.}\|_{\det,w})_{w\in\Omega})$ of $\overline{E}$, we select an $s\in\det(E)\smallsetminus\{0\}$, and define
$\widehat{\deg}(\overline{E})=-\int_{\Omega}\log\|s\|_{\det,w}\nu(\mathrm{d}w),$
which is well-defined due to [7, §4.2]. This is independent of the choice of $s$ by the product formula of $S$.
We define
$\widehat{\deg}_{+}(\overline{E})=\sup_{F\subseteq E}\widehat{\deg}(\overline{F}),$
where $\overline{F}$ runs over all the adelic sub-bundles of $\overline{E}$.
For a non-zero adelic bundle $\overline{E}$, we define the slope of $\overline{E}$ as
$\widehat{\mu}(\overline{E})=\frac{\widehat{\deg}(\overline{E})}{\dim(E)}.$
In addition, we define the maximal slope of $\overline{E}$ as
$\widehat{\mu}_{\max}(\overline{E})=\sup_{F\subseteq E}\widehat{\mu}(\overline{F}),$
where $\overline{F}$ runs over all adelic sub-bundles of $\overline{E}$. Similarly, we define the minimal slope of $\overline{E}$ as
$\widehat{\mu}_{\min}(\overline{E})=\inf_{E\twoheadrightarrow G}\widehat{\mu}(\overline{G}),$
where $\overline{G}$ runs over all adelic quotient bundles of $\overline{E}$.
We refer the readers to [7, §4.3] for more details and properties of these invariants.
2.2 Harder-Narasimhan filtration from $\mathbb{R}$-filtration
In this part, we introduce a method to filtrate an adelic vector bundle by its minimal slopes, which is called the Harder-Narasimhan filtration of an adelic vector bundle. See [3] for its origin and [7] for the version over adelic curve.
2.2.1 $\mathbb{R}$-filtrations on vector spaces
Let $K$ be a field, and $W$ be a $K$-vector space of finite dimension. For an $\mathbb{R}$-filtration on $W$, we mean a sequence $\mathcal{F}=\left(\mathcal{F}^{t}W\right)_{t\in\mathbb{R}}$ of vector subspaces of $W$, which satisfies the following conditions:
  1. (1)
    if $t\leqslant s$, then $\mathcal{F}^{s}W\subseteq\mathcal{F}^{t}W$;
  2. (2)
    $\mathcal{F}^{t}W=0$ for sufficiently large $t\in\mathbb{R}$, and $\mathcal{F}^{t}W=W$ for sufficiently small $t$;
  3. (3)
    the function
    $\begin{array}[]{rcl}\mathbb{R}&\longrightarrow&\mathbb{R}\\ t&\mapsto&\dim_{K}(\mathcal{F}^{t}W)\end{array}$
    is left continuous with respect to Euclidean topology.
2.2.2 A measure induced by $\mathbb{R}$-filtration
Let $K$ be a field. For a $K$-vector space $W$ of finite dimension equipped with an $\mathbb{R}$-filtration $\mathcal{F}$, we denote by $\nu_{(W,\mathcal{F})}$ (or by $\nu_{W}$ if there is no ambiguity) the Borel probability measure on $\mathbb{R}$, which is induced from the derivative of the function
$\begin{array}[]{rcl}\mathbb{R}&\longrightarrow&\mathbb{R}\\ t&\mapsto&-\frac{\dim(\mathcal{F}^{t}W)}{\dim(W)}.\end{array}$
In fact, $\nu_{W}$ is a linear combination of some Dirac measure. More precisely, if the $\mathbb{R}$-filtration $\mathcal{F}$ consists of a flag
$W=W_{0}\supsetneq W_{1}\supsetneq\cdots\supsetneq W_{n}=0$
together with the jumps at the real numbers $\lambda_{1}<\cdots<\lambda_{n}$, then we have
$\nu_{(W,\mathcal{F})}=\sum_{i=1}^{n}\frac{\dim\left(W_{i-1}/W_{i}\right)}{\dim(W)}\delta_{\lambda_{i}}.$
In addition, we define the function
$\begin{array}[]{rrcl}\lambda:&W&\longrightarrow&\mathbb{R}\cup\{+\infty\}\\ &x&\mapsto&\sup\{t\in\mathbb{R}\mid x\in\mathcal{F}^{t}W\}.\end{array}$ (1)
Then we have $\lambda(W)\subset\operatorname{supp}(\nu_{W})\cup\{+\infty\}$, and $\lambda(x)$ is finite if $x\neq 0$.
In addition, we define
$\lambda_{+}(W)=\int_{0}^{+\infty}x\nu_{W}(\mathrm{d}x)\text{, and }\lambda_{\max}(W)=\max\limits_{x\in W\smallsetminus\{0\}}\lambda(x).$ (2)
2.2.3 Harder-Narasimhan filtration
By the notion of $\mathbb{R}$-filtration, we introduce a filtration by minimal slopes of adelic bundles. This is original from an analogical study of vector bundles on a smooth projective curve [11].
[…]
Let $\overline{E}$ be a non-zero adelic vector bundle over a proper adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. For any $t\in\mathbb{R}$, let
$\mathcal{F}_{\operatorname{HN}}^{t}\overline{E}=\bigcap_{\epsilon>0}\sum_{\begin{subarray}{c}\{0\}\neq F\in\Theta(E)\\ \widehat{\mu}_{\min}(\overline{F})\geqslant t-\epsilon\end{subarray}}F=\sum_{\begin{subarray}{c}\{0\}\neq F\in\Theta(E)\\ \widehat{\mu}_{\min}(\overline{F})\geqslant t\end{subarray}}F,$
where $\Theta(E)$ denotes the set of all vector sub-spaces of $E$, and $\overline{F}$ is equipped with the sub-space norms. This $\mathbb{R}$-filtration $\mathcal{F}_{\operatorname{HN}}$ is called the Harder-Narasimhan filtration of $\overline{E}$.
The last equality in Definition 2.2.3 if from [7, Proposition 4.3.36]. In fact, the Harder-Narasimhan $\mathbb{R}$-filtration of $\overline{E}$ corresponds to an increasing flag of vector subspaces
$\{0\}=E_{1}\subsetneq E_{2}\subsetneq\cdots\subsetneq E_{n}=E$ (3)
and a decreasing sequence of real numbers $\mu_{1}>\cdots>\mu_{n}$ represents the jumps of the $\mathbb{R}$-filtration $(\mathcal{F}_{\operatorname{HN}}^{t}\overline{E})_{t\in\mathbb{R}}$, where
$\mu_{i}=\sup\left\{t\in\mathbb{R}\mid\dim_{K}(\mathcal{F}^{t}_{\operatorname{HN}}\overline{E})\geqslant i\right\}.$
In particular, still by [7, Proposition 4.3.36], we have $\mu_{n}=\widehat{\mu}_{\min}(\overline{E})$. The uniqueness of (3) is from [7, Theorem 4.3.58].
2.3 Filtrations on graded algebras
In this part, we study some asymptotic properties of graded linear algebras by $\mathbb{R}$-filtraion.
2.3.1 Approximable graded algebras
The notation of approximable graded algebra is from [2], whose idea is original from [16].
[…]
Let $K$ be a field, and $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ be a graded $K$-algebra. We define $E_{\bullet}$ is approximable if:
  1. (1)
    for each $D\in\mathbb{N}$, $E_{D}$ is of finite dimension, and $E_{D}\neq 0$ for sufficiently large $D\in\mathbb{N}$;
  2. (2)
    for every $\epsilon\in(0,1)$, there exists a $p_{0}\in\mathbb{N}^{+}$, such that for any $p\geqslant p_{0}$, we have
    $\liminf\limits_{n\to\infty}\frac{\dim\left(\operatorname{Im}\left(S^{n}E_{p}\rightarrow E_{np}\right)\right)}{\dim(E_{np})}>1-\epsilon,$
    where $S^{n}B_{p}\rightarrow E_{np}$ is the canonical homomorphism defined by the algebra structure on $E_{\bullet}$.
2.3.2 Convergence of measures on an approximable algebra
For a graded algebra, we equip an $\mathbb{R}$-filtration with all its homogeneous part. If it is an approximable graded algebras, we will have some properties of convergence.
Let $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ be an approximable graded algebra. Suppose for every $D\geqslant 0$, the vector space $E_{D}$ is equipped with an $\mathbb{R}$-filtration $\mathcal{F}$. We ignore the index $D\in\mathbb{N}$ since there will be no ambiguity.
[…]
With all the above notations and conditions. If for every homogeneous elements $x_{1},x_{2}\in E_{\bullet}$, we have
$\lambda(x_{1}x_{2})\geqslant\lambda(x_{1})+\lambda(x_{2}),$
where the function $\lambda(\raisebox{1.72218pt}{.})$ is defined at (1) and the homogeneous degrees are ignored, then we say $E_{\bullet}$ is filtrated with respect to $\mathcal{F}$.
For every $\epsilon>0$, let $T_{\epsilon}$ be the operator on the space of Borel probability measures on $\mathbb{R}$ by mapping $x\mapsto\epsilon x$.
We have the following result of filtrated graded algebras.
[…]
With all the above notations. Suppose that $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ is a filtrated graded algebra with respect to $\lambda(\raisebox{1.72218pt}{.})$ defined at (1), and
$\sup_{n\geqslant 1}\frac{\lambda_{\max}(E_{n})}{n}<+\infty,$
where $\lambda_{\max}(\raisebox{1.72218pt}{.})$ is defined at (2). Then the sequence $\left(\frac{\lambda_{\max}(E_{n})}{n}\right)_{n\in\mathbb{N}^{+}}$ converges in $\mathbb{R}$, and the sequence of measures $\left(T_{\frac{1}{n}}(\nu_{E_{n}})\right)_{n\geqslant 1}$ converges weakly to a Borel probability measure on $\mathbb{R}$.
2.3.3 Limits of a measure from filtrated algebras
We consider the filtrated approximable graded $K$-algebra $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$. Let $\nu_{E_{\bullet}}$ be the limit of the sequence of measures $\left(T_{\frac{1}{n}}(\nu_{E_{n}})\right)_{n\geqslant 1}$, and $\lambda^{\operatorname{asy}}_{\max}(E_{\bullet})=\lim\limits_{n\to\infty}\frac{\lambda_{\max}(E_{n})}{n}$, whose existences are assured by Proposition 2.3.2. In addition, we have
$\lambda_{+}^{\operatorname{asy}}(E_{\bullet}):=\int_{0}^{+\infty}x\nu_{E_{\bullet}}(\mathrm{d}x)=\lim\limits_{n\to\infty}\frac{\lambda_{+}(E_{n})}{n},$ (4)
where $\lambda_{+}(\raisebox{1.72218pt}{.})$ is defined at (2). See [2, §2.4] for more details.
We have the following result.
[…]
Let $E_{\bullet}$ be a filtrated approximable graded $K$-algebra. Then $\lambda_{+}^{\operatorname{asy}}(E_{\bullet})>0$ if and only if $\lambda_{\max}^{\operatorname{asy}}(E_{\bullet})>0$.
3 Graded linear series of global section spaces
Let $S$ be an adelic curve, $X$ be a projective scheme over $\operatorname{Spec}K$, and $\overline{L}$ be an adelic line bundle on $X$ over $S$. For our target, we will construct some graded linear series consisting of a particular kind of adelic vector bundles over $S$ from the above data.
3.1 Adelic line bundles on an arithmetic variety
In this part, we recall some notions about arithmetic positivity of the adelic line bundles on an arithmetic variety over adelic curve. We refer the readers to [9, §6 – §9] for a self-contained introduction to these notions.
3.1.1 Adelic line bundle
Let $S=(K,(\mathcal{A},\Omega,\nu),\phi)$ be an adelic curve, and $X$ be a projective scheme over $\operatorname{Spec}K$. We say that $\overline{L}=(L,\varphi)$ is an adelic line bundle on $X$, if $L$ is a line bundle on $X$, and $\varphi=(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ is a family of metrics on $L_{w}$, which is dominated and measurable.
In addition, we suppose that either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ contains a sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Then for every $w\in\Omega$ and $x\in X_{w}^{\operatorname{an}}$, we define
|.$\begin{array}[]{rrcl}|\raisebox{1.72218pt}{.}|_{\varphi_{w}}(x):&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&|s|_{\varphi_{w}}(x)\end{array}$
as the value of $s$ at the point $x$ with respect to the metric, and
.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\varphi_{w}}:&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&\sup\limits_{x\in X^{\operatorname{an}}_{w}}\{|s|_{\varphi_{w}}(x)\}.\end{array}$ (5)
We denote $\varphi=(\|\raisebox{1.72218pt}{.}\|_{\varphi_{w}})_{w\in\Omega}$ the norm family on $H^{0}(X,L)$ over $S$ if there is no ambiguity. By [7, Theorem 6.1.13, Theorem 6.1.32], $(H^{0}(X,L),\varphi)$ is an adelic vector bundle over $S$.
Let $\overline{L}=(L,\varphi)$ and $\overline{M}=(M,\psi)$ be two adelic line bundles on $X$. We denote by $\overline{L}\otimes\overline{M}=(L\otimes M,\varphi+\psi)$ the adelic line bundle, where $\varphi+\psi$ is the tensor product of metrics. In particular, for $n\in\mathbb{N}$, we denote by $\overline{L}^{\otimes n}=(L^{\otimes n},n\varphi)$ the adelic line bundle of $n$-times tensor product of $\overline{L}$, and by $(H^{0}(X,L^{\otimes n}),n\varphi)$ the induced adelic vector bundle over $S$. By [7, Proposition 6.1.28, Proposition 6.1.31], all the above constructions are well-defined.
We refer the readers to [7, §6.1] for a precise explanation of §3.1.1, see also [9, §2.6 – 2.8].
3.1.2 Semi-positive metric
Let $S=(K,(\mathcal{A},\Omega,\nu),\phi)$ be an adelic curve, $X$ be a projective scheme over $\operatorname{Spec}K$, $L$ be a line bundle on $X$, $\varphi=(\varphi_{w})_{w\in\Omega}$ and $\psi=(\psi_{w})_{w\in\Omega}$ be two families of metrics on $L$ such that $(L,\varphi)$ and $(L,\psi)$ are both adelic line bundles. Then we define the distance between $\varphi$ and $\psi$ as
$d(\varphi,\psi)=\int_{\Omega}\sup_{x\in X^{\operatorname{an}}_{w}}\left|\log\frac{|\raisebox{1.72218pt}{.}|_{\varphi_{w}}(x)}{|\raisebox{1.72218pt}{.}|_{\psi_{w}}(x)}\right|\nu(\mathrm{d}w).$
If $L$ is semiample, and if there exist a positive integer $m$ and a sequence $(\varphi_{n})_{n\in\mathbb{N}}$ of quotient metric families (where $\varphi_{n}$ a metric family of $L^{\otimes mn}$), such that
$\lim_{n\to\infty}\frac{1}{mn}d(nm\varphi,\varphi_{n})=0,$
we say that the metric family $\varphi$ is semi-positive.
3.1.3 Ampleness
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$. We say $\overline{L}$ is relatively ample if $L$ is ample and $\varphi$ is semi-positive.
In addition, an adelic line bundle $\overline{L}$ on $X$ is ample if it is relatively ample and if there exists an $\epsilon>0$ such that the inequality
$(\overline{L}|_{Y}^{\dim(Y)+1})\geqslant\epsilon\deg_{L}(Y)(\dim(Y)+1)$
holds for any integral closed subscheme $Y$ of $X$, where $(\overline{L}|_{Y}^{\dim(Y)+1})$ is the arithmetic intersection number studied in [8]. We also call it arithmetically ample to distinguish it from the geometric case.
[…]
Compared with the arithmetically ample definition of S. Zhang in [26], an arithmetically ample Hermitian line bundle on an arithmetic variety requires the metrics on the line bundle to be smooth, while the definition in [9, Definition 9.1.1] over an adelic curve only require continuous metrics. In this article, we follow the definition given in [9, Definition 9.1.1].
3.1.4 Nefness
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$. We say that $\overline{L}$ is relatively nef if there exists a relatively ample adelic line bundle $\overline{A}$ on $X$ and a positive integer $N$ such that, for any $n\geqslant N$, the tensor product $\overline{L}^{\otimes n}\otimes\overline{A}$ is relatively ample.
An adelic line bundle $\overline{L}$ is said to be nef or arithmetically nef if there exist an ample adelic line bundle $\overline{A}$ and a positive integer $M$ such that $\overline{L}^{\otimes m}\otimes\overline{A}$ is ample for any $m\geqslant M$ in the arithmetic sense.
3.1.5 Effectivity
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$. If $s\in H^{0}(X,L)$ is a non-zero global section such that $\|s\|_{\varphi_{w}}\leqslant 1$ for any $w\in\Omega$, we say that the global section $s$ is effective. We say that $\overline{L}$ is effective or arithmetically effective if it admits at least one effective global section by the above definition.
3.2 Algebraic construction of graded linear series
In this part we only consider the geometric setting, which means everything works over an arbitrary field $K$. For every $D\in\mathbb{N}^{+}$, we consider the global section space $H^{0}(X,L^{\otimes D})$, and the graded $K$-algebra $\bigoplus\limits_{n\geqslant 0}H^{0}(X,L^{\otimes n})$.
[…]
Let $X$ be a projective variety over $K$, and $L$ be a line bundle on $X$. For a graded linear series of $L$, we mean a graded sub-$K$-algebra of $\bigoplus\limits_{n\geqslant 0}H^{0}(X,L^{\otimes n})$.
3.2.1 Graded linear series containing an ample divisor
The following definition is original from [16, Definition 2.9], where it is called "Condition (C)". This notion plays a key role in the arithmetic Fujita approximation of [2].
[…]
Let $X$ be a projective variety, and $L$ be a line bundle on $X$. Suppose that there exists a $p\in\mathbb{N}^{+}$, an ample line bundle $A$, an effective line bundle $M$ on $X$, and a non-zero section $s\in H^{0}(X,M)$, such that $L^{\otimes p}=A\otimes M$. We say a graded linear series $W_{\bullet}=\bigoplus\limits_{n\geqslant 0}W_{n}$ of $L$ contains an ample divisor if the homomorphism of graded algebras
$\bigoplus_{n\geqslant 0}H^{0}(X,A^{\otimes n})\rightarrow\bigoplus_{n\geqslant 0}H^{0}(X,L^{\otimes np})$
induced by $s$ can factor through $\bigoplus\limits_{n\geqslant 0}W_{np}$.
In §2.3, we have seen that an approximable graded linear series has many convergent properties. The following property is original from [16], which is applied in [2] to judge whether a graded linear series is approximable.
[…]
If $W_{\bullet}=\bigoplus\limits_{n\geqslant 0}W_{n}$ contains an ample divisor and $W_{n}\neq 0$ for sufficiently large $n$, then $W_{\bullet}$ is approximable.
3.2.2 Big line bundles
Let $W_{\bullet}=\bigoplus\limits_{n\geqslant 0}W_{n}$ be a graded linear series of $L$. We denote
$\operatorname{vol}(W_{\bullet})=\limsup_{n\to\infty}\frac{\dim_{K}(W_{n})}{n^{\dim(X)}/\dim(X)!}.$
If $W_{n}=H^{0}(X,L^{\otimes n})$, we denote $\operatorname{vol}(L)=\operatorname{vol}(W)$.
If $\operatorname{vol}(L)>0$, we say that $L$ is a big line bundle on $X$.
3.3 Arithmetic invariants of graded linear series
In this part, we consider arithmetic constructions.
3.3.1
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over a proper adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ contains a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. In §3.1.1, we defined a norm family on $H^{0}(X,L^{\otimes D})$ by (5) indexed by $\Omega$, which is denoted by $D\varphi$ with no ambiguity. By [7, §6.4.1], .$\|\raisebox{1.72218pt}{.}\|_{D\varphi_{w}}$ is ultrametric for every $w\in\Omega_{f}$. Then $(H^{0}(X,L^{\otimes D}),D\varphi)$ is an adelic vector bundle over $S$.
Let $\overline{E}_{D}$ be an adelic vector sub-bundle $H^{0}(X,L^{\otimes D})$ over $S$. In this case, we say that $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ is a normed graded linear series over $S$ if $E_{D}\cdot E_{D^{\prime}}\subseteq E_{D+D^{\prime}}$ for $D,D^{\prime}\in\mathbb{N}$.
3.3.2
Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ be a normed graded linear series over $S$. We define the arithmetic volume of $\overline{E}_{\bullet}$ as
$\widehat{\operatorname{vol}}(\overline{E}_{\bullet})=\limsup_{n\to\infty}\frac{\widehat{\deg}_{+}\left(\overline{E}_{n}\right)}{n^{\dim(X)+1}/(\dim(X)+1)!}.$ (6)
In particular, if $\overline{E}_{D}=\left(H^{0}(X,L^{\otimes D}),D\varphi\right)$ as above, we denote $\widehat{\operatorname{vol}}(\overline{L})=\widehat{\operatorname{vol}}(\overline{E}_{\bullet})$ for simplicify.
If $\widehat{\operatorname{vol}}(\overline{L})>0$, we say that $\overline{L}$ is big on $X$, or we call it arithmetically big to distinguish it from the geometric case.
3.3.3
For every $D\in\mathbb{N}$, Let $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E}_{D})\right)_{t\in\mathbb{R}}$ be the $\mathbb{R}$-filtration induced by the Harder-Narasimhan filtration of $\overline{E}_{D}$ following Definition 2.2.3, where we ignore the homogeneous degree $D\in\mathbb{N}$.
We denote by $\nu_{\overline{E}_{D}}$ the Borel probability measure induced by the Harder-Narasimhan filtration of $\overline{E}_{D}$ constructed in the sense of §2.2.2. Then we have the result below from Proposition 4.3.50, Proposition 4.3.51 and Corollary 4.3.52 of [7], see also [9, §2.5.4].
[…]
With all the above notations and conditions. We have
$0\leqslant\widehat{\deg}_{+}(\overline{E}_{D})-\int_{0}^{+\infty}x\nu_{\overline{E}_{D}}(\mathrm{d}x)\leqslant\frac{1}{2}\nu(\Omega_{\infty})\dim_{K}(E_{D})\log(\dim_{K}(E_{D})).$
3.4 Graded linear series over trivially valued fields
Let $K$ be a field, $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a sub-field which is dense in $K_{w}$ for all $w\in\Omega$, and $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ be the adelic curve of trivially valued field introduced in Example 2.1.1.
In fact, for any normed vector space over $(K,|\raisebox{1.72218pt}{.}|_{0})$ of finite dimension, its norm takes only finitely many values (cf. [9, §3.1]).
In this part, we will study the adelic vector bundles and graded linear series over $S_{0}$ following that in [9, §3]. Some constructions are from those over $S$.
3.4.1
Let $\overline{V}$ be an adelic vector bundle over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, and $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{V})\right)_{t\in\mathbb{R}}$ be the Harder-Narasimhan filtration of $\overline{V}$ defined in Definition 2.2.3. We define the norm .$\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}$ on $V$ over $(K,|\raisebox{1.72218pt}{.}|_{0})$ as
.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}:&V&\longrightarrow&\mathbb{R}^{+}\\ &s&\mapsto&\exp\left(-\sup\{t\in\mathbb{R}\mid s\in\mathcal{F}^{t}_{\operatorname{HN}}(\overline{V})\}\right)\end{array}.$ (7)
Then $\overline{V}_{\operatorname{HN}}=\left(V,\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}\right)$ is an adelic vector bundle over $S_{0}$.
3.4.2
Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ be a normed graded linear series over $S=(K,(\Omega,\mathcal{A},\nu),\phi)$. For every $\overline{E}_{D}$, we denote by .$\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}}$ the norm (7) on $E_{D}$. Then $(E_{D},\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}})$ an Adélie vector bundle over $S_{0}$.
By [9, Proposition 4.1.3], for every $s\in E_{m}$ and $t\in E_{n}$, we have
$\|s\cdot t\|^{\operatorname{HN}}_{m+n}\leqslant\exp(f(m)+f(n))\|s\|^{\operatorname{HN}}_{m}\|t\|^{\operatorname{HN}}_{n},$
where $f(n)=\frac{3}{2}\nu(\Omega_{\infty})\log n$, and $\nu(\Omega_{\infty})$ is the measure of achimedean absolute values of $S=(K,(\Omega,\mathcal{A},\nu),\phi)$.
In order to obtain a new family of norms which are sub-multiplicative, we introduce the spectral norm .$\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ of .$\|\raisebox{1.72218pt}{.}\|^{\operatorname{HN}}_{D}$ as
.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}:&H^{0}(X,L^{\otimes D})&\longrightarrow&\mathbb{R}_{\geqslant 0}\\ &s&\mapsto&\lim\limits_{n\rightarrow\infty}\left(\|s^{n}\|_{D}^{\operatorname{HN}}\right)^{1/n}.\end{array}$ (8)
Then $\left(E_{D},\|\raisebox{1.72218pt}{.}\|_{D,\mathrm{sp}}\right)$ is an adelic vector bundle over $S_{0}$. By [9, Proposition 3.2.2 (3)], for every $s\in E_{m}$ and $t\in E_{n}$, we have
$\|s\cdot t\|_{\mathrm{sp},m+n}\leqslant\|s\|_{\mathrm{sp},m}\|t\|_{\mathrm{sp},n}.$
Let $\overline{E}_{\operatorname{HN},D}=\left(E_{D},\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}\right)$, then
$\overline{E}_{\operatorname{HN},\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{\operatorname{HN},D}$ (9)
is a filtrated graded linear series (see Definition 2.3.2 for the definition) over $S_{0}$, where we take $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$.
We denote by $\mathcal{F}_{\mathrm{sp}}$ the $\mathbb{R}$-filtration on $\overline{E}_{\operatorname{HN},D}$ induced by the filtrated structure $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$, where we ignore the index $D\in\mathbb{N}$.
3.4.3
Some asymptotic behaviors of $\overline{E}_{\operatorname{HN},\bullet}$ are useful for the study of $\overline{E}_{\bullet}$. Since $\overline{E}_{\operatorname{HN},\bullet}$ is filtrated, Proposition 2.3.2 and Proposition 2.3.3 will be applied in this process.
[…]
With all the notations in §3.4.2, we have
$\widehat{\operatorname{vol}}(\overline{E}_{\bullet})=\widehat{\operatorname{vol}}(\overline{E}_{\operatorname{HN},\bullet}).$
Proof.
It is from [9, Corollary 3.4.4, (4.1.2), Corollary-Definition 4.1.4] directly. ∎
By Proposition 2.3.2, the sequence $\{\frac{1}{n}\widehat{\mu}_{\max}(\overline{E}_{n})\}_{n\geqslant 0}$ converges to a real number, and we denote it by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\bullet})$. If $E_{n}=H^{0}(X,L^{\otimes n})$ equipped with the superior norms, it is denoted by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$ for simplicity.
We have the following result from [7, Proposition 6.4.18], which is an equivalent of arithmetically big adelic line bundle.
[…]
Let $\overline{L}$ be an adelic line bundle on $X$ over an adelic curve $S$. If $L$ is a big line bundle on $X$, then $\widehat{\operatorname{vol}}(\overline{L})>0$ if and only if $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})>0$.
From Proposition 3.3.3, we have the following result by an elementary calculation.
[…]
Let $\overline{E}^{\bullet}=\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$ be a normed graded linear series over $S$. If $\operatorname{vol}(E^{\bullet})>0$, then we have
$\lambda_{+}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN}}^{\bullet})=\frac{\widehat{\operatorname{vol}}(\overline{E}^{\bullet})}{(d+1)\operatorname{vol}(E^{\bullet})},$
where $\lambda_{+}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ is defined at (4) with respect to the canonical filtrated structure as (9) over $S_{0}$, and $d=\dim(X)$.
3.4.4
Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over an arbitrary curve $S$. We consider $\widehat{\mu}_{\min}(H^{0}(X,L^{\otimes n}),n\varphi)$, and we define
$\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})=\liminf_{n\to\infty}\frac{\widehat{\mu}_{\min}(H^{0}(X,L^{\otimes n}),n\varphi)}{n}.$
We recall some results on the criteria of arithmetic positivities via $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ in [9] below.
[…]
Let $\overline{L}$ be a relatively ample adelic line bundle. Then the following properties are equivalent:
  1. (1)
    $\overline{L}$ is ample;
  2. (2)
    $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})>0$;
  3. (3)
    there exists an $\epsilon>0$, such that for any closed sub-scheme $Y$ of $X$, we have $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L}|_{Y})>\epsilon$.
See §3.1.3 for the definitions of relatively ample and ample adelic line bundles.
[…]
Let $\overline{L}$ be an adelic line bundle. Then the following properties are equivalent:
  1. (1)
    $\overline{L}$ is nef;
  2. (2)
    $\overline{L}$ is relatively nef and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})\geqslant 0$.
See §3.1.4 for the definitions of relatively nef and nef adelic line bundles.
3.4.5
The following result shows that a power of an arithmetically big line bundle over $S_{0}$ can be decomposed as the sum of an ample and an effective line bundles, which is an analogue of [22, Corollary 2.4].
[…]
Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be arithmetically big on $X$ over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$. Then there exists a $p\in\mathbb{N}$, such that $\overline{L}^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{M}$ is arithmetically ample and $\overline{N}$ is arithmetically effective.
Proof.
By [14, Corollary 2.2.7], there exists a $p_{1}\in\mathbb{N}$, such that for all $t>p_{1}$, we are able to find an ample line bundle $M$ and an effective line bundle $M$ satisfying $L^{\otimes t}\cong A\otimes M$.
We equip $L$ with the metric $\varphi$ given in the statement such that $\overline{L}$ is arithmetically big, and equip $A$ with an arbitrary metric $\psi$ such that $\overline{M}$ is arithmetically ample. Then by [7, Corollary 6.4.10], there exists $p_{2}\in\mathbb{N}$, such that for all $t>p_{2}$, we have
$\widehat{\deg}_{+}\left(H^{0}(X,L^{\otimes t}\otimes M^{\vee}),t\varphi-\psi\right)>0.$
Since $S_{0}$ only has one place, then from the above inequality we deduce that $\overline{L}^{\otimes t}\otimes\overline{M}^{\vee}$ contains an effective section over $S_{0}$ in the sense of §3.1.5, so it is arithmetically effective over $S_{0}$ by the adelic structure of $S_{0}$.
We choose $p\geqslant\max\{p_{1},p_{2}\}$ which satisfying the requirement. ∎
[…]
The geometric version of Proposition 3.4.5 is firstly proved by K. Kodaira, see [14, Proposition 2.2.6, Corollary 2.2.7] for a brief introduction. An arithmetic version over number fields was proved by X. Yuan [22, Theorem 2.3, Corollary 2.4] via Siu’s inequality. It is an interesting question whether Proposition 3.4.5 is verified over a general adelic curve, or add some technical conditions. In fact, over a general adelic curve, we cannot deduce that an adelic line bundle $(L,\varphi)$ on $X$ contains an effective section from the inequality $\widehat{\deg}_{+}\left((H^{0}(X,L),\varphi)\right)>0$ directly.
4 Fujita approximation of graded linear series
In this section, we denote $E_{D}=H^{0}(X,L^{\otimes D})$ for simplicity, and all the constructions will be based on this notation.
4.1 Approximatible graded linear series
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable which is dense in $K_{w}$ for all $w\in\Omega$. Let $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ be the adelic curve of trivially valued field. In this part, we will consider some approximated properties of normed graded linear series over $S$ and $S_{0}$.
Let $\lambda\in\mathbb{R}_{\geqslant 0}$, $\overline{E}_{\operatorname{HN},0}^{\lambda}=K$, and
$\overline{E}_{\operatorname{HN},D}^{\lambda}=\operatorname{Vect}_{K}\left\{s\in H^{0}(X,L^{\otimes D})\mid\|s\|_{\mathrm{sp},D}\leqslant\exp(-D\lambda)\right\}$
for $D\geqslant 1$. So we have
$\overline{E}_{\operatorname{HN},D}^{\lambda}=\mathcal{F}_{\mathrm{sp}}^{D\lambda}\overline{E}_{\operatorname{HN},D},$
where the $\mathbb{R}$-filtration $\mathcal{F}_{\mathrm{sp}}$ is defined in §2.3.2, and $\overline{E}_{\operatorname{HN},D}$ follows the definition in (9) by taking $E_{D}=H^{0}(X,L^{\otimes D})$. In addition, we define
$\overline{E}^{\lambda}_{\operatorname{HN},\bullet}=\bigoplus_{D\geqslant 0}\overline{E}_{\operatorname{HN},D}^{\lambda},$ (10)
which is a graded linear series over $S_{0}$.
We will prove that (10) is approximable from the fact that it contains an ample divisor, where the strategy is similar to [2, Proposition 3.12].
4.1.1
We have the following proposition from the definition directly.
[…]
(10) is a filtrated graded linear series over $S_{0}$.
4.1.2
For $\lambda\in\mathbb{R}$, we define the adelic line bundle
$\overline{\mathcal{O}}_{\lambda}=(K,\|\raisebox{1.72218pt}{.}\|_{\lambda})$ (11)
over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$, where the norm satisfies $\|\mathbbm{1}\|_{\lambda}=e^{-\lambda}$.
[…]
Let $\pi:X\rightarrow\operatorname{Spec}K$ be a projective morphism, and $\overline{L}$ be an adelic line bundle over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$. With all the notations and assumption same as in (10). If $\overline{L}$ is arithmetically big, then for all $\lambda\in\left[0,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$, $\overline{E}^{\lambda,\bullet}_{\operatorname{HN}}$ contains an ample divisor.
Proof.
By the choice of $\lambda$, we have $\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L}\otimes\pi^{*}\overline{\mathcal{O}}_{-\lambda})>0$ by definition directly, so $\overline{L}\otimes\pi^{*}\overline{\mathcal{O}}_{-\lambda}$ is an arithmetically big line bundle on $X$ by Proposition 3.4.3. Then for a sufficiently large $D\in\mathbb{N}$, $L^{\otimes D}$ has a global section $s_{D}$ satisfying $\|s_{D}\|_{\mathrm{sp},D}\leqslant\exp(-D\lambda)$, which proves $\overline{E}_{\operatorname{HN},D}^{\lambda}\neq 0$ for $D\gg 0$.
Since $\overline{L}\otimes\pi^{*}\overline{\mathcal{O}}_{-\lambda}$ is arithmetically big on $X$, then by Propopsition 3.4.5, there exists $p\in\mathbb{N}$, such that $\left(\overline{L}\otimes\pi^{*}\overline{\mathcal{O}}_{-\lambda}\right)^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{M}$ is arithmetically ample and $\overline{N}$ is arithmetically effective, in other words $\overline{M}$ contains an effective section $s$.
The remainder part of this proof is quite similar to that of [2, Proposition 3.12]. We are able to take the above $p\in\mathbb{N}$ divisible enough such that the graded $K$-algebra $\bigoplus\limits_{n\in\mathbb{N}}H^{0}(X,M^{\otimes n})$ is generated by effective sections. These sections are considered as those of $\overline{M}\otimes\pi^{*}(\overline{\mathcal{O}}_{\lambda}^{\otimes p})$, whose norms are smaller than $\exp(-\lambda p)$. Then the homomorphism
$\bigoplus_{D\geqslant 0}H^{0}(X,M^{\otimes D})\longrightarrow\bigoplus_{D\geqslant 0}H^{0}(X,L^{\otimes pD})$
induced by $s$ factors through $\bigoplus\limits_{n\geqslant 0}\overline{E}_{\operatorname{HN},np}^{\lambda}$. ∎
Same as the argument in [2, Corollary 3.13], we have the following result.
[…]
The graded linear series (10) over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ is approximable for any $\lambda\in\left[0,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$.
Proof.
By Proposition 3.2.1, we have the assertion from Proposition 4.1.2 directly. ∎
4.2 Fujita approximation of graded linear series
Let $p\in\mathbb{N}$ such that $E_{\operatorname{HN},p}^{0}\neq 0$, and $\overline{E}_{\bullet}^{(p)}=\bigoplus\limits_{n\geqslant 0}\overline{E}^{(p)}_{\operatorname{HN},np}$ be the graded $K$-algebra generated by $E_{\operatorname{HN},p}^{0}$, which is considered as a normed graded linear series over $S_{0}$. At the same time, it is also able to be considered as a normed graded linear series over a general proper adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ equipped with the subspace norms of $\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$. In the both senses, they have the same arithmetic volume by Proposition 3.4.3 if $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
The following result is an analogue of [2, Theorem 4.1] over $S_{0}$, whose proof is quite similar to the ancient one.
[…]
Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be an arithmetically big line bundle on $X$ over a proper adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. With all the above notations and constructions, we have
$\widehat{\operatorname{vol}}(\overline{L})=\sup_{p\in\mathbb{N}}\widehat{\operatorname{vol}}(\overline{E}_{\bullet}^{(p)}),$
where $\overline{E}_{\bullet}^{(p)}$ is defined as above, which is a normed graded linear series over $S_{0}$.
Proof.
For any $n\in\mathbb{N}^{+}$, let $\nu_{\mathrm{sp},n}=T_{\frac{1}{n}}\nu_{\left(E_{\operatorname{HN},n},\mathcal{F}_{\mathrm{sp}}\right)}$, where the notation $T_{\frac{1}{n}}\nu_{\left(E_{\operatorname{HN},n},\mathcal{F}_{\mathrm{sp}}\right)}$ is introduced in §2.3.2, and the $\mathbb{R}$-filtration $\mathcal{F}_{\mathrm{sp}}$ is defined in §3.4.2 with respect to the spectral norms. Then by Proposition 2.3.2, the sequence $(\nu_{\mathrm{sp},n})_{n\geqslant 1}$ converges weakly to a Borel probability measure, which is denoted by $\nu_{\mathrm{sp}}$.
Similarly, let $\nu_{\mathrm{sp},n}^{(p)}=T_{\frac{1}{np}}\nu_{\left(E^{(p)}_{\operatorname{HN},np},\mathcal{F}_{\mathrm{sp}}\right)}$, which also converges weakly to a Borel probability measure noted by $\nu_{\mathrm{sp}}^{(p)}$ by Proposition 4.1.1.
In order to prove the assertion, we will prove that the restriction of $\nu_{\mathrm{sp}}$ on $[0,+\infty)$ is able to be approximated well by $\nu_{\mathrm{sp}}^{(p)}$.
Step I. – Simultaneous approximations.
Let
[…]
be a division of the interval $[0,\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L}))$ such that
$\nu_{\mathrm{sp}}(\{t_{1},\ldots,t_{m}\})=0.$ (12)
We define the function
$\begin{array}[]{rrcl}h_{D}:&\mathbb{R}&\longrightarrow&\mathbb{R}\\ &x&\mapsto&\sum\limits_{i=0}^{m-1}t_{i}\mathbbm{1}_{[t_{i},t_{i+1})}(x)+t_{m}\mathbbm{1}_{[t_{m},+\infty)}(x).\end{array}$
By Corollary 4.1.2 and Definition 2.3.1, for every $\epsilon>0$, there exists a $p\in\mathbb{N}$ sufficiently large, such that $E_{\bullet}^{(p)}$ approximates all $E^{t_{i}}_{\operatorname{HN},\bullet}$ defined at (10) simultaneously, where $i=0,\ldots,m$. In other words, there exists an $N_{0}\in\mathbb{N}$, such that for every $n\geqslant N_{0}$, we have
$\inf_{0\leqslant i\leqslant m}\frac{\dim_{K}\left(\operatorname{Im}\left(S^{n}E_{\operatorname{HN},p}^{t_{i}}\rightarrow E_{\operatorname{HN},np}^{t_{i}}\right)\right)}{\dim_{K}(E_{\operatorname{HN},np}^{t_{i}})}\geqslant 1-\epsilon$
for all $i=0,\ldots,m$. Then by the above inequality, we obtain
$\dim_{K}(\mathcal{F}_{\mathrm{sp}}^{npt_{i}}(E_{\operatorname{HN},np}^{(p)}))$$\geqslant$ $\dim_{K}\left(\operatorname{Im}\left(S^{n}E_{\operatorname{HN},p}^{t_{i}}\rightarrow E_{\operatorname{HN},np}^{t_{i}}\right)\right)$
$\geqslant$ $(1-\epsilon)\dim_{K}(E_{\operatorname{HN},np}^{t_{i}})$
for all $i=0,\ldots,m$.
Step II. – An upper bound of integrations.
We note that
$np\dim_{K}(E_{\operatorname{HN},np}^{(p)})\int_{0}^{+\infty}t\nu_{\mathrm{sp},np}^{(p)}(\mathrm{d}t)=-\int_{0}^{+\infty}t\mathrm{d}\dim_{K}(\mathcal{F}_{\mathrm{sp}}^{t}E_{\operatorname{HN},np}^{(p)})$
$=$ $\int_{0}^{+\infty}\dim_{K}(\mathcal{F}_{\mathrm{sp}}^{t}E_{\operatorname{HN},np}^{(p)})\mathrm{d}t\geqslant\sum_{i=1}^{m}np(t_{i}-t_{i-1})\dim_{K}(\mathcal{F}_{\mathrm{sp}}^{npt_{i}}E_{\operatorname{HN},np}^{(p)})$
$\geqslant$ $np(1-\epsilon)\sum_{i=1}^{m}(t_{i}-t_{i-1})\dim_{K}(E_{\operatorname{HN},np}^{t_{i}}),$
where the first inequality is from the decreasing of $\mathcal{F}_{\mathrm{sp}}$, and the last inequality is from (4.2) combined with the fact $E_{\operatorname{HN},np}^{t_{i}}=\mathcal{F}_{\mathrm{sp}}^{npt_{i}}E_{\operatorname{HN},np}$.
In addition, by applying Abel’s summation formula to the last term of the above inequality, we have
$\dim_{K}(E_{\operatorname{HN},np}^{(p)})\int_{0}^{+\infty}t\nu_{\mathrm{sp},np}^{(p)}(\mathrm{d}t)\geqslant(1-\epsilon)\dim_{K}(E_{\operatorname{HN},np})\int_{0}^{+\infty}h_{D}(t)\nu_{\mathrm{sp},np}(\mathrm{d}t).$
Step III. – Estimates of arithmetic volumes.
By Proposition 3.4.3, we have
$\lim_{n\to\infty}\frac{d+1}{(np)^{d+1}/(d+1)!}\dim_{K}(E_{\operatorname{HN},np}^{(p)})\int_{0}^{+\infty}t\nu_{\mathrm{sp},np}^{(p)}(\mathrm{d}t)$
$=$ $(d+1)\operatorname{vol}(E_{\bullet}^{(p)})\int_{0}^{+\infty}t\nu_{\mathrm{sp}}^{(p)}(\mathrm{d}t)=\widehat{\operatorname{vol}}(E_{\bullet}^{(p)}).$
Then we have
$\widehat{\operatorname{vol}}(E_{\bullet}^{(p)})$$\geqslant$ $\lim_{n\to\infty}\frac{d+1}{(np)^{d+1}/(d+1)!}(1-\epsilon)\dim_{K}(E_{\operatorname{HN},np})\int_{0}^{+\infty}h_{D}\mathrm{d}\nu_{\mathrm{sp},np}$
$=$ $(d+1)(1-\epsilon)\operatorname{vol}(L)\int_{\mathbb{R}}h_{D}\mathrm{d}\nu_{\mathrm{sp}},$
where the last equality is from [1, Chap. IV §5 $n^{\circ}$ 12 Proposition 22].
We choose a sequence of subdivisions $(D_{j})_{j\in\mathbb{N}}$ satisfying (12), such that the sequence of functions $(h_{D_{j}}(t))_{j\in\mathbb{N}}$ converges uniformly to
$\max\{t,0\}-\max\{t-\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L}),0\}$
when $j\to\infty$. By definition, the support of $\nu_{\mathrm{sp}}$ is bounded from above by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$, we obtain
$\sup_{p\in\mathbb{N}}\widehat{\operatorname{vol}}(E_{\bullet}^{(p)})\geqslant(d+1)(1-\epsilon)\operatorname{vol}(L)\int_{0}^{+\infty}t\nu_{\mathrm{sp}}(\mathrm{d}t)=(1-\epsilon)\widehat{\operatorname{vol}}(\overline{L}),$
where the last equality is from Proposition 3.4.3 on the equality of arithmetic volumes over $S$ and $S_{0}$. ∎
[…]
For a general graded linear series, if it satisfies the property of Proposition 4.2, we say that this graded linear series satisfies Fujita approximation. We would like to emphasize that this notation is different from those in [5, Definition 3.7] and [6, §6].
5 Fujita approximation of adelic line bundles
From this Fujita approximation of graded linear series proved in Proposition 4.2, we will prove the arithmetic Fujita approximation of an arithmetically big line bundle.
5.1 Fujita approximation of big adelic line bundles
In this part, we will prove the arithmetic Fujita approximation of usual adelic line bundles. First, we prove a nef version.
[…]
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big line bundle on $X$ over $S$. Then for every $\epsilon>0$, there exists a birational morphism $v:X^{\prime}\rightarrow X$ together with a positive integer $p$, such that the decomposition $v^{*}(\overline{L}^{\otimes p})=\overline{M}\otimes\overline{G}$ satisfies:
  1. (1)
    $\overline{M}$ is effective and $\overline{G}$ is arithmetically nef;
  2. (2)
    $p^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon$.
Proof.
We devide the proof into three parts.
Step I. – Determining the geometric decomposition.
Let $\pi:X\rightarrow\operatorname{Spec}K$ be the structural morphism. For any $p\geqslant 1$ such that $E_{\operatorname{HN},p}^{0}\neq 0$ (see Proposition 4.1.2 for the existence of such $p\in\mathbb{N}$), let
$\phi_{p}:X_{p}\rightarrow X$
be the blow-up along the base locus of $E_{\operatorname{HN},p}^{0}$, which means
$X_{p}=\operatorname{Proj}\left(\operatorname{Im}\left(\bigoplus_{n\geqslant 0}\pi^{*}(E_{\operatorname{HN},np}^{p})\longrightarrow\bigoplus_{n\geqslant 0}L^{\otimes np}\right)\right).$
Let $G_{p}=\mathcal{O}_{X_{p}}(1)$, and $M_{p}$ be the line bundle on $X_{p}$ with respect to the exceptional divisor. Let $s$ be a global section of $M_{p}$ which trivializes $M_{p}$ outside the exceptional divisor. Then by definition, we have
$\phi_{p}^{*}(L^{\otimes p})=G_{p}\otimes M_{p}.$
The canonical morphism
$\phi_{p}^{*}\pi^{*}(E_{\operatorname{HN},p}^{0})\rightarrow G_{p}$
induces a projective embedding $i_{p}:X_{p}\rightarrow\mathbb{P}\left(E_{\operatorname{HN},p}^{0}\right)$ with $i^{*}\left(\mathcal{O}_{E_{\operatorname{HN},p}^{0}}(1)\right)=G_{p}$. Then the restriction of global sections of $\mathcal{O}_{E_{\operatorname{HN},p}^{0}}(n)$ on $X_{p}$ gives an injective homomorphism
$\operatorname{Im}\left(S^{n}E_{\operatorname{HN},p}^{0}\rightarrow E_{\operatorname{HN},np}\right)=E_{\operatorname{HN},np}^{(p)}\rightarrow H^{0}(X_{p},G_{p}^{\otimes n}),$ (14)
where $H^{0}(X_{p},G_{p}^{\otimes n})$ is considered as a sub-$K$-vector space of $H^{0}(X_{p},\phi_{p}^{*}(L^{\otimes p}))$ via the section $s$.
Step II. – Constructing the metrics of line bundles.
For any $n\geqslant 1$ and $w\in\Omega$, let .$\|\raisebox{1.72218pt}{.}\|_{n,w}$ be the quotient norm on $G_{p,w}=G_{p}\times_{\operatorname{Spec}K}\operatorname{Spec}K_{w}$ induced by the surjective homomorphism
$\phi_{p}^{*}\pi^{*}(E_{\operatorname{HN},np}^{(p)})\rightarrow G_{p}^{\otimes n},$
where $E_{\operatorname{HN},np}^{(p)}$ is equipped with the superior norm .$\|\raisebox{1.72218pt}{.}\|_{\sup,w}$ with respect to $w\in\Omega$.
In fact, for any $x\in X_{p,w}(\overline{K}_{w})$ outside the exceptional divisor, it corresponds to a $1$-dimensional quotient of $E_{\operatorname{HN},p}^{0}$. For any $n\geqslant 1$, this quotient induces a $1$-dimensional quotient $\ell_{n,x}$ of $E_{\operatorname{HN},np}^{(p)}$. By Hahn-Banach theorem, every linear form on $\ker(E_{\operatorname{HN},np}^{(p)}\rightarrow\ell_{n,x})$ can be extended to a linear form on $E_{\operatorname{HN},np}^{(p)}$. In other words, there exists a section $w\in E_{\operatorname{HN},np}^{(p)}$ whose image in $G_{p,w}^{\otimes n}(x)$ has norm $\|w\|_{\sup,w}$. As a section of $L_{w}^{\otimes np}$ on $X_{v}(\overline{K}_{w})$, we have $\|w(x)\|_{w}\leqslant\|w\|_{\sup,w}$, where .$\|\raisebox{1.72218pt}{.}\|_{w}$ is the metric on $\overline{L}$ indexed by $v\in\Omega$. Then for any section $u\in G_{p,w}$ over a neighborhood of $x$, we may assume that $u(x)^{\otimes n}$ equals to the image of $w(x)$ in $G_{p,w}^{\otimes n}(x)$ by dilation, then we have
$\|u(x)\|_{n,w}=\|w\|_{\sup,w}^{1/n}\geqslant\|w(x)\|_{w}^{1/n}=\|u(x)\otimes s(x)\|_{w}.$
Let $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{W\in\Omega}$ be an adelic metric on $G_{p}$, and we define $(M_{p},\beta_{n})=\phi_{p}^{*}(\overline{L})\otimes(G_{p},\alpha_{n})^{\vee}$. In this sense, $s$ is an effective section in $(M_{p},\beta_{n})$, then it is an effective line bundle.
Step III. – The relation among arithmetic volumes.
By the construction in Step II, for any $n\geqslant 1$, we have
$p^{-(d+1)}\widehat{\operatorname{vol}}(G_{p},\alpha_{n})\geqslant\widehat{\operatorname{vol}}(E_{\bullet}^{(p)},\alpha_{n}),$
where
$\widehat{\operatorname{vol}}(E_{\bullet}^{(p)},\alpha_{n})=\lim_{m\rightarrow\infty}\frac{\widehat{\deg}_{+}(E_{\operatorname{HN},mp}^{(p)},(\|\raisebox{1.72218pt}{.}\|_{n,w,\sup})_{w\in\Omega})}{m^{d+1}/(d+1)!},$
and .$\|\raisebox{1.72218pt}{.}\|_{n,w,\sup}$ is the superior norm induced by $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ on $G_{p}$. Then by [2, Corollary A.2], we have
$\limsup_{n\to\infty}\widehat{\operatorname{vol}}(E_{\bullet}^{(p)},\alpha_{n})\geqslant\widehat{\operatorname{vol}}(\overline{E}_{\bullet}^{(p)}).$
By Proposition 4.2, for any $\epsilon>0$, there exists a $p\in\mathbb{N}$, such that
$\limsup_{n\to\infty}p^{-(d+1)}\widehat{\operatorname{vol}}(G_{p},\alpha_{n})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\frac{\epsilon}{2}.$
Then there exists an $N_{0}\in\mathbb{N}$, such that
$p^{-(d+1)}\widehat{\operatorname{vol}}(G_{p},\alpha_{n})-\frac{\epsilon}{2}\geqslant\widehat{\operatorname{vol}}(\overline{L})-\frac{\epsilon}{2}$
for infinitely many $n\geqslant N_{0}$. In addition, by Proposition 3.4.4 we obtain that $(G_{p},\alpha_{n})$ is arithmetically nef, since $\widehat{\mu}_{\min}(\overline{E}_{\operatorname{HN},np}^{(p)})\geqslant 0$ for all $n\geqslant 1$, then we deduce the positive asymptotic minimal slope.
To sum up, we obtain the assertion. ∎
Next, we obtain the following ample version from Theorem 5.1 directly.
[…]
We keep all the notations and conditions in Theorem 5.1. Then we are able to require $\overline{G}$ to be arithmetically ample in the term (1) of Theorem 5.1.
Proof.
Let $\overline{B}$ be an arithmetically big line bundle on $X$. Then by [7, Corollary 6.4.10], for $m\in\mathbb{N}$ sufficiently large, we have $\overline{L}^{\otimes m}\otimes\overline{B}^{\vee}$ is arithmetically big. By Theorem 5.1, for any $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$, a $p\in\mathbb{N}$, and a decomposition
$\phi^{*}(\overline{L}^{\otimes mp}\otimes\overline{B}^{\otimes(-p)})=\overline{G}\otimes\overline{M}$
with $\overline{G}$ arithmetically nef and $\overline{M}$ effective, such that
$p^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G})\geqslant\widehat{\operatorname{vol}}(\overline{L}^{\otimes m}\otimes\overline{B}^{\vee})-\epsilon.$
Let $\overline{B}^{\prime}$ be an arithmetically ample line bundle on $X^{\prime}$. Since $\phi^{*}(\overline{B})$ is arithmetically big, there exists a $q\in\mathbb{N}$, such that $\phi^{*}(\overline{B}^{\otimes q})\otimes\overline{B}^{\prime\vee}$ is effective. We consider
$\phi^{*}(\overline{L}^{\otimes mpq})=(\overline{G}^{\otimes q}\otimes\overline{B}^{\prime})\otimes(\overline{M}^{\otimes q}\otimes\phi^{*}(\overline{B}^{\otimes pq})\otimes\overline{B}^{\prime\vee}).$
In the above decomposition, $\overline{G}^{\prime}:=\overline{G}^{\otimes q}\otimes\overline{B}^{\prime}$ is arithmetically ample, $\overline{M}^{\otimes q}\otimes\phi^{*}(\overline{B}^{\otimes pq})\otimes\overline{B}^{\prime\vee}$ is effective, and
$(mpq)^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G}^{\prime})\geqslant(mpq)^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G}^{\otimes q})\geqslant m^{-(d+1)}(\widehat{\operatorname{vol}}(\overline{L}^{\otimes m}\otimes\overline{B}^{\vee})-\epsilon).$
By the continuity of the arithmetic volume function (see [7, Corollary 6.4.10]), we have the result. ∎
5.2 Case of $\mathbb{K}$-line bundles
In this part we will prove the arithmetic Fujita approximation is verified for adelic $\mathbb{Q}$-line bundles and $\mathbb{R}$-line bundles over adelic curve.
5.2.1
First, we prove the case of adelic $\mathbb{Q}$-line bundle, which is a quite straightforward from the usual case.
[…]
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$ with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big adelic $\mathbb{Q}$-line bundle on $X$ over $S$. For every $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$ together with a $p\in\mathbb{N}^{+}$, such that the decomposition $\phi^{*}(\overline{L}^{\otimes p})=\overline{M}\otimes\overline{G}$ satisfies:
  1. (1)
    $\overline{M}$ and $\overline{G}$ are adelic $\mathbb{Q}$-line bundles, where $\overline{M}$ is effective and $\overline{G}$ is arithmetically ample;
  2. (2)
    $p^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon$.
Proof.
By definition, there exists an $m\in\mathbb{N}$, such that $\overline{L}^{\otimes m}$ is a usual arithmetically big adelic line bundle. Then we obtain the assertion by applying Corollary 5.1 to $\overline{L}^{\otimes m}$. ∎
5.2.2
Next, we prove the case of adelic $\mathbb{R}$-line bundle, and see [7, §6.4.3] for an introduction to this notion. For this target, we prove an auxiliary lemma first, which allows us to approximate an arithmetically big $\mathbb{R}$-line bundle by an arithmetically big $\mathbb{Q}$-line bundle along an effective direction.
[…]
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, and $X$ be a projective variety over $\operatorname{Spec}K$. Let $\overline{L}$ be an arithmetically big $\mathbb{R}$-line bundle on $X$. Then for all $\epsilon>0$, there exists an arithmetically big $\mathbb{Q}$-line bundle $\overline{L}^{\prime}$, such that $\overline{L}-\overline{L}^{\prime}$ is effective and
$0<\widehat{\operatorname{vol}}(\overline{L})-\widehat{\operatorname{vol}}(\overline{L}^{\prime})<\epsilon.$
Proof.
The adelic $\mathbb{R}$-line bundle $\overline{L}$ is arithmetically big, so $L$ is a big $\mathbb{R}$-line bundle. By an equivalent definition [14, Definition 2.2.21], we have $L=\sum\limits_{i=1}^{r}a_{i}L_{i}$, where $a_{1},\ldots,a_{r}>0$ and $L_{1},\ldots,L_{r}$ are usual big line bundles. By Kodaira’s decomposition [14, Corollary 2.2.7], there exists an $m\in\mathbb{N}$, such that for all $i=1,\ldots,r$, we have $mL_{i}=A_{i}+M_{i}$, where $A_{i}$ is ample and $M_{i}$ is effective.
We equip metrics on $A_{1},\ldots,A_{r},M_{1},\ldots,M_{r}$ over $S$, such that $\overline{A}_{1},\ldots,\overline{A}_{r}$ are arithmetically ample, and $\overline{M}_{1},\ldots,\overline{M}_{r}$ are arithmetically effective. Then for any $b_{1},\ldots,b_{r},c_{1},\ldots,c_{r}\in\mathbb{R}^{+}$, we have $b_{1}\overline{A}_{1}+\cdots+b_{r}\overline{A}_{r}+c_{1}\overline{M}_{1}+\cdots+c_{r}\overline{M}_{r}$ is effective.
By the continuity of $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ from [7, Corollary 6.4.24], for all $\epsilon>0$, there exist $\delta_{1},\ldots,\delta_{r},\eta_{1},\ldots,\eta_{r}>0$, such that for all $t_{i}\in(0,\delta_{i})$, $s_{i}\in(0,\eta_{i})$ for $i=1,\ldots,r$, we always have
$0<\widehat{\operatorname{vol}}(m\overline{L})-\widehat{\operatorname{vol}}\left(m\overline{L}-\sum_{i=1}^{r}t_{i}\overline{A}_{i}-\sum_{i=1}^{r}s_{i}\overline{M}_{i}\right)<\epsilon.$
If there does not exist an arithmetically big $\mathbb{Q}$-line bundle satisfying the requirements of the lemma, then $m\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{A}_{i}-\sum\limits_{i=1}^{r}s_{i}\overline{M}_{i}$ can never be a $\mathbb{Q}$-line bundle for all $t_{i}\in(0,\delta_{i})$ and $s_{i}\in(0,\eta_{i})$ $i=1,\ldots,r$. It means there is no $\mathbb{Q}$-point in the region
$(ma_{1}-\delta_{1},ma_{1})\times\cdots\times(ma_{r}-\delta_{r},ma_{r})\times(ma_{1}-\eta_{1},ma_{1})\times\cdots\times(ma_{r}-\eta_{r},ma_{r})\subset\mathbb{R}^{2r},$
which has a strict positive Lebesgue measure. Then it contradicts to the fact that $\mathbb{Q}^{2r}$ is dense in $\mathbb{R}^{2r}$ with respect to the Euclidean topology.
By choosing the constant $\epsilon$ sufficiently small, we pick a family $t_{1},\ldots,t_{r},s_{1},\ldots,s_{r}\in\mathbb{R}^{+}$ such that $m\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{A}_{i}-\sum\limits_{i=1}^{r}s_{i}\overline{M}_{i}$ is an arithmetically big $\mathbb{Q}$-line bundle. By the above construction, $\sum\limits_{i=1}^{r}t_{i}\overline{A}_{i}+\sum\limits_{i=1}^{r}s_{i}\overline{M}_{i}$ is effective. Let
$\overline{L}^{\prime}=\overline{L}-\frac{1}{m}\sum\limits_{i=1}^{r}t_{i}\overline{A}_{i}-\frac{1}{m}\sum\limits_{i=1}^{r}s_{i}\overline{M}_{i},$
which satisfies the requirement. ∎
We prove the Fujita approximation of $\mathbb{R}$-line bundles via Lemma 5.2.2.
[…]
Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$. Let $X$ be a projective variety over $\operatorname{Spec}K$ with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big adelic $\mathbb{R}$-line bundle on $X$ over $S$. For every $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$ together with a $p\in\mathbb{N}^{+}$, such that the decomposition $\phi^{*}(\overline{L}^{\otimes p})=\overline{M}\otimes\overline{G}$ satisfies:
  1. (1)
    $\overline{M}$ and $\overline{G}$ are adelic $\mathbb{R}$-line bundles, where $\overline{M}$ is effective and $\overline{G}$ is arithmetically ample;
  2. (2)
    $p^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon$.
Proof.
By Lemma 5.2.2, for all $\epsilon>0$, there exists an arithmetically big $\mathbb{Q}$-line bundle $\overline{L}^{\prime}$, such that $\overline{L}-\overline{L}^{\prime}$ is effective and
$0<\widehat{\operatorname{vol}}(\overline{L})-\widehat{\operatorname{vol}}(\overline{L}^{\prime})<\epsilon.$
Then we apply Corollary 5.2.1 to $\overline{L}^{\prime}$, and we obtain the assertion. ∎

### v2

@@
 Abstract
-In this paper, we will prove an analogue of Fujita’s approximation theorem under the framework of Arakelov theory over adelic curves, which proves a conjecture of Huayi Chen and Atsushi Moriwaki.
+In this paper, we will prove an analogue of Fujita’s approximation under the framework of Arakelov theory over adelic curves, which gives a positive response to a conjecture of Huayi Chen and Atsushi Moriwaki.
 {altabstract}
-Dans cet article, on démontrera une analogie du théorème de l’approximation de Fujita sous le cadre de la théorie d’Arakelov sur courbes adéliques, qui démontre une conjecture de Huayi Chen et Atsushi Moriwaki.
+Dans cet article, on démontrera une analogie de l’approximation de Fujita sous le cadre de la théorie d’Arakelov sur courbes adéliques, qui donne une réponse positive d’une conjecture de Huayi Chen et Atsushi Moriwaki.
 1 Introduction
-Fujita approximation is an approximative version of Zariski decomposition of pseudo-effective divisors [25].
-More precisely, let $X$ be a projective variety over a field $k$ and $L$ be a line bundle on $X$.
+Fujita approximation is an approximative version of Zariski decomposition of pseudo-effective divisors, and see [28] for more details.
+More precisely, let $X$ be a projective variety over a field $k$, and $L$ be a line bundle on $X$.
 The volume of $L$ is defined as
-$\operatorname{vol}(L)=\limsup_{D\rightarrow\infty}\frac{\dim_{k}H^{0}(X,L^{\otimes D})}{D^{\dim(X)}/(\dim(X))!}.$
+$\operatorname{vol}(L)=\limsup_{D\rightarrow\infty}\frac{\dim_{k}H^{0}(X,L^{\otimes D})}{D^{\dim(X)}/\dim(X)!}.$
 We say that $L$ is a big line bundle if $\operatorname{vol}(L)>0$.
-Fujita’s approximation theorem states that, for all $\epsilon>0$, there exist a $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(L^{\otimes p})$ can be decomposed the tensor product of an ample line bundle $A$ and an effective line bundle $M$, which satisfy
+Fujita’s approximation theorem states that, for all $\epsilon>0$, there exist a $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(L^{\otimes p})$ can be decomposed as the tensor product of an ample line bundle $A$ and an effective line bundle $M$, which satisfy
 $p^{-\dim(X)}\operatorname{vol}(A)\geqslant\operatorname{vol}(L)-\epsilon.$
-In fact, this result asserts that a big line bundle can be arbitrarily closely approximated by ample line bundles.
+Intuitively, this result asserts that a big line bundle can be arbitrarily closely approximated by ample line bundles along an effective direction.
 This approximated decomposition theorem is proved by T.
@@
 Takagi [24] for an arbitrary characteristic.
-We refer the readers to [15, §11.4] for a survey on this topic.
+We refer the readers to [17, §11.4] for a survey on this topic.
 1.1 Arithmetic analogue over number fields
-It is quite nature to find an analogue of Fujita’s approximation theorem under the framework of Arakelov geometry, since we can define arithmetic positivities and arithmetic volume of a normed line bundle over arithmetic varieties.
+It is quite nature to ask whether there is an analogue of Fujita’s approximation theorem under the framework of Arakelov geometry, since we can define the arithmetic positivity and arithmetic volume of an adelic line bundle over arithmetic varieties.
 1.1.1
@@
 We define
-$\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes D})=\log\#\{s\in H^{0}(\mathscr{X},\mathscr{L}^{\otimes D})\mid\|s\|_{\sup,\sigma}\leqslant 1,\;\forall\sigma\in M_{K,\infty}\},$
+$\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes D})=\log\#\left\{s\in H^{0}(\mathscr{X},\mathscr{L}^{\otimes D})\mid\|s\|_{\sup,v}\leqslant 1,\;\forall\;v\in M_{K,\infty}\right\},$
 and
 $\widehat{\operatorname{vol}}(\overline{\mathscr{L}})=\limsup_{n\to\infty}\frac{\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes n})}{n^{\dim(\mathscr{X})}/\dim(\mathscr{X})!}.$
-By the works mentioned above, the arithmetic volume function $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ has many similar behaviors to those of the geometric setting $\operatorname{vol}(\raisebox{1.72218pt}{.})$.
-So it is expected that an arithmetically big line bundle $\overline{\mathscr{L}}$, which satisfies $\widehat{\operatorname{vol}}(\overline{\mathscr{L}})>0$, also has the Fujita approximation.
+By the works mentioned above, the arithmetic volume function $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ has many similar properties to those of the geometric setting $\operatorname{vol}(\raisebox{1.72218pt}{.})$, so it is expected that an arithmetically big line bundle $\overline{\mathscr{L}}$, which means $\widehat{\operatorname{vol}}(\overline{\mathscr{L}})>0$, also has the property of Fujita approximation.
 In fact, this used to be a conjecture of A.
@@
 Yuan proved the arithmetic Fujita approximation over number field at almost the same time.
-Although they both applied [16, §3], but their methods are essential different.
+Although they both applied [18, §3], but their methods are essential different.
 X.
-Yuan [23] takes pluri-subharmonic envelope to deal with the infinite part, which is well-known in complex analysis.
-H.
+Yuan [26, Theorem C] takes pluri-subharmonic envelope to deal with the infinite part, which is well-known in complex analysis.
+In [4], H.
 Chen applied his $\mathbb{R}$-filtration method developed in [5].
-More precisely, by considering the successive minima on $E_{n}=H^{0}(\mathscr{X},\mathscr{L}^{\otimes n})$, we can make that $\bigoplus\limits_{n\geqslant 0}E_{n}$ is a filtrated graded linear series.
-Then by combining some properties of approximable graded linear series in [16], we apply some convergent properties of $\mathbb{R}$-filtrations to construct an ample line bundle which approximates $\overline{\mathscr{L}}$ well.
+More precisely, by considering the successive minima on $H^{0}(\mathscr{X},\mathscr{L}^{\otimes n})$, we can make $\bigoplus\limits_{n\geqslant 0}H^{0}(\mathscr{X},\mathscr{L}^{\otimes n})$ to be a filtrated graded linear series.
+Then by combining some results of approximable graded linear series in [18], we are able to apply some convergent properties of $\mathbb{R}$-filtrations to construct an ample line bundle which approximates $\overline{\mathscr{L}}$ well.
+This is the main idea of the proof in [4].
 1.2 Arakelov geometry over adelic curves
@@
 Chen and A.
-Moriwaki developed a kind of Arakelov theory over so-called adelic curves.
-Based on this, they built the arithmetic intersection theory in [8], and studied the arithmetic positivities in [9].
-Under this setting, we can study the arithmetic geometry over number fields, function fields and other bases with adelic structures by a united method.
+Moriwaki developed a kind of Arakelov theory over the so-called adelic curve.
+Based on this framework, they built an arithmetic intersection theory in [10], and studied the arithmetic positivities in [11].
+Under this setting, we can study the arithmetic geometry over number fields, function fields and other bases with adelic structures under a same framework.
 1.2.1
-Briefly speaking, an adelic curve $S=(K,(\mathcal{A},\Omega,\nu),\phi)$ is an arbitrary field $K$ which is equipped with a parameter set $\Omega$ of all its absolute values.
-We require the set $\Omega$ parameterized by $\phi$ has a $\sigma$-algebra structure $\mathcal{A}$ and a measure $\nu$.
-Under this setting, we can generalize many classic results and topics on Arakelov geometry, including the arithmetic volume function, arithmetic positivity, and Hilbert-Samuel formula.
-Since the adelic structure becomes more complicated, the obstructions original from measurability, integrability, and non-discreteness make the developing process more technique.
-In fact, some classic properties are not valid any longer in this framework.
+Briefly speaking, an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ consists of an arbitrary field $K$ which is equipped with a parameter set $\Omega$ of all its absolute values.
+We require the set $\Omega$ parameterized by $\phi$ has a $\sigma$-algebra structure $\mathcal{A}$ and a measure $\nu$ compatible with $\mathcal{A}$.
+Under this setting, we can generalize many classic results and subjects of Arakelov geometry, including the arithmetic volume function, arithmetic positivity, and Hilbert-Samuel formula.
+Since the adelic structure becomes more complicated than those of classic cases, the obstructions original from measurability, integrability, and non-discreteness make the developing process more technique.
+In fact, some classic properties are not valid any longer in this framework, though we have developed many in [9, 10, 11], and see also [23].
 1.2.2
 Inspired by [4, 26], it is quite natural to ask whether the Fujita approximation is still valid over adelic curves.
-Over a general adelic curve, we are lack of the Minkowskian property, so we cannot apply the filtrated method by successive minina.
-At the same time, some original constructions do not work any longer without the help of integral models.
-In [9, §8.5], H.
+Over a general adelic curve, we are lack of the Minkowskian property, so we cannot apply the filtrated method by successive minina directly in [4] At the same time, some constructions do not work any longer without the help of integral models.
+In [11, §8.5], H.
 Chen and A.
 Moriwaki proved a similar result to the arithmetic Fujita approximation over adelic curves, which is called the relative Fujita approximation.
-More precisely, let an adelic line bundle $\overline{L}$ where $L$ is big on $X$.
-Then for any $t<\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$, there exists a $p\in\mathbb{N}$ sufficiently large and a birational morphism $\varphi:X^{\prime}\rightarrow X$, such that $(\varphi^{*}L)^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{N}$ is relatively ample and $\overline{M}$ is effective, and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})>pt$.
-Compared with this relative version, the classic arithmetic Fujita approximation seems to take an "average" at every minima of minimal slope, and approximates all of them simultaneously.
-In fact, this is a conjecture proposed by Chen-Moriwaki in [9, Remark 9.2.7].
+More precisely, let $\overline{L}$ be an adelic line bundle, where $L$ is big on $X$.
+Then for any $t<\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$, there exists a sufficiently large $p\in\mathbb{N}$ and a birational morphism $\varphi:X^{\prime}\rightarrow X$, such that $(\varphi^{*}\overline{L})^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{N}$ is relatively ample and $\overline{M}$ is effective, and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})>pt$.
+Here $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ denote the asymptotic maximal and minimal slopes respectively.
+Compared with this relative version above, the classic arithmetic Fujita approximation seems to take an "average" at every minima of minimal slope, and approximates all of them simultaneously, not only the maximal one.
+In fact, this is conjectured by Chen-Moriwaki in [11, Remark 9.2.7].
 1.3 Main result
-In this article, we prove the following result on the arithmetic Fujita approximation over adelic curves under the framework [7, 9].
-More general than the conjecture proposed in [9, Remark 9.2.7], we prove the case of adelic $\mathbb{R}$-line bundle.
+In this article, we prove the following result on the arithmetic Fujita approximation over adelic curves under the framework [9, 11], which gives a positive response to the conjecture [11, Remark 9.2.7].
+More generally, we prove a version of adelic $\mathbb{R}$-line bundle below at the same time, which is also a generalization of [4, Theorem 4.3] and [26, Theorem C].
 […]
-Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, which satisfies either $\mathcal{A}$ is discrete or has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
-Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be an arithmetically big adelic $\mathbb{R}$-line bundle on $X$ over $S$.
-Then for any $\epsilon>0$, there exist a $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(\overline{L}^{\otimes p})=\overline{A}\otimes\overline{M}$, where $\overline{A}$ is arithmetically ample and $\overline{M}$ is arithmetically effective.
-Moreover, the inequality
-$p^{-(\dim(X)+1)}\widehat{\operatorname{vol}}(\overline{A})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon$
-is verified.
+Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, which satisfies either $\mathcal{A}$ is discrete or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
+Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be an arithmetically big $\mathbb{R}$-line bundle on $X$ over $S$.
+Then for any $\epsilon>0$, there exist a sufficiently large $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(\overline{L}^{\otimes p})=\overline{A}\otimes\overline{M}$, where $\overline{A}$ is arithmetically ample and $\overline{M}$ is an arithmetically effective $\mathbb{R}$-line bundle.
+Moreover, we have
+$p^{-(\dim(X)+1)}\widehat{\operatorname{vol}}(\overline{A})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon.$
 1.3.1
-In [9], they develop a method which construct a graded linear series over the adelic curve of trivially valued field from the general setting by its Harder-Narasimhan filtration.
-In addition, we can modify the original norms by its spectral norm, which makes the graded linear series to be filtrated.
+In [11, §3], they develop a method which constructs a graded linear series over the adelic curve of trivially valued field from general one by its Harder-Narasimhan filtration.
+Next, we can modify the original norms by its spectral norm, which makes the graded linear series to be filtrated.
 Under this construction, the new graded linear series has some asymptotic invariants same as the original one.
 Then some ideas in [4, 18] can be absorbed in this work.
-Once we prove the case of integral and $\mathbb{Q}$-line bundles, we can approximate an $\mathbb{R}$-line bundle by a $\mathbb{Q}$-line bundle in an effective direction, and then we can prove the case of $\mathbb{R}$-line bundle.
+Once we prove the case of $\mathbb{Q}$-line bundle (in Corollary 5.2.1) from the case of usual adelic line bundle (in Corollary 5.1), we can approximate an arithmetically big $\mathbb{R}$-line bundle by an arithmetically big $\mathbb{Q}$-line bundle in an effective direction, and then we can prove the case of $\mathbb{R}$-line bundle (in Theorem 5.2.2).
 1.3.2
 In the application of $\mathbb{R}$-filtration method, we need to take some limits of some graded linear series and measures, and most operations only work when the objects are measurable and integrable.
-If either $\mathcal{A}$ is discrete or has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$, there will be no obstructions on the measurability or integrability, and see [7, §4] for more details of the necessity.
-Of course, in [2, 23], there is no such problem since the above condition is satisfied automatically.
+If either $\mathcal{A}$ is discrete or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$, there will be no obstructions on the measurability or integrability, and see [9, §4] for more details of the necessity.
+In order to study some arithmetic positivities by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$, we need to suppose the base field $K$ is perfect.
+For example, for an adelic line bundle $\overline{L}$ where $L$ is nef, we need to apply $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})\geqslant 0$ to deduce $\overline{L}$ is arithmetically nef by [11, Proposition 9.1.7], where we need the perfect condition.
+See [11, §9] for more details of the necessity.
+Of course, in [4, 26], there is no such problem since for the case of number field, because $\mathcal{A}$ is discrete in the language of adelic curve and the base field is perfect by its characteristic 00.
 1.4 Organization of article
-In §2, we introduce the notions of Arakelov theory over adelic curves and $\mathbb{R}$-filtration.
+In §2, we introduce the required notions of Arakelov theory over adelic curves and $\mathbb{R}$-filtration.
 In §3, we consider graded linear series over a general adelic curve and construct those over the adelic curve of trivially valued field filtrated by the Harder-Narasimhan filtration.
 In §4, we prove the property of Fujita approximation of certain graded linear series.
-In §5, we prove the arithmetic Fujita approximation of adelic line bundles.
+In §5, we prove the arithmetic Fujita approximation of adelic line bundles in the usual case, the case of $\mathbb{Q}$-line bundle and the case of $\mathbb{R}$-line bundle sequentially.
 Acknowledgements
@@
 Wenbin Luo for the advices on some key techniques.
-Parts of this work was accomplished during my visit to Institute for Theoretical Sciences, Westlake University, and I would like to thank their generous reception.
+Part of this work was accomplished during my visit to Institute for Theoretical Sciences, Westlake University, and I would like to thank their generous invitation and reception.
 Chunhui Liu was supported by National Science Foundation of China No.
@@
 2 Arakelov theory over adelic curves
-In this section, we will introduce the foundations of Arakelov over adelic curves and $\mathbb{R}$-filtration. The main references are [7, 9].
+In this section, we will introduce the foundations of Arakelov theory over adelic curves and of $\mathbb{R}$-filtration. The main references are [9, 11].
 2.1 Foundations of Arakelov geometry over adelic curves
-In this part, we recall some useful knowledge on the Arakelov geometry over adelic curves, see [7] for a systematical introduction to this subject.
+First, we recall some useful knowledge on the Arakelov geometry over adelic curves, and see [9] for a systematical introduction to this subject.
 2.1.1 Adelic curve
@@
 […]
-We say $(\Omega,\mathcal{A},\nu)$ an adelic structure on $K$ equipped with a mapping
+We say that $(\Omega,\mathcal{A},\nu)$ is an adelic structure on $K$ equipped with a mapping
 $\begin{array}[]{rrcl}\phi:&\Omega&\longrightarrow&M_{K}\\ &\omega&\mapsto&|\raisebox{1.72218pt}{.}|_{\omega}\end{array}$
@@
     $\begin{array}[]{rcl}\Omega&\longrightarrow&\mathbb{R}\\ w&\mapsto&\log|a|_{w}\end{array}$
-    is $\mathcal{A}$-measurable, integrable with respect to $\nu$.
+    is $\mathcal{A}$-measurable and integrable with respect to $\nu$.
 In addition, the data $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ is called an adelic curve.
-We denote by $\Omega_{f}$ the subset of $\Omega$ whose images in $M_{K}$ with respect to $\phi$ is non-archimedean, and by $\Omega_{\infty}$ the sbuset of $\Omega$ whose image in $M_{K}$ is archimedean.
+We denote by $\Omega_{f}$ the subset of $\Omega$ whose images in $M_{K}$ with respect to $\phi$ is non-archimedean, and by $\Omega_{\infty}$ the subset of $\Omega$ whose image in $M_{K}$ is archimedean.
 We have $\Omega=\Omega_{f}\sqcup\Omega_{\infty}$ by definition directly.
@@
 We say that an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ is proper if for all $a\in K^{\times}$, we always have
-$\int_{S}\log|a|_{w}\nu(\mathrm{d}w)=0,$
+$\int_{\Omega}\log|a|_{w}\nu(\mathrm{d}w)=0,$
 and the above equality is called the product formula of $S$.
-We refer the readers to [7, §3] for more details on this notion.
+We refer the readers to [9, §3] for more details on these notions.
 […]
-Let $K$ be a number field or a function field. Then the usual adelic ring $\mathbb{A}_{K}$ gives an adelic structure, and constructs examples of proper adelic curves.
-Let $K=\mathbb{Q}(\varphi_{1},\ldots,\varphi_{d})$ be a finitely transcendental extension over $\mathbb{Q}$.
+Let $K$ be a number field or a function field. Then the usual adèle ring of $K$ gives an adelic structure, and constructs examples of proper adelic curves.
+Let $K$ be a finitely transcendental extension over $\mathbb{Q}$, which is a sub-field of $\mathbb{C}$.
 In [19], A.
@@
 $|a|_{0}=\begin{cases}1,\text{ if }a\neq 0,\\ 0,\text{ if }a=0.\end{cases}$
-We note $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ the adelic curve which counts the absolute value |.$|\raisebox{1.72218pt}{.}|_{0}$ only once. This example will be applied later.
-We refer the readers to [7, §3.2] for more examples of adelic curve.
+We note $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ the adelic curve which counts the absolute value |.$|\raisebox{1.72218pt}{.}|_{0}$ once. By definition, $S_{0}$ is a proper adelic curve.
+Example 2.1.1 will be used in the article later. We refer the readers to [9, §3.2] for more examples of adelic curve.
 2.1.2 Adelic vector bundle
@@
 We say that an adelic vector bundle $\overline{F}=(F,(\|\raisebox{1.72218pt}{.}\|^{\prime}_{w})_{w\in\Omega})$ is an adelic sub-bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $F$ is a $K$-vector subspace of $E$, and .$\|\raisebox{1.72218pt}{.}\|^{\prime}_{w}$ is the restriction of .$\|\raisebox{1.72218pt}{.}\|_{w}$ on $F$ for each $w\in\Omega$.
-An adelic bundle $\overline{G}=(G,(\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w})_{w\in\Omega})$ is said to be an adelic quotient bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $G$ is a quotient vector space of $E$, and .$\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w}$ is the quotient norm .$\|\raisebox{1.72218pt}{.}\|_{w}$ with respect to the canonical homomorphism $E\rightarrow G$ for each $w\in\Omega$.
-If $G$ is the quotient space of $E$ modulo $F$, and equip with the quotient metrics from $\overline{F}$, then we denote $\overline{G}=\overline{E}/\overline{F}=\overline{E/F}$.
-For an adelic vector bundle $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, we define $\det(\overline{E})=(\det(E),(\|\raisebox{1.72218pt}{.}\|_{\det,w})_{w\in\Omega})$ as follows: the vector $K$-space $\det(E)=\bigwedge^{\dim(E)}(E)$ is the determinant space of $E$, and .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$ is the determinant norm on $\det(E_{w})$ of .$\|\raisebox{1.72218pt}{.}\|_{w}$ for each $w\in\Omega$.
-We refer the readers to [7, §1.2.3] for the details of the construction of .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$.
+An adelic bundle $\overline{G}=(G,(\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w})_{w\in\Omega})$ is said to be an adelic quotient bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $G$ is a quotient vector space of $E$, and .$\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w}$ is the quotient norm .$\|\raisebox{1.72218pt}{.}\|_{w}$ with respect to the canonical quotient homomorphism $E\rightarrow G$ for each $w\in\Omega$.
+If $G$ is the quotient space of $E$ modulo $F$, and is equipped with the quotient metrics from $\overline{F}$, then we denote $\overline{G}=\overline{E}/\overline{F}=\overline{E/F}$.
+For an adelic vector bundle $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, we define $\det(\overline{E})=(\det(E),(\|\raisebox{1.72218pt}{.}\|_{\det,w})_{w\in\Omega})$ as follows: the $K$-vector space $\det(E)=\bigwedge^{\dim(E)}(E)$ is the determinant space of $E$, and .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$ is the determinant norm on $\det(E)_{w}$ of .$\|\raisebox{1.72218pt}{.}\|_{w}$ for each $w\in\Omega$.
+We refer the readers to [9, §1.2.3] for the details of the construction of .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$.
 […]
 In general, for an adelic vector bundle $\overline{E}$ over $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, we cannot make sure that all its subspace equipped with the subspace norms is still an adelic vector bundle, and its quotient spaces admit the same situation.
-However, if $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$, we will get adelic vector bundles from all the above constructions.
-See [7, Proposition 4.1.24] for more details.
+However, if $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$, we will get adelic vector bundles from all the above constructions, and see [9, Proposition 4.1.24] for more details.
+In the whole article, we will always assume these conditions.
 2.1.3 Some Arakelov invariants
@@
 $\widehat{\deg}(\overline{E})=-\int_{\Omega}\log\|s\|_{\det,w}\nu(\mathrm{d}w),$
-which is well-defined due to [7, §4.2]. This is independent of the choice of $s$ by the product formula of $S$.
-We define
+which is well-defined due to [9, §4.2]. This is independent of the choice of $s$ by the product formula of $S$.
+We define the positive Arakelov degree of $\overline{E}$ as
 $\widehat{\deg}_{+}(\overline{E})=\sup_{F\subseteq E}\widehat{\deg}(\overline{F}),$
@@
 where $\overline{G}$ runs over all adelic quotient bundles of $\overline{E}$.
-We refer the readers to [7, §4.3] for more details and properties of these invariants.
+We refer the readers to [9, §4.3] for more details and properties of these invariants.
 2.2 Harder-Narasimhan filtration from $\mathbb{R}$-filtration
 In this part, we introduce a method to filtrate an adelic vector bundle by its minimal slopes, which is called the Harder-Narasimhan filtration of an adelic vector bundle.
+We will use the language of $\mathbb{R}$-filtration to formulate this construction.
 See [5] for its origin and [9] for the version over adelic curve.
@@
   2. (2)
-    $\mathcal{F}^{t}W=0$ for sufficiently large $t\in\mathbb{R}$, and $\mathcal{F}^{t}W=W$ for sufficiently small $t$;
+    $\mathcal{F}^{t}W=0$ for sufficiently large $t\in\mathbb{R}$, and $\mathcal{F}^{t}W=W$ for sufficiently small $t\in\mathbb{R}$;
   3. (3)
@@
 $\begin{array}[]{rcl}\mathbb{R}&\longrightarrow&\mathbb{R}\\ t&\mapsto&-\frac{\dim(\mathcal{F}^{t}W)}{\dim(W)}.\end{array}$
-In fact, $\nu_{W}$ is a linear combination of some Dirac measure. More precisely, if the $\mathbb{R}$-filtration $\mathcal{F}$ consists of a flag
+In fact, $\nu_{(W,\mathcal{F})}$ is a linear combination of some Dirac measures. More precisely, if the $\mathbb{R}$-filtration $\mathcal{F}$ consists of a flag
 $W=W_{0}\supsetneq W_{1}\supsetneq\cdots\supsetneq W_{n}=0$
 together with the jumps at the real numbers $\lambda_{1}<\cdots<\lambda_{n}$, then we have
-$\nu_{(W,\mathcal{F})}=\sum_{i=1}^{n}\frac{\dim\left(W_{i-1}/W_{i}\right)}{\dim(W)}\delta_{\lambda_{i}}.$
-In addition, we define the function
+$\nu_{(W,\mathcal{F})}=\sum_{i=1}^{n}\frac{\dim\left(W_{i-1}/W_{i}\right)}{\dim(W)}\delta_{\lambda_{i}},$
+where $\delta_{\lambda_{i}}$ is the Dirac operator at the real number $\lambda_{i}$.
+Next, we define the function
 $\begin{array}[]{rrcl}\lambda:&W&\longrightarrow&\mathbb{R}\cup\{+\infty\}\\ &x&\mapsto&\sup\{t\in\mathbb{R}\mid x\in\mathcal{F}^{t}W\}.\end{array}$ (1)
@@
 2.2.3 Harder-Narasimhan filtration
-By the notion of $\mathbb{R}$-filtration, we introduce a filtration by minimal slopes of adelic bundles. This is original from an analogical study of vector bundles on a smooth projective curve [11].
+By the notion of $\mathbb{R}$-filtration, we introduce a filtration by the minimal slopes of adelic bundles.
+This is original from an analogical study of vector bundles on a smooth projective curve [13].
 […]
@@
 $\mathcal{F}_{\operatorname{HN}}^{t}\overline{E}=\bigcap_{\epsilon>0}\sum_{\begin{subarray}{c}\{0\}\neq F\in\Theta(E)\\ \widehat{\mu}_{\min}(\overline{F})\geqslant t-\epsilon\end{subarray}}F=\sum_{\begin{subarray}{c}\{0\}\neq F\in\Theta(E)\\ \widehat{\mu}_{\min}(\overline{F})\geqslant t\end{subarray}}F,$
-where $\Theta(E)$ denotes the set of all vector sub-spaces of $E$, and $\overline{F}$ is equipped with the sub-space norms.
+where $\Theta(E)$ denotes the set of all vector subspaces of $E$, and $\overline{F}$ is equipped with the subspace norms.
 This $\mathbb{R}$-filtration $\mathcal{F}_{\operatorname{HN}}$ is called the Harder-Narasimhan filtration of $\overline{E}$.
-The last equality in Definition 2.2.3 if from [7, Proposition 4.3.36]. In fact, the Harder-Narasimhan $\mathbb{R}$-filtration of $\overline{E}$ corresponds to an increasing flag of vector subspaces
+The last equality in Definition 2.2.3 if from [9, Proposition 4.3.36]. In fact, the Harder-Narasimhan filtration of $\overline{E}$ corresponds to an increasing flag of vector subspaces
 $\{0\}=E_{1}\subsetneq E_{2}\subsetneq\cdots\subsetneq E_{n}=E$ (3)
@@
 $\mu_{i}=\sup\left\{t\in\mathbb{R}\mid\dim_{K}(\mathcal{F}^{t}_{\operatorname{HN}}\overline{E})\geqslant i\right\}.$
-In particular, still by [7, Proposition 4.3.36], we have $\mu_{n}=\widehat{\mu}_{\min}(\overline{E})$. The uniqueness of (3) is from [7, Theorem 4.3.58].
+In particular, still by [9, Proposition 4.3.36], we have $\mu_{n}=\widehat{\mu}_{\min}(\overline{E})$. The uniqueness of the flag (3) is from [9, Theorem 4.3.58].
 2.3 Filtrations on graded algebras
@@
 […]
-Let $K$ be a field, and $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ be a graded $K$-algebra. We define $E_{\bullet}$ is approximable if:
+Let $K$ be a field, and $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ be a graded $K$-algebra. We define that $E_{\bullet}$ is approximable if:
   1. (1)
@@
     for every $\epsilon\in(0,1)$, there exists a $p_{0}\in\mathbb{N}^{+}$, such that for any $p\geqslant p_{0}$, we have
-    $\liminf\limits_{n\to\infty}\frac{\dim\left(\operatorname{Im}\left(S^{n}E_{p}\rightarrow E_{np}\right)\right)}{\dim(E_{np})}>1-\epsilon,$
+    $\liminf\limits_{n\to\infty}\frac{\dim_{K}\left(\operatorname{Im}\left(S^{n}E_{p}\rightarrow E_{np}\right)\right)}{\dim_{K}(E_{np})}>1-\epsilon,$
     where $S^{n}B_{p}\rightarrow E_{np}$ is the canonical homomorphism defined by the algebra structure on $E_{\bullet}$.
 2.3.2 Convergence of measures on an approximable algebra
-For a graded algebra, we equip an $\mathbb{R}$-filtration with all its homogeneous part. If it is an approximable graded algebras, we will have some properties of convergence.
+For a graded algebra, we equip an $\mathbb{R}$-filtration with its each homogeneous part. If it is an approximable graded algebras, we will have some convergent properties.
 Let $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ be an approximable graded algebra.
@@
 $\lambda(x_{1}x_{2})\geqslant\lambda(x_{1})+\lambda(x_{2}),$
-where the function $\lambda(\raisebox{1.72218pt}{.})$ is defined at (1) and the homogeneous degrees are ignored, then we say $E_{\bullet}$ is filtrated with respect to $\mathcal{F}$.
+where the function $\lambda(\raisebox{1.72218pt}{.})$ is defined at (1) and the homogeneous degrees of $x_{1}$ and $x_{2}$ are ignored, then we say $E_{\bullet}$ is filtrated with respect to $\mathcal{F}$.
 For every $\epsilon>0$, let $T_{\epsilon}$ be the operator on the space of Borel probability measures on $\mathbb{R}$ by mapping $x\mapsto\epsilon x$.
@@
 […]
-With all the above notations. Suppose that $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ is a filtrated graded algebra with respect to $\lambda(\raisebox{1.72218pt}{.})$ defined at (1), and
+With all the above notations.
+Suppose that $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ is a filtrated approximable graded algebra with respect to $\lambda(\raisebox{1.72218pt}{.})$ defined at (1), and
 $\sup_{n\geqslant 1}\frac{\lambda_{\max}(E_{n})}{n}<+\infty,$
@@
 $\lambda_{+}^{\operatorname{asy}}(E_{\bullet}):=\int_{0}^{+\infty}x\nu_{E_{\bullet}}(\mathrm{d}x)=\lim\limits_{n\to\infty}\frac{\lambda_{+}(E_{n})}{n},$ (4)
-where $\lambda_{+}(\raisebox{1.72218pt}{.})$ is defined at (2). See [2, §2.4] for more details.
+where $\lambda_{+}(\raisebox{1.72218pt}{.})$ is defined at (2). See [4, §2.4] for more details.
 We have the following result.
@@
 Let $S$ be an adelic curve, $X$ be a projective scheme over $\operatorname{Spec}K$, and $\overline{L}$ be an adelic line bundle on $X$ over $S$.
-For our target, we will construct some graded linear series consisting of a particular kind of adelic vector bundles over $S$ from the above data.
+For our purpose, we will construct some graded linear series consisting of some particular adelic vector bundles over $S$ from above data.
 3.1 Adelic line bundles on an arithmetic variety
-In this part, we recall some notions about arithmetic positivity of the adelic line bundles on an arithmetic variety over adelic curve.
-We refer the readers to [9, §6 – §9] for a self-contained introduction to these notions.
+First, we recall some notions about arithmetic positivities of the adelic line bundles on arithmetic variety over an adelic curve, which are original from [8, 11].
+We refer the readers to [11, §6 – §9] for a self-contained introduction to these notions.
 3.1.1 Adelic line bundle
-Let $S=(K,(\mathcal{A},\Omega,\nu),\phi)$ be an adelic curve, and $X$ be a projective scheme over $\operatorname{Spec}K$.
-We say that $\overline{L}=(L,\varphi)$ is an adelic line bundle on $X$, if $L$ is a line bundle on $X$, and $\varphi=(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ is a family of metrics on $L_{w}$, which is dominated and measurable.
-In addition, we suppose that either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ contains a sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
-Then for every $w\in\Omega$ and $x\in X_{w}^{\operatorname{an}}$, we define
+Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, and $X$ be a projective scheme over $\operatorname{Spec}K$.
+We say that $\overline{L}=(L,\varphi)$ is an adelic line bundle on $X$, if $L$ is a line bundle on $X$, and $\varphi=(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ is a family of metrics on $L_{w}$ for each $w\in\Omega$, which is dominated and measurable.
+In addition, we suppose that either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ contains a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
+Then we construct a family of norms on $H^{0}(X,L)$ in the way below: for every $w\in\Omega$ and $x\in X_{w}^{\operatorname{an}}$, define
 |.$\begin{array}[]{rrcl}|\raisebox{1.72218pt}{.}|_{\varphi_{w}}(x):&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&|s|_{\varphi_{w}}(x)\end{array}$
-as the value of $s$ at the point $x$ with respect to the metric, and
+as the value of the global section $s$ at the point $x$ with respect to the given metric, and
 .$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\varphi_{w}}:&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&\sup\limits_{x\in X^{\operatorname{an}}_{w}}\{|s|_{\varphi_{w}}(x)\}.\end{array}$ (5)
 We denote $\varphi=(\|\raisebox{1.72218pt}{.}\|_{\varphi_{w}})_{w\in\Omega}$ the norm family on $H^{0}(X,L)$ over $S$ if there is no ambiguity.
-By [7, Theorem 6.1.13, Theorem 6.1.32], $(H^{0}(X,L),\varphi)$ is an adelic vector bundle over $S$.
+By [9, Theorem 6.1.13, Theorem 6.1.32], $(H^{0}(X,L),\varphi)$ is an adelic vector bundle over $S$.
 Let $\overline{L}=(L,\varphi)$ and $\overline{M}=(M,\psi)$ be two adelic line bundles on $X$.
-We denote by $\overline{L}\otimes\overline{M}=(L\otimes M,\varphi+\psi)$ the adelic line bundle, where $\varphi+\psi$ is the tensor product of metrics.
-In particular, for $n\in\mathbb{N}$, we denote by $\overline{L}^{\otimes n}=(L^{\otimes n},n\varphi)$ the adelic line bundle of $n$-times tensor product of $\overline{L}$, and by $(H^{0}(X,L^{\otimes n}),n\varphi)$ the induced adelic vector bundle over $S$.
-By [7, Proposition 6.1.28, Proposition 6.1.31], all the above constructions are well-defined.
-We refer the readers to [7, §6.1] for a precise explanation of §3.1.1, see also [9, §2.6 – 2.8].
+We denote by $\overline{L}\otimes\overline{M}=(L\otimes M,\varphi+\psi)$ the tensor adelic line bundle, where $\varphi+\psi$ is the tensor product of metrics.
+In particular, for $n\in\mathbb{N}$, we denote by $\overline{L}^{\otimes n}=(L^{\otimes n},n\varphi)$ the adelic line bundle of $n$-times tensor product of $\overline{L}$, and by $(H^{0}(X,L^{\otimes n}),n\varphi)$ the adelic vector bundle over $S$ defined above.
+By [9, Proposition 6.1.28, Proposition 6.1.31], all the above constructions are all well-defined.
+We refer the readers to [9, §6.1] for a precise explanation of §3.1.1, see also [11, §2.6 – §2.8].
 3.1.2 Semi-positive metric
-Let $S=(K,(\mathcal{A},\Omega,\nu),\phi)$ be an adelic curve, $X$ be a projective scheme over $\operatorname{Spec}K$, $L$ be a line bundle on $X$, $\varphi=(\varphi_{w})_{w\in\Omega}$ and $\psi=(\psi_{w})_{w\in\Omega}$ be two families of metrics on $L$ such that $(L,\varphi)$ and $(L,\psi)$ are both adelic line bundles.
+Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, $X$ be a projective scheme over $\operatorname{Spec}K$, $L$ be a line bundle on $X$, $\varphi=(\varphi_{w})_{w\in\Omega}$ and $\psi=(\psi_{w})_{w\in\Omega}$ be two families of metrics on $L$ such that $(L,\varphi)$ and $(L,\psi)$ are both adelic line bundles.
 Then we define the distance between $\varphi$ and $\psi$ as
 $d(\varphi,\psi)=\int_{\Omega}\sup_{x\in X^{\operatorname{an}}_{w}}\left|\log\frac{|\raisebox{1.72218pt}{.}|_{\varphi_{w}}(x)}{|\raisebox{1.72218pt}{.}|_{\psi_{w}}(x)}\right|\nu(\mathrm{d}w).$
-If $L$ is semiample, and if there exist a positive integer $m$ and a sequence $(\varphi_{n})_{n\in\mathbb{N}}$ of quotient metric families (where $\varphi_{n}$ a metric family of $L^{\otimes mn}$), such that
+If $L$ is semi-ample, and if there exist a positive integer $m$ and a sequence $(\varphi_{n})_{n\in\mathbb{N}}$ of quotient metric families (where $\varphi_{n}$ a metric family of $L^{\otimes mn}$), such that
 $\lim_{n\to\infty}\frac{1}{mn}d(nm\varphi,\varphi_{n})=0,$
@@
 Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$. We say $\overline{L}$ is relatively ample if $L$ is ample and $\varphi$ is semi-positive.
-In addition, an adelic line bundle $\overline{L}$ on $X$ is ample if it is relatively ample and if there exists an $\epsilon>0$ such that the inequality
+In addition, an adelic line bundle $\overline{L}$ on $X$ is ample, if it is relatively ample and if there exists an $\epsilon>0$ such that the inequality
 $(\overline{L}|_{Y}^{\dim(Y)+1})\geqslant\epsilon\deg_{L}(Y)(\dim(Y)+1)$
@@
 […]
-Compared with the arithmetically ample definition of S.
-Zhang in [26], an arithmetically ample Hermitian line bundle on an arithmetic variety requires the metrics on the line bundle to be smooth, while the definition in [9, Definition 9.1.1] over an adelic curve only require continuous metrics.
-In this article, we follow the definition given in [9, Definition 9.1.1].
+We compare the above definition with the arithmetically ample definition of S.
+Zhang in [29], where an arithmetically ample Hermitian line bundle on an arithmetic variety requires the metrics on the line bundle to be smooth.
+The definition in [11, Definition 9.1.1] over an adelic curve only require continuous metrics.
+In this article, we follow the definition given in [11, Definition 9.1.1] as above.
 3.1.4 Nefness
 Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$.
-We say that $\overline{L}$ is relatively nef if there exists a relatively ample adelic line bundle $\overline{A}$ on $X$ and a positive integer $N$ such that, for any $n\geqslant N$, the tensor product $\overline{L}^{\otimes n}\otimes\overline{A}$ is relatively ample.
-An adelic line bundle $\overline{L}$ is said to be nef or arithmetically nef if there exist an ample adelic line bundle $\overline{A}$ and a positive integer $M$ such that $\overline{L}^{\otimes m}\otimes\overline{A}$ is ample for any $m\geqslant M$ in the arithmetic sense.
+We say that $\overline{L}$ is relatively nef if there exists a relatively ample adelic line bundle $\overline{A}$ on $X$ and a positive integer $N$ such that, for any $n\geqslant N$, the tensor product of adelic line bundles $\overline{L}^{\otimes n}\otimes\overline{A}$ is relatively ample.
+An adelic line bundle $\overline{L}$ is said to be nef or arithmetically nef, if there exist an arithmetically ample line bundle $\overline{A}$ and a positive integer $M$ such that $\overline{L}^{\otimes m}\otimes\overline{A}$ is ample for any $m\geqslant M$ in the arithmetic sense.
 3.1.5 Effectivity
 Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$.
-If $s\in H^{0}(X,L)$ is a non-zero global section such that $\|s\|_{\varphi_{w}}\leqslant 1$ for any $w\in\Omega$, we say that the global section $s$ is effective.
-We say that $\overline{L}$ is effective or arithmetically effective if it admits at least one effective global section by the above definition.
+If $s\in H^{0}(X,L)$ is a non-zero global section which satisfies $\|s\|_{\varphi_{w}}\leqslant 1$ for all $w\in\Omega$, we say that the global section $s$ is effective.
+We say that an adelic line bundle $\overline{L}$ is effective or arithmetically effective, if it admits at least one effective global section by the above definition.
 3.2 Algebraic construction of graded linear series
-In this part we only consider the geometric setting, which means everything works over an arbitrary field $K$.
+In this part we will only consider the geometric setting, which means everything works over an arbitrary field with no metric.
 For every $D\in\mathbb{N}^{+}$, we consider the global section space $H^{0}(X,L^{\otimes D})$, and the graded $K$-algebra $\bigoplus\limits_{n\geqslant 0}H^{0}(X,L^{\otimes n})$.
@@
 3.2.1 Graded linear series containing an ample divisor
-The following definition is original from [16, Definition 2.9], where it is called "Condition (C)". This notion plays a key role in the arithmetic Fujita approximation of [2].
+The following definition is original from [18, Definition 2.9], where it is called "Condition (C)".
+This notion plays a key role in the arithmetic Fujita approximation of [4], and we follow the notation of [4, Definition 3.1].
 […]
@@
 Suppose that there exists a $p\in\mathbb{N}^{+}$, an ample line bundle $A$, an effective line bundle $M$ on $X$, and a non-zero section $s\in H^{0}(X,M)$, such that $L^{\otimes p}=A\otimes M$.
-We say a graded linear series $W_{\bullet}=\bigoplus\limits_{n\geqslant 0}W_{n}$ of $L$ contains an ample divisor if the homomorphism of graded algebras
+In this case, we say a graded linear series $W_{\bullet}=\bigoplus\limits_{n\geqslant 0}W_{n}$ of $L$ contains an ample divisor if the homomorphism of graded algebras
 $\bigoplus_{n\geqslant 0}H^{0}(X,A^{\otimes n})\rightarrow\bigoplus_{n\geqslant 0}H^{0}(X,L^{\otimes np})$
 induced by $s$ can factor through $\bigoplus\limits_{n\geqslant 0}W_{np}$.
-In §2.3, we have seen that an approximable graded linear series has many convergent properties.
+In §2.3.2, we have seen that a filtrated approximable graded linear series has many convergent properties.
 The following property is original from [18], which is applied in [4] to judge whether a graded linear series is approximable.
@@
 $\operatorname{vol}(W_{\bullet})=\limsup_{n\to\infty}\frac{\dim_{K}(W_{n})}{n^{\dim(X)}/\dim(X)!}.$
-If $W_{n}=H^{0}(X,L^{\otimes n})$, we denote $\operatorname{vol}(L)=\operatorname{vol}(W)$.
+If $W_{n}=H^{0}(X,L^{\otimes n})$, we denote $\operatorname{vol}(L)=\operatorname{vol}(W_{\bullet})$.
 If $\operatorname{vol}(L)>0$, we say that $L$ is a big line bundle on $X$.
 3.3 Arithmetic invariants of graded linear series
-In this part, we consider arithmetic constructions.
+In this part, we consider some arithmetic constructions.
 3.3.1
@@
 In §3.1.1, we defined a norm family on $H^{0}(X,L^{\otimes D})$ by (5) indexed by $\Omega$, which is denoted by $D\varphi$ with no ambiguity.
-By [7, §6.4.1], .$\|\raisebox{1.72218pt}{.}\|_{D\varphi_{w}}$ is ultrametric for every $w\in\Omega_{f}$.
+By [9, §6.4.1], .$\|\raisebox{1.72218pt}{.}\|_{D\varphi_{w}}$ is ultrametric for every $w\in\Omega_{f}$.
 Then $(H^{0}(X,L^{\otimes D}),D\varphi)$ is an adelic vector bundle over $S$.
-Let $\overline{E}_{D}$ be an adelic vector sub-bundle $H^{0}(X,L^{\otimes D})$ over $S$.
+For every $D\in\mathbb{N}$, let $\overline{E}_{D}$ be an adelic vector sub-bundle $H^{0}(X,L^{\otimes D})$ over $S$.
 In this case, we say that $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ is a normed graded linear series over $S$ if $E_{D}\cdot E_{D^{\prime}}\subseteq E_{D+D^{\prime}}$ for $D,D^{\prime}\in\mathbb{N}$.
@@
 $\widehat{\operatorname{vol}}(\overline{E}_{\bullet})=\limsup_{n\to\infty}\frac{\widehat{\deg}_{+}\left(\overline{E}_{n}\right)}{n^{\dim(X)+1}/(\dim(X)+1)!}.$ (6)
-In particular, if $\overline{E}_{D}=\left(H^{0}(X,L^{\otimes D}),D\varphi\right)$ as above, we denote $\widehat{\operatorname{vol}}(\overline{L})=\widehat{\operatorname{vol}}(\overline{E}_{\bullet})$ for simplicify.
-If $\widehat{\operatorname{vol}}(\overline{L})>0$, we say that $\overline{L}$ is big on $X$, or we call it arithmetically big to distinguish it from the geometric case.
+In particular, if $\overline{E}_{D}=\left(H^{0}(X,L^{\otimes D}),D\varphi\right)$ as above, we denote $\widehat{\operatorname{vol}}(\overline{L})=\widehat{\operatorname{vol}}(\overline{E}_{\bullet})$ for simplicity.
+If $\widehat{\operatorname{vol}}(\overline{L})>0$, we say that $\overline{L}$ is big on $X$, or arithmetically big to distinguish it from the geometric case.
 3.3.3
-For every $D\in\mathbb{N}$, Let $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E}_{D})\right)_{t\in\mathbb{R}}$ be the $\mathbb{R}$-filtration induced by the Harder-Narasimhan filtration of $\overline{E}_{D}$ following Definition 2.2.3, where we ignore the homogeneous degree $D\in\mathbb{N}$.
+For every $D\in\mathbb{N}$, let $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E}_{D})\right)_{t\in\mathbb{R}}$ be the $\mathbb{R}$-filtration induced by the Harder-Narasimhan filtration of $\overline{E}_{D}$ following Definition 2.2.3, where we ignore the homogeneous degree $D\in\mathbb{N}$ of $\overline{E}_{D}$.
 We denote by $\nu_{\overline{E}_{D}}$ the Borel probability measure induced by the Harder-Narasimhan filtration of $\overline{E}_{D}$ constructed in the sense of §2.2.2.
-Then we have the result below from Proposition 4.3.50, Proposition 4.3.51 and Corollary 4.3.52 of [7], see also [9, §2.5.4].
+Then we have the result below from Proposition 4.3.50, Proposition 4.3.51 and Corollary 4.3.52 of [9], see also [11, §2.5.4].
 […]
@@
 Let $K$ be a field, $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a sub-field which is dense in $K_{w}$ for all $w\in\Omega$, and $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ be the adelic curve of trivially valued field introduced in Example 2.1.1.
-In fact, for any normed vector space over $(K,|\raisebox{1.72218pt}{.}|_{0})$ of finite dimension, its norm takes only finitely many values (cf. [9, §3.1]).
-In this part, we will study the adelic vector bundles and graded linear series over $S_{0}$ following that in [9, §3]. Some constructions are from those over $S$.
+In fact, for any normed vector space over $(K,|\raisebox{1.72218pt}{.}|_{0})$ of finite dimension, its norm takes only finitely many values (cf. [11, §3.1]).
+In this part, we will study the adelic vector bundles and graded linear series over $S_{0}$ following that in [11, §3]. Some constructions are from those over $S$.
 3.4.1
-Let $\overline{V}$ be an adelic vector bundle over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, and $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{V})\right)_{t\in\mathbb{R}}$ be the Harder-Narasimhan filtration of $\overline{V}$ defined in Definition 2.2.3.
-We define the norm .$\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}$ on $V$ over $(K,|\raisebox{1.72218pt}{.}|_{0})$ as
-.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}:&V&\longrightarrow&\mathbb{R}^{+}\\ &s&\mapsto&\exp\left(-\sup\{t\in\mathbb{R}\mid s\in\mathcal{F}^{t}_{\operatorname{HN}}(\overline{V})\}\right)\end{array}.$ (7)
-Then $\overline{V}_{\operatorname{HN}}=\left(V,\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}\right)$ is an adelic vector bundle over $S_{0}$.
+Let $\overline{E}$ be an adelic vector bundle over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, and $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E})\right)_{t\in\mathbb{R}}$ be the Harder-Narasimhan filtration of $\overline{E}$ defined in Definition 2.2.3.
+We define the norm .$\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}$ on $E$ over $(K,|\raisebox{1.72218pt}{.}|_{0})$ as
+.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}:&E&\longrightarrow&\mathbb{R}^{+}\\ &s&\mapsto&\exp\left(-\sup\{t\in\mathbb{R}\mid s\in\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E})\}\right).\end{array}$ (7)
+Then $\overline{E}_{\operatorname{HN}}=\left(E,\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}\right)$ is an adelic vector bundle over $S_{0}$.
 3.4.2
-Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ be a normed graded linear series over $S=(K,(\Omega,\mathcal{A},\nu),\phi)$.
-For every $\overline{E}_{D}$, we denote by .$\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}}$ the norm (7) on $E_{D}$.
-Then $(E_{D},\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}})$ an Adélie vector bundle over $S_{0}$.
-By [9, Proposition 4.1.3], for every $s\in E_{m}$ and $t\in E_{n}$, we have
+Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ be a normed graded linear series over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$.
+For every $\overline{E}_{D}$, we denote by .$\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}}$ the norm (7) on $E_{D}$ indexed by $D\in\mathbb{N}$.
+Then $(E_{D},\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}})$ an adelic vector bundle over $S_{0}$.
+By [11, Proposition 4.1.3], for every $s\in E_{m}$ and $t\in E_{n}$, we have
 $\|s\cdot t\|^{\operatorname{HN}}_{m+n}\leqslant\exp(f(m)+f(n))\|s\|^{\operatorname{HN}}_{m}\|t\|^{\operatorname{HN}}_{n},$
-where $f(n)=\frac{3}{2}\nu(\Omega_{\infty})\log n$, and $\nu(\Omega_{\infty})$ is the measure of achimedean absolute values of $S=(K,(\Omega,\mathcal{A},\nu),\phi)$.
+where $f(n)=\frac{3}{2}\nu(\Omega_{\infty})\log n$, and $\nu(\Omega_{\infty})$ is the measure of archimedean absolute values of $S=(K,(\Omega,\mathcal{A},\nu),\phi)$.
+The term $f(n)$ is original from the partial response of the semi-stable conjecture of Bost [1, Theorem A], see also [8, §5.6].
 In order to obtain a new family of norms which are sub-multiplicative, we introduce the spectral norm .$\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ of .$\|\raisebox{1.72218pt}{.}\|^{\operatorname{HN}}_{D}$ as
-.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}:&H^{0}(X,L^{\otimes D})&\longrightarrow&\mathbb{R}_{\geqslant 0}\\ &s&\mapsto&\lim\limits_{n\rightarrow\infty}\left(\|s^{n}\|_{D}^{\operatorname{HN}}\right)^{1/n}.\end{array}$ (8)
-Then $\left(E_{D},\|\raisebox{1.72218pt}{.}\|_{D,\mathrm{sp}}\right)$ is an adelic vector bundle over $S_{0}$. By [9, Proposition 3.2.2 (3)], for every $s\in E_{m}$ and $t\in E_{n}$, we have
+.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}:&H^{0}(X,L^{\otimes D})&\longrightarrow&\mathbb{R}_{\geqslant 0}\\ &s&\mapsto&\lim\limits_{n\rightarrow\infty}\left(\|s^{n}\|_{nD}^{\operatorname{HN}}\right)^{1/n}.\end{array}$ (8)
+In this case, $\left(E_{D},\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}\right)$ is an adelic vector bundle over $S_{0}$. By [11, Proposition 3.2.2 (3)], for every $s\in E_{m}$ and $t\in E_{n}$, we have
 $\|s\cdot t\|_{\mathrm{sp},m+n}\leqslant\|s\|_{\mathrm{sp},m}\|t\|_{\mathrm{sp},n}.$
+3.4.3
 Let $\overline{E}_{\operatorname{HN},D}=\left(E_{D},\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}\right)$, then
-$\overline{E}_{\operatorname{HN},\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{\operatorname{HN},D}$ (9)
-is a filtrated graded linear series (see Definition 2.3.2 for the definition) over $S_{0}$, where we take $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$.
-We denote by $\mathcal{F}_{\mathrm{sp}}$ the $\mathbb{R}$-filtration on $\overline{E}_{\operatorname{HN},D}$ induced by the filtrated structure $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$, where we ignore the index $D\in\mathbb{N}$.
-3.4.3
+$\overline{E}_{\operatorname{HN},\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{\operatorname{HN},D}$
+is a filtrated graded linear series (see Definition 2.3.2 for the definition) over $S_{0}$, where we take $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ in Definition 2.3.2.
+We denote by $\mathcal{F}_{\mathrm{sp}}$ the $\mathbb{R}$-filtration on $\overline{E}_{\operatorname{HN},D}$ induced by the filtrated structure $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$, where we omit the index $D\in\mathbb{N}$ in the above notation.
+The existence of such $\mathcal{F}_{\mathrm{sp}}$ is ensured by the finiteness of the values of .$\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ on $\overline{E}_{\operatorname{HN},D}$.
+3.4.4
 Some asymptotic behaviors of $\overline{E}_{\operatorname{HN},\bullet}$ are useful for the study of $\overline{E}_{\bullet}$.
@@
 […]
-With all the notations in §3.4.2, we have
+With all the notations in §3.4.2 and §3.4.3, we have
 $\widehat{\operatorname{vol}}(\overline{E}_{\bullet})=\widehat{\operatorname{vol}}(\overline{E}_{\operatorname{HN},\bullet}).$
 Proof.
-It is from [9, Corollary 3.4.4, (4.1.2), Corollary-Definition 4.1.4] directly. ∎
-By Proposition 2.3.2, the sequence $\{\frac{1}{n}\widehat{\mu}_{\max}(\overline{E}_{n})\}_{n\geqslant 0}$ converges to a real number, and we denote it by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\bullet})$.
+It is from [11, Corollary 3.4.4, (4.1.2), Corollary-Definition 4.1.4] directly. ∎
+By Proposition 2.3.2, the sequence $\left(\frac{1}{n}\widehat{\mu}_{\max}(\overline{E}_{n})\right)_{n\geqslant 0}$ converges to a real number, and we denote it by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\bullet})$.
 If $E_{n}=H^{0}(X,L^{\otimes n})$ equipped with the superior norms, it is denoted by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$ for simplicity.
-We have the following result from [7, Proposition 6.4.18], which is an equivalent of arithmetically big adelic line bundle.
+The following result is an equivalent property of arithmetically bigness.
 […]
-Let $\overline{L}$ be an adelic line bundle on $X$ over an adelic curve $S$.
+Let $\overline{L}$ be an adelic line bundle on $X$ over a proper adelic curve $S$.
 If $L$ is a big line bundle on $X$, then $\widehat{\operatorname{vol}}(\overline{L})>0$ if and only if $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})>0$.
-From Proposition 3.3.3, we have the following result by an elementary calculation.
+In addition, we have
+$\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})=\lambda_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}).$
+Proof.
+The equivalence of $\widehat{\operatorname{vol}}(\overline{L})>0$ and $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})>0$ is from [9, Proposition 6.4.18] directly.
+The equality is from [11, Proposition 3.4.1].
+∎
+From Proposition 3.3.3 and Proposition 3.4.4, we have the following result by an elementary calculation.
 […]
-Let $\overline{E}^{\bullet}=\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$ be a normed graded linear series over $S$. If $\operatorname{vol}(E^{\bullet})>0$, then we have
-$\lambda_{+}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN}}^{\bullet})=\frac{\widehat{\operatorname{vol}}(\overline{E}^{\bullet})}{(d+1)\operatorname{vol}(E^{\bullet})},$
-where $\lambda_{+}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ is defined at (4) with respect to the canonical filtrated structure as (9) over $S_{0}$, and $d=\dim(X)$.
-3.4.4
-Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over an arbitrary curve $S$. We consider $\widehat{\mu}_{\min}(H^{0}(X,L^{\otimes n}),n\varphi)$, and we define
-$\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})=\liminf_{n\to\infty}\frac{\widehat{\mu}_{\min}(H^{0}(X,L^{\otimes n}),n\varphi)}{n}.$
-We recall some results on the criteria of arithmetic positivities via $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ in [9] below.
+Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$ be a normed graded linear series over $S$.
+If $\operatorname{vol}(E_{\bullet})>0$ following the definition in §3.2.2, then we have
+$\lambda_{+}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet})=\frac{\widehat{\operatorname{vol}}(\overline{E}_{\bullet})}{\left(\dim(X)+1\right)\operatorname{vol}(E_{\bullet})},$
+where $\lambda_{+}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ is defined at (4) with respect to the canonical filtrated structure introduced in §3.4.3 over $S_{0}$.
+3.4.5
+Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over an arbitrary proper curve $S$. We consider $\widehat{\mu}_{\min}(H^{0}(X,L^{\otimes n}),n\varphi)$, and we define
+$\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})=\liminf_{n\to\infty}\frac{\widehat{\mu}_{\min}(H^{0}(X,L^{\otimes n}),n\varphi)}{n},$
+where the adelic vector bundle $(H^{0}(X,L^{\otimes n}),n\varphi)$ over $S$ is defined in §3.1.1.
+We recall some results on the criteria of arithmetic positivities via $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ in [11, §9] below, where we always suppose that $K$ is perfect.
 […]
@@
 See §3.1.4 for the definitions of relatively nef and nef adelic line bundles.
-3.4.5
-The following result shows that a power of an arithmetically big line bundle over $S_{0}$ can be decomposed as the sum of an ample and an effective line bundles, which is an analogue of [22, Corollary 2.4].
-[…]
-Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be arithmetically big on $X$ over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$.
-Then there exists a $p\in\mathbb{N}$, such that $\overline{L}^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{M}$ is arithmetically ample and $\overline{N}$ is arithmetically effective.
-Proof.
-By [14, Corollary 2.2.7], there exists a $p_{1}\in\mathbb{N}$, such that for all $t>p_{1}$, we are able to find an ample line bundle $M$ and an effective line bundle $M$ satisfying $L^{\otimes t}\cong A\otimes M$.
-We equip $L$ with the metric $\varphi$ given in the statement such that $\overline{L}$ is arithmetically big, and equip $A$ with an arbitrary metric $\psi$ such that $\overline{M}$ is arithmetically ample.
-Then by [7, Corollary 6.4.10], there exists $p_{2}\in\mathbb{N}$, such that for all $t>p_{2}$, we have
-$\widehat{\deg}_{+}\left(H^{0}(X,L^{\otimes t}\otimes M^{\vee}),t\varphi-\psi\right)>0.$
-Since $S_{0}$ only has one place, then from the above inequality we deduce that $\overline{L}^{\otimes t}\otimes\overline{M}^{\vee}$ contains an effective section over $S_{0}$ in the sense of §3.1.5, so it is arithmetically effective over $S_{0}$ by the adelic structure of $S_{0}$.
-We choose $p\geqslant\max\{p_{1},p_{2}\}$ which satisfying the requirement. ∎
-[…]
-The geometric version of Proposition 3.4.5 is firstly proved by K.
-Kodaira, see [14, Proposition 2.2.6, Corollary 2.2.7] for a brief introduction.
-An arithmetic version over number fields was proved by X.
-Yuan [22, Theorem 2.3, Corollary 2.4] via Siu’s inequality.
-It is an interesting question whether Proposition 3.4.5 is verified over a general adelic curve, or add some technical conditions.
-In fact, over a general adelic curve, we cannot deduce that an adelic line bundle $(L,\varphi)$ on $X$ contains an effective section from the inequality $\widehat{\deg}_{+}\left((H^{0}(X,L),\varphi)\right)>0$ directly.
 4 Fujita approximation of graded linear series
-In this section, we denote $E_{D}=H^{0}(X,L^{\otimes D})$ for simplicity, and all the constructions will be based on this notation.
+In this section, we will always denote $E_{D}=H^{0}(X,L^{\otimes D})$ for simplicity, and all the constructions will be based on this notation.
 4.1 Approximatible graded linear series
 Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable which is dense in $K_{w}$ for all $w\in\Omega$.
-Let $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ be the adelic curve of trivially valued field.
-In this part, we will consider some approximated properties of normed graded linear series over $S$ and $S_{0}$.
+Let $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ be the adelic curve of trivially valued field defined in Example 2.1.1.
+In this part, we will consider some approximated properties of normed graded linear series over $S_{0}$ constructed from that over $S$.
+4.1.1
 Let $\lambda\in\mathbb{R}_{\geqslant 0}$, $\overline{E}_{\operatorname{HN},0}^{\lambda}=K$, and
-$\overline{E}_{\operatorname{HN},D}^{\lambda}=\operatorname{Vect}_{K}\left\{s\in H^{0}(X,L^{\otimes D})\mid\|s\|_{\mathrm{sp},D}\leqslant\exp(-D\lambda)\right\}$
+$\overline{E}_{\operatorname{HN},D}^{\lambda}=\operatorname{Vect}_{K}\left\{s\in H^{0}(X,L^{\otimes D})\mid\|s\|_{\mathrm{sp},D}\leqslant\exp(-D\lambda)\right\}$ (9)
 for $D\geqslant 1$. So we have
 $\overline{E}_{\operatorname{HN},D}^{\lambda}=\mathcal{F}_{\mathrm{sp}}^{D\lambda}\overline{E}_{\operatorname{HN},D},$
-where the $\mathbb{R}$-filtration $\mathcal{F}_{\mathrm{sp}}$ is defined in §2.3.2, and $\overline{E}_{\operatorname{HN},D}$ follows the definition in (9) by taking $E_{D}=H^{0}(X,L^{\otimes D})$.
+where the $\mathbb{R}$-filtration $\mathcal{F}_{\mathrm{sp}}$ is defined in §2.3.2, and $\overline{E}_{\operatorname{HN},D}$ follows the definition in §3.4.3 by taking $E_{D}=H^{0}(X,L^{\otimes D})$.
 In addition, we define
@@
 which is a graded linear series over $S_{0}$.
-We will prove that (10) is approximable from the fact that it contains an ample divisor, where the strategy is similar to [2, Proposition 3.12].
-4.1.1
-We have the following proposition from the definition directly.
+4.1.2
+By the properties of $\overline{E}_{\operatorname{HN},\bullet}$ in §3.4.3, we have the following proposition.
 […]
-(10) is a filtrated graded linear series over $S_{0}$.
-4.1.2
-For $\lambda\in\mathbb{R}$, we define the adelic line bundle
-$\overline{\mathcal{O}}_{\lambda}=(K,\|\raisebox{1.72218pt}{.}\|_{\lambda})$ (11)
-over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$, where the norm satisfies $\|\mathbbm{1}\|_{\lambda}=e^{-\lambda}$.
+The graded linear series (10) is a filtrated graded linear series over $S_{0}$.
+4.1.3
+Next, we prove that $\overline{E}_{\operatorname{HN},D}^{\lambda}$ is approximable when $\lambda$ is of some particular values.
+For this target, we will prove it contains an ample divisor, and see Definition 4.1.3.
 […]
-Let $\pi:X\rightarrow\operatorname{Spec}K$ be a projective morphism, and $\overline{L}$ be an adelic line bundle over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$.
+Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
+Let $\pi:X\rightarrow\operatorname{Spec}K$ be a projective morphism, and $\overline{L}$ be an adelic line bundle over $S$.
 With all the notations and assumption same as in (10).
-If $\overline{L}$ is arithmetically big, then for all $\lambda\in\left[0,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$, $\overline{E}^{\lambda,\bullet}_{\operatorname{HN}}$ contains an ample divisor.
+If $\overline{L}$ is big, then for all $\lambda\in\left(-\infty,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$, $\overline{E}^{\lambda}_{\operatorname{HN},\bullet}$ contains an ample divisor.
 Proof.
-By the choice of $\lambda$, we have $\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L}\otimes\pi^{*}\overline{\mathcal{O}}_{-\lambda})>0$ by definition directly, so $\overline{L}\otimes\pi^{*}\overline{\mathcal{O}}_{-\lambda}$ is an arithmetically big line bundle on $X$ by Proposition 3.4.3.
-Then for a sufficiently large $D\in\mathbb{N}$, $L^{\otimes D}$ has a global section $s_{D}$ satisfying $\|s_{D}\|_{\mathrm{sp},D}\leqslant\exp(-D\lambda)$, which proves $\overline{E}_{\operatorname{HN},D}^{\lambda}\neq 0$ for $D\gg 0$.
-Since $\overline{L}\otimes\pi^{*}\overline{\mathcal{O}}_{-\lambda}$ is arithmetically big on $X$, then by Propopsition 3.4.5, there exists $p\in\mathbb{N}$, such that $\left(\overline{L}\otimes\pi^{*}\overline{\mathcal{O}}_{-\lambda}\right)^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{M}$ is arithmetically ample and $\overline{N}$ is arithmetically effective, in other words $\overline{M}$ contains an effective section $s$.
-The remainder part of this proof is quite similar to that of [2, Proposition 3.12].
-We are able to take the above $p\in\mathbb{N}$ divisible enough such that the graded $K$-algebra $\bigoplus\limits_{n\in\mathbb{N}}H^{0}(X,M^{\otimes n})$ is generated by effective sections.
-These sections are considered as those of $\overline{M}\otimes\pi^{*}(\overline{\mathcal{O}}_{\lambda}^{\otimes p})$, whose norms are smaller than $\exp(-\lambda p)$.
+By Proposition 3.4.4 and Proposition 3.4.4, from the fact that $\overline{L}$ is big, we have $\overline{E}_{\operatorname{HN},\bullet}$ is big.
+In addition, by the choice of $\lambda$, we have $\lambda^{\operatorname{asy}}_{\max}(\overline{E}_{\operatorname{HN},\bullet}^{\lambda})>0$ by definition directly.
+Then for a sufficiently large $n\in\mathbb{N}$, $L^{\otimes n}$ has a global section $s\in H^{0}(X,L^{\otimes n})$ satisfying $\|s\|_{\mathrm{sp},n}\leqslant\exp(-n\lambda)$, which proves $\overline{E}_{\operatorname{HN},n}^{\lambda}\neq 0$ for $n\in\mathbb{N}$ sufficiently large.
+By Kodaira decomposition [16, Corollary 2.2.7], there exists a $p_{0}\in\mathbb{N}$, such that for every $p\in\mathbb{N}\cap[p_{0},+\infty)$, there always exist an ample line bundle $M$ and an effective line bundle $N$, which satisfies $L^{\otimes p}=M\otimes N$.
+We may choose the above $p$ large enough, such that $\bigoplus\limits_{D\geqslant 0}H^{0}(X,M^{\otimes D})$ is generated by global sections $t_{1},\ldots,t_{l}\in H^{0}(X,M)$, which is also considered as sections in $H^{0}(X,L^{\otimes d})$.
+We denote by $\lambda_{0}=\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$ for simplicity.
+Then for all $\epsilon>0$, there exist a $k\in\mathbb{N}$ and an $s\in H^{0}(X,L^{\otimes k})$, such that
+$\|s\|_{\mathrm{sp},pk}<\exp\left(-(\lambda_{0}-\epsilon/2)k\right).$
+Then there exists a $d\in\mathbb{N}$, such that
+$\|t_{i}s^{d}\|_{\mathrm{sp},kd+p}<\exp\left(-(\lambda_{0}-\epsilon)(kd+p)\right),$
+where $i=1,\ldots,l$.
+Let $G_{\bullet}^{\epsilon}$ be the graded linear series generated by $s^{D}t_{1},\ldots,s^{D}t_{l}$, which is a sub-graded algebra of $E_{\operatorname{HN},\bullet}$.
+Then for $D\in\mathbb{N}$, we have
+$G_{D}^{\epsilon}=\operatorname{Vect}_{K}\left\{(s^{d}t_{1})^{a_{1}}\cdots(s^{d}t_{l})^{a_{l}}\mid a_{1}+\cdots+a_{l}=D\right\}=s^{dD}H^{0}(X,M^{\otimes D}),$
+which is a sub-$K$-algebra of $E_{\operatorname{HN},(kd+p)D}$.
+Then for all $\lambda\in(-\infty,\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L}))$, we embed $H^{0}(X,M^{\otimes D})$ into $E_{\operatorname{HN},pD}^{\lambda}$ via $s^{dD}$.
 Then the homomorphism
 $\bigoplus_{D\geqslant 0}H^{0}(X,M^{\otimes D})\longrightarrow\bigoplus_{D\geqslant 0}H^{0}(X,L^{\otimes pD})$
-induced by $s$ factors through $\bigoplus\limits_{n\geqslant 0}\overline{E}_{\operatorname{HN},np}^{\lambda}$. ∎
-Same as the argument in [2, Corollary 3.13], we have the following result.
+factors through $\bigoplus\limits_{n\geqslant 0}E_{\operatorname{HN},np}^{\lambda}$. ∎
 […]
-The graded linear series (10) over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ is approximable for any $\lambda\in\left[0,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$.
+In [4, Proposition 3.11], a similar result of Proposition 4.1.3 is proved by an arithmetic analogue of Kodaira decomposition [25, Corollary 2.4].
+Until the author’s knowledge, we do not know whether it is still valid over a general adelic curve.
+To avoid this obstruction, we use a method quite similar to that of [2], and see also [9, §6.3].
+Same as the argument in [4, Corollary 3.13], we have the following result.
+[…]
+We keep all the notations and conditions in Proposition 4.1.3.
+The graded linear series (10) over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ is approximable for any $\lambda\in\left(-\infty,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$.
 Proof.
@@
 4.2 Fujita approximation of graded linear series
-Let $p\in\mathbb{N}$ such that $E_{\operatorname{HN},p}^{0}\neq 0$, and $\overline{E}_{\bullet}^{(p)}=\bigoplus\limits_{n\geqslant 0}\overline{E}^{(p)}_{\operatorname{HN},np}$ be the graded $K$-algebra generated by $E_{\operatorname{HN},p}^{0}$, which is considered as a normed graded linear series over $S_{0}$.
-At the same time, it is also able to be considered as a normed graded linear series over a general proper adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ equipped with the subspace norms of $\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$.
-In the both senses, they have the same arithmetic volume by Proposition 3.4.3 if $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
-The following result is an analogue of [2, Theorem 4.1] over $S_{0}$, whose proof is quite similar to the ancient one.
+Let $p\in\mathbb{N}$ such that $E_{\operatorname{HN},p}^{0}\neq 0$, where we follow the notation in (9).
+In this case, we define $\overline{E}_{\bullet}^{(p)}=\bigoplus\limits_{n\geqslant 0}\overline{E}^{(p)}_{\operatorname{HN},np}$ to be the graded $K$-algebra generated by $E_{\operatorname{HN},p}^{0}$, which is considered as a normed graded linear series over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ defined in Example 2.1.1.
+The following result is an analogue of [4, Theorem 4.1] over $S_{0}$, whose proof is quite similar to the ancient one.
 […]
@@
 With all the above notations and constructions, we have
-$\widehat{\operatorname{vol}}(\overline{L})=\sup_{p\in\mathbb{N}}\widehat{\operatorname{vol}}(\overline{E}_{\bullet}^{(p)}),$
-where $\overline{E}_{\bullet}^{(p)}$ is defined as above, which is a normed graded linear series over $S_{0}$.
+$\widehat{\operatorname{vol}}(\overline{L})=\sup_{p\in\mathbb{N}}\widehat{\operatorname{vol}}(\overline{E}_{\bullet}^{(p)}).$
 Proof.
@@
 Similarly, let $\nu_{\mathrm{sp},n}^{(p)}=T_{\frac{1}{np}}\nu_{\left(E^{(p)}_{\operatorname{HN},np},\mathcal{F}_{\mathrm{sp}}\right)}$, which also converges weakly to a Borel probability measure noted by $\nu_{\mathrm{sp}}^{(p)}$ by Proposition 4.1.2.
-In order to prove the assertion, we will prove that the restriction of $\nu_{\mathrm{sp}}$ on $[0,+\infty)$ is able to be approximated well by $\nu_{\mathrm{sp}}^{(p)}$.
+In order to prove the assertion, we will prove that the restriction of $\nu_{\mathrm{sp}}$ on $[0,+\infty)$ can be approximated well by $\nu_{\mathrm{sp}}^{(p)}$ for $p\in\mathbb{N}$.
 Step I. – Simultaneous approximations.
@@
 […]
-be a division of the interval $[0,\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L}))$ such that
+be a division of the interval $[0,\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}))$ such that
 $\nu_{\mathrm{sp}}(\{t_{1},\ldots,t_{m}\})=0.$ (11)
@@
 Step III. – Estimates of arithmetic volumes.
-By Proposition 3.4.3, we have
+Let $d=\dim(X)$. By Proposition 3.4.4, we have
 $\lim_{n\to\infty}\frac{d+1}{(np)^{d+1}/(d+1)!}\dim_{K}(E_{\operatorname{HN},np}^{(p)})\int_{0}^{+\infty}t\nu_{\mathrm{sp},np}^{(p)}(\mathrm{d}t)$
@@
 $=$ $(d+1)(1-\epsilon)\operatorname{vol}(L)\int_{\mathbb{R}}h_{D}\mathrm{d}\nu_{\mathrm{sp}},$
-where the last equality is from [1, Chap. IV §5 $n^{\circ}$ 12 Proposition 22].
+where the last equality is from [3, Chap. IV §5 $n^{\circ}$ 12 Proposition 22].
 We choose a sequence of subdivisions $(D_{j})_{j\in\mathbb{N}}$ satisfying (11), such that the sequence of functions $(h_{D_{j}}(t))_{j\in\mathbb{N}}$ converges uniformly to
-$\max\{t,0\}-\max\{t-\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L}),0\}$
-when $j\to\infty$. By definition, the support of $\nu_{\mathrm{sp}}$ is bounded from above by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$, we obtain
+$\max\{t,0\}-\max\{t-\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}),0\}$
+when $j\to\infty$. By definition, the support of $\nu_{\mathrm{sp}}$ is bounded from above by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet})$, we obtain
 $\sup_{p\in\mathbb{N}}\widehat{\operatorname{vol}}(E_{\bullet}^{(p)})\geqslant(d+1)(1-\epsilon)\operatorname{vol}(L)\int_{0}^{+\infty}t\nu_{\mathrm{sp}}(\mathrm{d}t)=(1-\epsilon)\widehat{\operatorname{vol}}(\overline{L}),$
@@
 For a general graded linear series, if it satisfies the property of Proposition 4.2, we say that this graded linear series satisfies Fujita approximation.
-We would like to emphasize that this notation is different from those in [5, Definition 3.7] and [6, §6].
+We would like to emphasize that this notion is different from those in [7, Definition 3.7] or [8, §6], though they are somewhat similar.
 5 Fujita approximation of adelic line bundles
-From this Fujita approximation of graded linear series proved in Proposition 4.2, we will prove the arithmetic Fujita approximation of an arithmetically big line bundle.
+From the Fujita approximation of graded linear series proved in Proposition 4.2, we will prove the arithmetic Fujita approximation of an arithmetically big line bundle with coefficients in $\mathbb{Z}$, $\mathbb{Q}$ and $\mathbb{R}$.
 5.1 Fujita approximation of big adelic line bundles
-In this part, we will prove the arithmetic Fujita approximation of usual adelic line bundles. First, we prove a nef version.
+First, we will prove the arithmetic Fujita approximation of usual adelic line bundles.
+We will prove the nef version first, and apply it to prove the ample version.
+In order to apply Proposition 3.4.5, we assume the base field is perfect.
 […]
-Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
+Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
 Let $X$ be a projective variety with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big line bundle on $X$ over $S$.
@@
 Proof.
-We devide the proof into three parts.
+We devide the proof into three steps.
 Step I. – Determining the geometric decomposition.
@@
 be the blow-up along the base locus of $E_{\operatorname{HN},p}^{0}$, which means
-$X_{p}=\operatorname{Proj}\left(\operatorname{Im}\left(\bigoplus_{n\geqslant 0}\pi^{*}(E_{\operatorname{HN},np}^{p})\longrightarrow\bigoplus_{n\geqslant 0}L^{\otimes np}\right)\right).$
+$X_{p}=\operatorname{Proj}\left(\operatorname{Im}\left(\bigoplus_{n\geqslant 0}\pi^{*}(E_{\operatorname{HN},np}^{(p)})\longrightarrow\bigoplus_{n\geqslant 0}L^{\otimes np}\right)\right).$
 Let $G_{p}=\mathcal{O}_{X_{p}}(1)$, and $M_{p}$ be the line bundle on $X_{p}$ with respect to the exceptional divisor.
@@
 Step II. – Constructing the metrics of line bundles.
+We consider $\overline{E}^{(p)}_{\bullet}$ as a normed graded linear series over $S$, which is equipped with the subspace norms of $\overline{E}_{\bullet}=\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$.
 For any $n\geqslant 1$ and $w\in\Omega$, let .$\|\raisebox{1.72218pt}{.}\|_{n,w}$ be the quotient norm on $G_{p,w}=G_{p}\times_{\operatorname{Spec}K}\operatorname{Spec}K_{w}$ induced by the surjective homomorphism
@@
 By Hahn-Banach theorem, every linear form on $\ker(E_{\operatorname{HN},np}^{(p)}\rightarrow\ell_{n,x})$ can be extended to a linear form on $E_{\operatorname{HN},np}^{(p)}$.
-In other words, there exists a section $w\in E_{\operatorname{HN},np}^{(p)}$ whose image in $G_{p,w}^{\otimes n}(x)$ has norm $\|w\|_{\sup,w}$.
+In other words, there exists a section $w\in E_{\operatorname{HN},np}^{(p)}$, whose image in $G_{p,w}^{\otimes n}(x)$ has norm $\|w\|_{\sup,w}$.
 As a section of $L_{w}^{\otimes np}$ on $X_{v}(\overline{K}_{w})$, we have $\|w(x)\|_{w}\leqslant\|w\|_{\sup,w}$, where .$\|\raisebox{1.72218pt}{.}\|_{w}$ is the metric on $\overline{L}$ indexed by $v\in\Omega$.
@@
 $\|u(x)\|_{n,w}=\|w\|_{\sup,w}^{1/n}\geqslant\|w(x)\|_{w}^{1/n}=\|u(x)\otimes s(x)\|_{w}.$
-Let $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{W\in\Omega}$ be an adelic metric on $G_{p}$, and we define $(M_{p},\beta_{n})=\phi_{p}^{*}(\overline{L})\otimes(G_{p},\alpha_{n})^{\vee}$.
-In this sense, $s$ is an effective section in $(M_{p},\beta_{n})$, then it is an effective line bundle.
+Let $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ be a family of adelic metrics on $G_{p}$, and we define $(M_{p},\beta_{n})=\phi_{p}^{*}(\overline{L})\otimes(G_{p},\alpha_{n})^{\vee}$.
+In this sense, $s$ is an effective section in $(M_{p},\beta_{n})$, then it is an effective adelic line bundle.
 Step III. – The relation among arithmetic volumes.
@@
 where
-$\widehat{\operatorname{vol}}(E_{\bullet}^{(p)},\alpha_{n})=\lim_{m\rightarrow\infty}\frac{\widehat{\deg}_{+}(E_{\operatorname{HN},mp}^{(p)},(\|\raisebox{1.72218pt}{.}\|_{n,w,\sup})_{w\in\Omega})}{m^{d+1}/(d+1)!},$
-and .$\|\raisebox{1.72218pt}{.}\|_{n,w,\sup}$ is the superior norm induced by $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ on $G_{p}$. Then by [2, Corollary A.2], we have
+$\widehat{\operatorname{vol}}(E_{\bullet}^{(p)},\alpha_{n})=\lim_{m\rightarrow\infty}\frac{\widehat{\deg}_{+}(E_{\operatorname{HN},mp}^{(p)},(\|\raisebox{1.72218pt}{.}\|_{n,\sup,w})_{w\in\Omega})}{m^{d+1}/(d+1)!},$
+and .$\|\raisebox{1.72218pt}{.}\|_{n,\sup,w}$ is the superior norm induced by $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ on $G_{p}$. Then by [4, Corollary A.2], we have
 $\limsup_{n\to\infty}\widehat{\operatorname{vol}}(E_{\bullet}^{(p)},\alpha_{n})\geqslant\widehat{\operatorname{vol}}(\overline{E}_{\bullet}^{(p)}).$
@@
 for infinitely many $n\geqslant N_{0}$.
-In addition, by Proposition 3.4.4 we obtain that $(G_{p},\alpha_{n})$ is arithmetically nef, since $\widehat{\mu}_{\min}(\overline{E}_{\operatorname{HN},np}^{(p)})\geqslant 0$ for all $n\geqslant 1$, then we deduce the positive asymptotic minimal slope.
+In addition, since $\widehat{\mu}_{\min}(\overline{E}_{\operatorname{HN},np}^{(p)})\geqslant 0$ for all $n\geqslant 1$, then it has a positive asymptotic minimal slope.
+Then by Proposition 3.4.5, we obtain that $(G_{p},\alpha_{n})$ is arithmetically nef with the fact that $K$ is perfect.
 To sum up, we obtain the assertion. ∎
-Next, we obtain the following ample version from Theorem 5.1 directly.
+Next, we obtain the following ample version from Theorem 5.1 directly. The idea is absorbed from [4, Remark 4.4] and [11, Corollary 8.5.7].
 […]
@@
 Let $\overline{B}$ be an arithmetically big line bundle on $X$.
-Then by [7, Corollary 6.4.10], for $m\in\mathbb{N}$ sufficiently large, we have $\overline{L}^{\otimes m}\otimes\overline{B}^{\vee}$ is arithmetically big.
+Then by [9, Corollary 6.4.10], for $m\in\mathbb{N}$ sufficiently large, we have $\overline{L}^{\otimes m}\otimes\overline{B}^{\vee}$ is arithmetically big.
 By Theorem 5.1, for any $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$, a $p\in\mathbb{N}$, and a decomposition
@@
 $(mpq)^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G}^{\prime})\geqslant(mpq)^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G}^{\otimes q})\geqslant m^{-(d+1)}(\widehat{\operatorname{vol}}(\overline{L}^{\otimes m}\otimes\overline{B}^{\vee})-\epsilon).$
-By the continuity of the arithmetic volume function (see [7, Corollary 6.4.10]), we have the result. ∎
+By the continuity of the arithmetic volume function (see [9, Corollary 6.4.10]), we have the result. ∎
 5.2 Case of $\mathbb{K}$-line bundles
-In this part we will prove the arithmetic Fujita approximation is verified for adelic $\mathbb{Q}$-line bundles and $\mathbb{R}$-line bundles over adelic curve.
+In this part we will prove the arithmetic Fujita approximation is verified for adelic $\mathbb{Q}$-line bundles and $\mathbb{R}$-line bundles over adelic curves.
 5.2.1
-First, we prove the case of adelic $\mathbb{Q}$-line bundle, which is a quite straightforward from the usual case.
+First, we prove the case of adelic $\mathbb{Q}$-line bundle, which is a quite straightforward corollary from the usual case.
 […]
-Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
-Let $X$ be a projective variety over $\operatorname{Spec}K$ with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big adelic $\mathbb{Q}$-line bundle on $X$ over $S$.
+Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
+Let $X$ be a projective variety over $\operatorname{Spec}K$ with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big $\mathbb{Q}$-line bundle on $X$ over $S$.
 For every $\epsilon>0$, there exists a birational morphism $\phi:X^{\prime}\rightarrow X$ together with a $p\in\mathbb{N}^{+}$, such that the decomposition $\phi^{*}(\overline{L}^{\otimes p})=\overline{M}\otimes\overline{G}$ satisfies:
   1. (1)
-    $\overline{M}$ and $\overline{G}$ are adelic $\mathbb{Q}$-line bundles, where $\overline{M}$ is effective and $\overline{G}$ is arithmetically ample;
+    $\overline{M}$ and $\overline{G}$ are adelic line bundles, where $\overline{M}$ is effective and $\overline{G}$ is arithmetically ample;
   2. (2)
@@
 Proof.
-By definition, there exists an $m\in\mathbb{N}$, such that $\overline{L}^{\otimes m}$ is a usual arithmetically big adelic line bundle.
+By definition, there exists an $m\in\mathbb{N}$, such that $\overline{L}^{\otimes m}$ is a usual arithmetically big line bundle.
 Then we obtain the assertion by applying Corollary 5.1 to $\overline{L}^{\otimes m}$.
@@
 5.2.2
-Next, we prove the case of adelic $\mathbb{R}$-line bundle, and see [7, §6.4.3] for an introduction to this notion.
+Next, we prove the case of adelic $\mathbb{R}$-line bundle, and see [9, §6.4.3] for an introduction to this notion.
 For this target, we prove an auxiliary lemma first, which allows us to approximate an arithmetically big $\mathbb{R}$-line bundle by an arithmetically big $\mathbb{Q}$-line bundle along an effective direction.
@@
 The adelic $\mathbb{R}$-line bundle $\overline{L}$ is arithmetically big, so $L$ is a big $\mathbb{R}$-line bundle.
-By an equivalent definition [14, Definition 2.2.21], we have $L=\sum\limits_{i=1}^{r}a_{i}L_{i}$, where $a_{1},\ldots,a_{r}>0$ and $L_{1},\ldots,L_{r}$ are usual big line bundles.
-By Kodaira’s decomposition [14, Corollary 2.2.7], there exists an $m\in\mathbb{N}$, such that for all $i=1,\ldots,r$, we have $mL_{i}=A_{i}+M_{i}$, where $A_{i}$ is ample and $M_{i}$ is effective.
-We equip metrics on $A_{1},\ldots,A_{r},M_{1},\ldots,M_{r}$ over $S$, such that $\overline{A}_{1},\ldots,\overline{A}_{r}$ are arithmetically ample, and $\overline{M}_{1},\ldots,\overline{M}_{r}$ are arithmetically effective.
-Then for any $b_{1},\ldots,b_{r},c_{1},\ldots,c_{r}\in\mathbb{R}^{+}$, we have $b_{1}\overline{A}_{1}+\cdots+b_{r}\overline{A}_{r}+c_{1}\overline{M}_{1}+\cdots+c_{r}\overline{M}_{r}$ is effective.
-By the continuity of $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ from [7, Corollary 6.4.24], for all $\epsilon>0$, there exist $\delta_{1},\ldots,\delta_{r},\eta_{1},\ldots,\eta_{r}>0$, such that for all $t_{i}\in(0,\delta_{i})$, $s_{i}\in(0,\eta_{i})$ for $i=1,\ldots,r$, we always have
-$0<\widehat{\operatorname{vol}}(m\overline{L})-\widehat{\operatorname{vol}}\left(m\overline{L}-\sum_{i=1}^{r}t_{i}\overline{A}_{i}-\sum_{i=1}^{r}s_{i}\overline{M}_{i}\right)<\epsilon.$
-If there does not exist an arithmetically big $\mathbb{Q}$-line bundle satisfying the requirements of the lemma, then $m\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{A}_{i}-\sum\limits_{i=1}^{r}s_{i}\overline{M}_{i}$ can never be a $\mathbb{Q}$-line bundle for all $t_{i}\in(0,\delta_{i})$ and $s_{i}\in(0,\eta_{i})$ $i=1,\ldots,r$.
+By an equivalent definition [16, Definition 2.2.21], we have $L=\sum\limits_{i=1}^{r}a_{i}L_{i}$, where $a_{1},\ldots,a_{r}>0$ and $L_{1},\ldots,L_{r}$ are usual big line bundles.
+By definition, $L_{1},\ldots,L_{r}$ are all effective in the geometric sense.
+We equip metrics on $L_{1},\ldots,L_{r}$ over $S$, such that $\overline{L}_{1},\ldots,\overline{L}_{r}$ are arithmetically effective.
+Then for any $b_{1},\ldots,b_{r}\in\mathbb{R}^{+}$, we have $b_{1}\overline{L}_{1}+\cdots+b_{r}\overline{L}_{r}$ is effective.
+By the continuity of $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ from [9, Corollary 6.4.24], for all $\epsilon>0$, there exist $\delta_{1},\ldots,\delta_{r}>0$, such that for all $t_{i}\in(0,\delta_{i})$, $i=1,\ldots,r$, we always have
+$0<\widehat{\operatorname{vol}}(\overline{L})-\widehat{\operatorname{vol}}\left(\overline{L}-\sum_{i=1}^{r}t_{i}\overline{L}_{i}\right)<\epsilon.$
+If there does not exist any arithmetically big $\mathbb{Q}$-line bundles satisfying the requirements of the lemma, then $\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ can never be a $\mathbb{Q}$-line bundle for all $t_{i}\in(0,\delta_{i})$, $i=1,\ldots,r$.
 It means there is no $\mathbb{Q}$-point in the region
-$(ma_{1}-\delta_{1},ma_{1})\times\cdots\times(ma_{r}-\delta_{r},ma_{r})\times(ma_{1}-\eta_{1},ma_{1})\times\cdots\times(ma_{r}-\eta_{r},ma_{r})\subset\mathbb{R}^{2r},$
-which has a strict positive Lebesgue measure. Then it contradicts to the fact that $\mathbb{Q}^{2r}$ is dense in $\mathbb{R}^{2r}$ with respect to the Euclidean topology.
-By choosing the constant $\epsilon$ sufficiently small, we pick a family $t_{1},\ldots,t_{r},s_{1},\ldots,s_{r}\in\mathbb{R}^{+}$ such that $m\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{A}_{i}-\sum\limits_{i=1}^{r}s_{i}\overline{M}_{i}$ is an arithmetically big $\mathbb{Q}$-line bundle.
-By the above construction, $\sum\limits_{i=1}^{r}t_{i}\overline{A}_{i}+\sum\limits_{i=1}^{r}s_{i}\overline{M}_{i}$ is effective.
+$(a_{1}-\delta_{1},a_{1})\times\cdots\times(a_{r}-\delta_{r},a_{r})\subset\mathbb{R}^{r},$
+which has a strict positive Lebesgue measure. Then it contradicts to the fact that $\mathbb{Q}^{r}$ is dense in $\mathbb{R}^{r}$ with respect to the Euclidean topology.
+By choosing the constant $\epsilon$ sufficiently small, we pick a family $t_{1},\ldots,t_{r}\in\mathbb{R}^{+}$ such that $\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ is an arithmetically big $\mathbb{Q}$-line bundle.
+By the above construction, $\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ is effective.
 Let
-$\overline{L}^{\prime}=\overline{L}-\frac{1}{m}\sum\limits_{i=1}^{r}t_{i}\overline{A}_{i}-\frac{1}{m}\sum\limits_{i=1}^{r}s_{i}\overline{M}_{i},$
+$\overline{L}^{\prime}=\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i},$
 which satisfies the requirement. ∎
@@
 […]
-Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
+Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve with a perfect $K$, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
 Let $X$ be a projective variety over $\operatorname{Spec}K$ with $\dim(X)=d$, and $\overline{L}$ be an arithmetically big adelic $\mathbb{R}$-line bundle on $X$ over $S$.
@@
   1. (1)
-    $\overline{M}$ and $\overline{G}$ are adelic $\mathbb{R}$-line bundles, where $\overline{M}$ is effective and $\overline{G}$ is arithmetically ample;
+    $\overline{M}$ is an arithmetically effective $\mathbb{R}$-line bundle, and $\overline{G}$ is an arithmetically ample line bundle;
   2. (2)

### v3

@@
 L’approximation arithmétique de Fujita sur courbes adéliques
-Abstract
+[…]
+L’approximation arithmétique de Fujita sur courbes adéliques
+Arithmetic Fujita approximation over adelic curves
+Chunhui Liu Institute for Advanced Study in Mathematics
+Harbin Institute of Technology
+150001 Harbin
+P. R. China
+chunhui.liu@hit.edu.cn
+(8 janvier 2026)
+Résumé
 In this paper, we will prove an analogue of Fujita’s approximation under the framework of Arakelov theory over adelic curves, which gives a positive response to a conjecture of Huayi Chen and Atsushi Moriwaki.
@@
 We say that $L$ is a big line bundle if $\operatorname{vol}(L)>0$.
-Fujita’s approximation theorem states that, for all $\epsilon>0$, there exist a $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(L^{\otimes p})$ can be decomposed as the tensor product of an ample line bundle $A$ and an effective line bundle $M$, which satisfy
+Fujita’s approximation theorem states that, for all $\epsilon>0$, there exist a $p\in\mathbb{N}^{+}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(L^{\otimes p})$ can be decomposed as the tensor product of an ample line bundle $A$ and an effective line bundle $M$, which satisfy
 $p^{-\dim(X)}\operatorname{vol}(A)\geqslant\operatorname{vol}(L)-\epsilon.$
@@
 This approximated decomposition theorem is proved by T.
-Fujita [12] for the case of characteristic 00, and by S.
+Fujita [12] for the case of characteristic 0, and by S.
 Takagi [24] for an arbitrary characteristic.
@@
 Let $K$ be a number field, $\mathscr{X}$ be an arithmetic variety over $\operatorname{Spec}\mathcal{O}_{K}$ of dimension $d$, and $\overline{\mathscr{L}}$ be a Hermitian line bundle on $\mathscr{X}$.
-We define
+For a $D\in\mathbb{N}$, we define
 $\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes D})=\log\#\left\{s\in H^{0}(\mathscr{X},\mathscr{L}^{\otimes D})\mid\|s\|_{\sup,v}\leqslant 1,\;\forall\;v\in M_{K,\infty}\right\},$
 and
-$\widehat{\operatorname{vol}}(\overline{\mathscr{L}})=\limsup_{n\to\infty}\frac{\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes n})}{n^{\dim(\mathscr{X})}/\dim(\mathscr{X})!}.$
-By the works mentioned above, the arithmetic volume function $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ has many similar properties to those of the geometric setting $\operatorname{vol}(\raisebox{1.72218pt}{.})$, so it is expected that an arithmetically big line bundle $\overline{\mathscr{L}}$, which means $\widehat{\operatorname{vol}}(\overline{\mathscr{L}})>0$, also has the property of Fujita approximation.
-In fact, this used to be a conjecture of A.
-Moriwaki.
+$\widehat{\operatorname{vol}}(\overline{\mathscr{L}})=\limsup_{D\to\infty}\frac{\widehat{h}^{0}(\mathscr{X},\overline{\mathscr{L}}^{\otimes D})}{D^{\dim(\mathscr{X})}/\dim(\mathscr{X})!}.$
+[…]
 1.1.2
@@
 We require the set $\Omega$ parameterized by $\phi$ has a $\sigma$-algebra structure $\mathcal{A}$ and a measure $\nu$ compatible with $\mathcal{A}$.
-Under this setting, we can generalize many classic results and subjects of Arakelov geometry, including the arithmetic volume function, arithmetic positivity, and Hilbert-Samuel formula.
+Under this setting, we are able to generalize many classic results and subjects of Arakelov geometry, including the arithmetic volume function, arithmetic positivity, and Hilbert-Samuel formula.
 Since the adelic structure becomes more complicated than those of classic cases, the obstructions original from measurability, integrability, and non-discreteness make the developing process more technique.
-In fact, some classic properties are not valid any longer in this framework, though we have developed many in [9, 10, 11], and see also [23].
+In fact, some classic properties are not valid any longer in this framework, though many of these properties have been developed in [9, 10, 11], and see also [23].
 1.2.2
 Inspired by [4, 26], it is quite natural to ask whether the Fujita approximation is still valid over adelic curves.
-Over a general adelic curve, we are lack of the Minkowskian property, so we cannot apply the filtrated method by successive minina directly in [4] At the same time, some constructions do not work any longer without the help of integral models.
-In [11, §8.5], H.
-Chen and A.
-Moriwaki proved a similar result to the arithmetic Fujita approximation over adelic curves, which is called the relative Fujita approximation.
-More precisely, let $\overline{L}$ be an adelic line bundle, where $L$ is big on $X$.
-Then for any $t<\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$, there exists a sufficiently large $p\in\mathbb{N}$ and a birational morphism $\varphi:X^{\prime}\rightarrow X$, such that $(\varphi^{*}\overline{L})^{\otimes p}=\overline{M}\otimes\overline{N}$, where $\overline{N}$ is relatively ample and $\overline{M}$ is effective, and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})>pt$.
-Here $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ denote the asymptotic maximal and minimal slopes respectively.
+Over a general adelic curve, we are lack of the Minkowskian property, so we cannot apply the filtrated method by successive minina directly in [4].
+At the same time, some constructions and properties do not work any longer without the help of integral models or Minkowskian property, and see Example 4.1.3 for such an example.
+[…]
 Compared with this relative version above, the classic arithmetic Fujita approximation seems to take an "average" at every minima of minimal slope, and approximates all of them simultaneously, not only the maximal one.
@@
 More generally, we prove a version of adelic $\mathbb{R}$-line bundle below at the same time, which is also a generalization of [4, Theorem 4.3] and [26, Theorem C].
+In the statement of the main result, instead of considering an effective line bundle under the arithmetic setting, we use the notation that a line bundle has a non-zero small section over an adelic curve, and see §3.1.5 for the precise definition.
 […]
@@
 Let $X$ be a projective variety over $\operatorname{Spec}K$, and $\overline{L}$ be an arithmetically big $\mathbb{R}$-line bundle on $X$ over $S$.
-Then for any $\epsilon>0$, there exist a sufficiently large $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(\overline{L}^{\otimes p})=\overline{A}\otimes\overline{M}$, where $\overline{A}$ is arithmetically ample and $\overline{M}$ is an arithmetically effective $\mathbb{R}$-line bundle.
+Then for any $\epsilon>0$, there exist a sufficiently large $p\in\mathbb{N}$ and a birational morphism $v:X^{\prime}\rightarrow X$, such that $v^{*}(\overline{L}^{\otimes p})=\overline{A}\otimes\overline{M}$, where $\overline{A}$ is arithmetically ample and $\overline{M}$ is an adelic $\mathbb{R}$-line bundle have a non-zero small section.
 Moreover, we have
@@
 If either $\mathcal{A}$ is discrete or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$, there will be no obstructions on the measurability or integrability, and see [9, §4] for more details of the necessity.
-In order to study some arithmetic positivities by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$, we need to suppose the base field $K$ is perfect.
-For example, for an adelic line bundle $\overline{L}$ where $L$ is nef, we need to apply $\widehat{\mu}_{\min}^{\operatorname{asy}}(\overline{L})\geqslant 0$ to deduce $\overline{L}$ is arithmetically nef by [11, Proposition 9.1.7], where we need the perfect condition.
-See [11, §9] for more details of the necessity.
-Of course, in [4, 26], there is no such problem since for the case of number field, because $\mathcal{A}$ is discrete in the language of adelic curve and the base field is perfect by its characteristic 00.
+[…]
+Of course, in [4, 26], there is no such problem since for the case of number field, because $\mathcal{A}$ is discrete in the language of adelic curve and the base field is perfect by its characteristic 0.
 1.4 Organization of article
@@
 I would like to thank Prof.
-Atsushi Moriwaki for introducing this subject to me.
+Atsushi Moriwaki for introducing this subject to me, and suggesting me to use the notation "having a non-zero small section".
 In addition, I would like to thank Prof.
@@
 Wenbin Luo for the advices on some key techniques.
+In particular, I would like to thank the anonymous referee for his or her valuable comments on this article.
 Part of this work was accomplished during my visit to Institute for Theoretical Sciences, Westlake University, and I would like to thank their generous invitation and reception.
@@
 […]
-Let $K$ be an arbitrary field, and |.$|\raisebox{1.72218pt}{.}|_{0}$ be the trivial absolute value over $K$, which means
-$|a|_{0}=\begin{cases}1,\text{ if }a\neq 0,\\ 0,\text{ if }a=0.\end{cases}$
-We note $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ the adelic curve which counts the absolute value |.$|\raisebox{1.72218pt}{.}|_{0}$ once. By definition, $S_{0}$ is a proper adelic curve.
-Example 2.1.1 will be used in the article later. We refer the readers to [9, §3.2] for more examples of adelic curve.
+Let $K$ be an arbitrary field, and $|\raisebox{1.72218pt}{.}|_{0}$ be the trivial absolute value over $K$, which means
+[…]
+We note $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ the adelic curve which counts the absolute value $|\raisebox{1.72218pt}{.}|_{0}$ once. By definition, $S_{0}$ is a proper adelic curve.
+Example 2.1.1 will be used in the article later. We refer the readers to [9, §3.2] for more examples of adelic curves.
 2.1.2 Adelic vector bundle
 Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, and $E$ be a vector space over $K$.
-For a family of norms (.$(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ on $E_{w}=E\otimes_{K}K_{w}$ over $K_{w}$, if the norm family (.$(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ on $E$ and its dual norm family (.$(\|\raisebox{1.72218pt}{.}\|^{\vee}_{w})_{w\in\Omega}$ on $E^{\vee}$ are both $\mathcal{A}$-measurable on $\Omega$, then we say $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ is an adelic vector bundle over $S$.
+For a family of norms $(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ on $E_{w}=E\otimes_{K}K_{w}$ over $K_{w}$, if the norm family $(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ on $E$ and its dual norm family $(\|\raisebox{1.72218pt}{.}\|^{\vee}_{w})_{w\in\Omega}$ on $E^{\vee}$ are both $\mathcal{A}$-measurable on $\Omega$, and $(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega}$ is dominated, then we say $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ is an adelic vector bundle over $S$.
 If $\dim(E)=1$, we say that $\overline{E}$ is an adelic line bundle over $S$.
-We say that an adelic vector bundle $\overline{F}=(F,(\|\raisebox{1.72218pt}{.}\|^{\prime}_{w})_{w\in\Omega})$ is an adelic sub-bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $F$ is a $K$-vector subspace of $E$, and .$\|\raisebox{1.72218pt}{.}\|^{\prime}_{w}$ is the restriction of .$\|\raisebox{1.72218pt}{.}\|_{w}$ on $F$ for each $w\in\Omega$.
-An adelic bundle $\overline{G}=(G,(\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w})_{w\in\Omega})$ is said to be an adelic quotient bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $G$ is a quotient vector space of $E$, and .$\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w}$ is the quotient norm .$\|\raisebox{1.72218pt}{.}\|_{w}$ with respect to the canonical quotient homomorphism $E\rightarrow G$ for each $w\in\Omega$.
+See [9, Definition 4.1.28] for more details on this notion.
+We say that an adelic vector bundle $\overline{F}=(F,(\|\raisebox{1.72218pt}{.}\|^{\prime}_{w})_{w\in\Omega})$ is an adelic sub-bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $F$ is a $K$-vector subspace of $E$, and $\|\raisebox{1.72218pt}{.}\|^{\prime}_{w}$ is the restriction of $\|\raisebox{1.72218pt}{.}\|_{w}$ on $F$ for each $w\in\Omega$.
+An adelic bundle $\overline{G}=(G,(\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w})_{w\in\Omega})$ is said to be an adelic quotient bundle of $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, if $G$ is a quotient vector space of $E$, and $\|\raisebox{1.72218pt}{.}\|^{\prime\prime}_{w}$ is the quotient norm $\|\raisebox{1.72218pt}{.}\|_{w}$ with respect to the canonical quotient homomorphism $E\rightarrow G$ for each $w\in\Omega$.
 If $G$ is the quotient space of $E$ modulo $F$, and is equipped with the quotient metrics from $\overline{F}$, then we denote $\overline{G}=\overline{E}/\overline{F}=\overline{E/F}$.
-For an adelic vector bundle $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, we define $\det(\overline{E})=(\det(E),(\|\raisebox{1.72218pt}{.}\|_{\det,w})_{w\in\Omega})$ as follows: the $K$-vector space $\det(E)=\bigwedge^{\dim(E)}(E)$ is the determinant space of $E$, and .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$ is the determinant norm on $\det(E)_{w}$ of .$\|\raisebox{1.72218pt}{.}\|_{w}$ for each $w\in\Omega$.
-We refer the readers to [9, §1.2.3] for the details of the construction of .$\|\raisebox{1.72218pt}{.}\|_{\det,w}$.
+For an adelic vector bundle $\overline{E}=(E,(\|\raisebox{1.72218pt}{.}\|_{w})_{w\in\Omega})$ over $S$, we define $\det(\overline{E})=(\det(E),(\|\raisebox{1.72218pt}{.}\|_{\det,w})_{w\in\Omega})$ as follows : the $K$-vector space $\det(E)=\bigwedge^{\dim(E)}(E)$ is the determinant space of $E$, and $\|\raisebox{1.72218pt}{.}\|_{\det,w}$ is the determinant norm on $\det(E)_{w}$ of $\|\raisebox{1.72218pt}{.}\|_{w}$ for each $w\in\Omega$.
+We refer the readers to [9, §1.1.13, §1.2.3] for the details of the construction of $\|\raisebox{1.72218pt}{.}\|_{\det,w}$.
 […]
@@
 $\nu_{(W,\mathcal{F})}=\sum_{i=1}^{n}\frac{\dim\left(W_{i-1}/W_{i}\right)}{\dim(W)}\delta_{\lambda_{i}},$
-where $\delta_{\lambda_{i}}$ is the Dirac operator at the real number $\lambda_{i}$.
+where $\delta_{\lambda_{i}}$ is the Dirac measure at $\lambda_{i}\in\mathbb{R}$.
 Next, we define the function
@@
 This $\mathbb{R}$-filtration $\mathcal{F}_{\operatorname{HN}}$ is called the Harder-Narasimhan filtration of $\overline{E}$.
-The last equality in Definition 2.2.3 if from [9, Proposition 4.3.36]. In fact, the Harder-Narasimhan filtration of $\overline{E}$ corresponds to an increasing flag of vector subspaces
+The last equality in Definition 2.2.3 is from [9, Proposition 4.3.36]. In fact, the Harder-Narasimhan filtration of $\overline{E}$ corresponds to an increasing flag of vector subspaces
 $\{0\}=E_{1}\subsetneq E_{2}\subsetneq\cdots\subsetneq E_{n}=E$ (3)
@@
 2.3 Filtrations on graded algebras
-In this part, we study some asymptotic properties of graded linear algebras by $\mathbb{R}$-filtraion.
+In this part, we study some asymptotic properties of graded linear algebras by $\mathbb{R}$-filtration.
 2.3.1 Approximable graded algebras
@@
     $\liminf\limits_{n\to\infty}\frac{\dim_{K}\left(\operatorname{Im}\left(S^{n}E_{p}\rightarrow E_{np}\right)\right)}{\dim_{K}(E_{np})}>1-\epsilon,$
-    where $S^{n}B_{p}\rightarrow E_{np}$ is the canonical homomorphism defined by the algebra structure on $E_{\bullet}$.
+    where $S^{n}E_{p}\rightarrow E_{np}$ is the canonical homomorphism defined by the algebra structure on $E_{\bullet}$.
 2.3.2 Convergence of measures on an approximable algebra
@@
 $\lambda(x_{1}x_{2})\geqslant\lambda(x_{1})+\lambda(x_{2}),$
-where the function $\lambda(\raisebox{1.72218pt}{.})$ is defined at (1) and the homogeneous degrees of $x_{1}$ and $x_{2}$ are ignored, then we say $E_{\bullet}$ is filtrated with respect to $\mathcal{F}$.
+[…]
 For every $\epsilon>0$, let $T_{\epsilon}$ be the operator on the space of Borel probability measures on $\mathbb{R}$ by mapping $x\mapsto\epsilon x$.
@@
 […]
-With all the above notations.
-Suppose that $E_{\bullet}=\bigoplus\limits_{D\geqslant 0}E_{D}$ is a filtrated approximable graded algebra with respect to $\lambda(\raisebox{1.72218pt}{.})$ defined at (1), and
+[…]
 $\sup_{n\geqslant 1}\frac{\lambda_{\max}(E_{n})}{n}<+\infty,$
-where $\lambda_{\max}(\raisebox{1.72218pt}{.})$ is defined at (2).
-Then the sequence $\left(\frac{\lambda_{\max}(E_{n})}{n}\right)_{n\in\mathbb{N}^{+}}$ converges in $\mathbb{R}$, and the sequence of measures $\left(T_{\frac{1}{n}}(\nu_{E_{n}})\right)_{n\geqslant 1}$ converges weakly to a Borel probability measure on $\mathbb{R}$.
+[…]
 2.3.3 Limits of a measure from filtrated algebras
@@
 $\lambda_{+}^{\operatorname{asy}}(E_{\bullet}):=\int_{0}^{+\infty}x\nu_{E_{\bullet}}(\mathrm{d}x)=\lim\limits_{n\to\infty}\frac{\lambda_{+}(E_{n})}{n},$ (4)
-where $\lambda_{+}(\raisebox{1.72218pt}{.})$ is defined at (2). See [4, §2.4] for more details.
+[…]
 We have the following result.
@@
 Then we construct a family of norms on $H^{0}(X,L)$ in the way below : for every $w\in\Omega$ and $x\in X_{w}^{\operatorname{an}}$, define
-|.$\begin{array}[]{rrcl}|\raisebox{1.72218pt}{.}|_{\varphi_{w}}(x):&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&|s|_{\varphi_{w}}(x)\end{array}$
+$\begin{array}[]{rrcl}|\raisebox{1.72218pt}{.}|_{\varphi_{w}}(x):&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&|s|_{\varphi_{w}}(x)\end{array}$
 as the value of the global section $s$ at the point $x$ with respect to the given metric, and
-.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\varphi_{w}}:&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&\sup\limits_{x\in X^{\operatorname{an}}_{w}}\{|s|_{\varphi_{w}}(x)\}.\end{array}$ (5)
+$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\varphi_{w}}:&H^{0}(X_{w},L_{w})&\longrightarrow&\mathbb{R}\\ &s&\mapsto&\sup\limits_{x\in X^{\operatorname{an}}_{w}}\{|s|_{\varphi_{w}}(x)\}.\end{array}$ (5)
 We denote $\varphi=(\|\raisebox{1.72218pt}{.}\|_{\varphi_{w}})_{w\in\Omega}$ the norm family on $H^{0}(X,L)$ over $S$ if there is no ambiguity.
@@
 $\lim_{n\to\infty}\frac{1}{mn}d(nm\varphi,\varphi_{n})=0,$
-we say that the metric family $\varphi$ is semi-positive.
+we say that the metric family $\varphi$ is semi-positive. See [9, Definition 2.2.21] for the details of the construction of a quotient norm family.
 3.1.3 Ampleness
@@
 Let $\overline{L}=(L,\varphi)$ be an adelic line bundle on $X$ over $S$.
-If $s\in H^{0}(X,L)$ is a non-zero global section which satisfies $\|s\|_{\varphi_{w}}\leqslant 1$ for all $w\in\Omega$, we say that the global section $s$ is effective.
-We say that an adelic line bundle $\overline{L}$ is effective or arithmetically effective, if it admits at least one effective global section by the above definition.
+If $s\in H^{0}(X,L)$ is a non-zero global section which satisfies $\|s\|_{\varphi_{w}}\leqslant 1$ for all $w\in\Omega$, we say that the global section $s$ is small, or $s$ is a small section.
+We say that an adelic line bundle $\overline{L}$ has a non-zero small section, if it admits at least one small global section by the above definition.
+[…]
+In [11, Definition 8.5.4], the above $\overline{L}$ is said to be effective in the arithmetic sense, but this notion is only used in §8.5 and §8.6 of [11].
+In order to avoid the confusion, we use the word $\overline{L}$ has a non-zero small section” instead of [11, Definition 8.5.4] in this article, because some multiple of an arithmetically big adelic line bundle is not necessarily effective.
+This is different from the classic case over a number field.
 3.2 Algebraic construction of graded linear series
@@
 In §3.1.1, we defined a norm family on $H^{0}(X,L^{\otimes D})$ by (5) indexed by $\Omega$, which is denoted by $D\varphi$ with no ambiguity.
-By [9, §6.4.1], .$\|\raisebox{1.72218pt}{.}\|_{D\varphi_{w}}$ is ultrametric for every $w\in\Omega_{f}$.
+By [9, §6.4.1], $\|\raisebox{1.72218pt}{.}\|_{D\varphi_{w}}$ is ultrametric for every $w\in\Omega_{f}$.
 Then $(H^{0}(X,L^{\otimes D}),D\varphi)$ is an adelic vector bundle over $S$.
-For every $D\in\mathbb{N}$, let $\overline{E}_{D}$ be an adelic vector sub-bundle $H^{0}(X,L^{\otimes D})$ over $S$.
+For every $D\in\mathbb{N}$, let $\overline{E}_{D}$ be an adelic vector sub-bundle of $H^{0}(X,L^{\otimes D})$ over $S$.
 In this case, we say that $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ is a normed graded linear series over $S$ if $E_{D}\cdot E_{D^{\prime}}\subseteq E_{D+D^{\prime}}$ for $D,D^{\prime}\in\mathbb{N}$.
@@
 $\widehat{\operatorname{vol}}(\overline{E}_{\bullet})=\limsup_{n\to\infty}\frac{\widehat{\deg}_{+}\left(\overline{E}_{n}\right)}{n^{\dim(X)+1}/(\dim(X)+1)!}.$ (6)
-In particular, if $\overline{E}_{D}=\left(H^{0}(X,L^{\otimes D}),D\varphi\right)$ as above, we denote $\widehat{\operatorname{vol}}(\overline{L})=\widehat{\operatorname{vol}}(\overline{E}_{\bullet})$ for simplicity.
+In particular, if $\overline{E}_{n}=\left(H^{0}(X,L^{\otimes n}),n\varphi\right)$ as above, we denote $\widehat{\operatorname{vol}}(\overline{L})=\widehat{\operatorname{vol}}(\overline{E}_{\bullet})$ for simplicity.
 If $\widehat{\operatorname{vol}}(\overline{L})>0$, we say that $\overline{L}$ is big on $X$, or arithmetically big to distinguish it from the geometric case.
@@
 Let $\overline{E}$ be an adelic vector bundle over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$, and $\left(\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E})\right)_{t\in\mathbb{R}}$ be the Harder-Narasimhan filtration of $\overline{E}$ defined in Definition 2.2.3.
-We define the norm .$\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}$ on $E$ over $(K,|\raisebox{1.72218pt}{.}|_{0})$ as
-.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}:&E&\longrightarrow&\mathbb{R}^{+}\\ &s&\mapsto&\exp\left(-\sup\{t\in\mathbb{R}\mid s\in\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E})\}\right).\end{array}$ (7)
+We define the norm $\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}$ on $E$ over $(K,|\raisebox{1.72218pt}{.}|_{0})$ as
+$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}:&E&\longrightarrow&\mathbb{R}^{+}\\ &s&\mapsto&\exp\left(-\sup\{t\in\mathbb{R}\mid s\in\mathcal{F}^{t}_{\operatorname{HN}}(\overline{E})\}\right).\end{array}$ (7)
 Then $\overline{E}_{\operatorname{HN}}=\left(E,\|\raisebox{1.72218pt}{.}\|_{\operatorname{HN}}\right)$ is an adelic vector bundle over $S_{0}$.
@@
 Let $\overline{E}_{\bullet}=\bigoplus\limits_{D\in\mathbb{N}}\overline{E}_{D}$ be a normed graded linear series over an adelic curve $S=(K,(\Omega,\mathcal{A},\nu),\phi)$.
-For every $\overline{E}_{D}$, we denote by .$\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}}$ the norm (7) on $E_{D}$ indexed by $D\in\mathbb{N}$.
+For every $\overline{E}_{D}$, we denote by $\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}}$ the norm (7) on $E_{D}$ indexed by $D\in\mathbb{N}$.
 Then $(E_{D},\|\raisebox{1.72218pt}{.}\|_{D}^{\operatorname{HN}})$ an adelic vector bundle over $S_{0}$.
 By [11, Proposition 4.1.3], for every $s\in E_{m}$ and $t\in E_{n}$, we have
-$\|s\cdot t\|^{\operatorname{HN}}_{m+n}\leqslant\exp(f(m)+f(n))\|s\|^{\operatorname{HN}}_{m}\|t\|^{\operatorname{HN}}_{n},$
+$\|s\cdot t\|^{\operatorname{HN}}_{m+n}\leqslant\exp\left(f(\dim E_{m})+f(\dim E_{n})\right)\|s\|^{\operatorname{HN}}_{m}\|t\|^{\operatorname{HN}}_{n},$
 where $f(n)=\frac{3}{2}\nu(\Omega_{\infty})\log n$, and $\nu(\Omega_{\infty})$ is the measure of archimedean absolute values of $S=(K,(\Omega,\mathcal{A},\nu),\phi)$.
 The term $f(n)$ is original from the partial response of the semi-stable conjecture of Bost [1, Theorem A], see also [8, §5.6].
-In order to obtain a new family of norms which are sub-multiplicative, we introduce the spectral norm .$\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ of .$\|\raisebox{1.72218pt}{.}\|^{\operatorname{HN}}_{D}$ as
-.$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}:&H^{0}(X,L^{\otimes D})&\longrightarrow&\mathbb{R}_{\geqslant 0}\\ &s&\mapsto&\lim\limits_{n\rightarrow\infty}\left(\|s^{n}\|_{nD}^{\operatorname{HN}}\right)^{1/n}.\end{array}$ (8)
+In order to obtain a new family of norms which are sub-multiplicative, we introduce the spectral norm $\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ of $\|\raisebox{1.72218pt}{.}\|^{\operatorname{HN}}_{D}$ as
+$\begin{array}[]{rrcl}\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}:&H^{0}(X,L^{\otimes D})&\longrightarrow&\mathbb{R}_{\geqslant 0}\\ &s&\mapsto&\lim\limits_{n\rightarrow\infty}\left(\|s^{n}\|_{nD}^{\operatorname{HN}}\right)^{1/n}.\end{array}$ (8)
 In this case, $\left(E_{D},\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}\right)$ is an adelic vector bundle over $S_{0}$. By [11, Proposition 3.2.2 (3)], for every $s\in E_{m}$ and $t\in E_{n}$, we have
@@
 We denote by $\mathcal{F}_{\mathrm{sp}}$ the $\mathbb{R}$-filtration on $\overline{E}_{\operatorname{HN},D}$ induced by the filtrated structure $\lambda(\raisebox{1.72218pt}{.})=-\log\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$, where we omit the index $D\in\mathbb{N}$ in the above notation.
-The existence of such $\mathcal{F}_{\mathrm{sp}}$ is ensured by the finiteness of the values of .$\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ on $\overline{E}_{\operatorname{HN},D}$.
+The existence of such $\mathcal{F}_{\mathrm{sp}}$ is ensured by the finiteness of the values of $\|\raisebox{1.72218pt}{.}\|_{\mathrm{sp},D}$ on $\overline{E}_{\operatorname{HN},D}$.
 3.4.4
@@
 $\widehat{\operatorname{vol}}(\overline{E}_{\bullet})=\widehat{\operatorname{vol}}(\overline{E}_{\operatorname{HN},\bullet}).$
-Proof.
+Démonstration.
 It is from [11, Corollary 3.4.4, (4.1.2), Corollary-Definition 4.1.4] directly. ∎
 By Proposition 2.3.2, the sequence $\left(\frac{1}{n}\widehat{\mu}_{\max}(\overline{E}_{n})\right)_{n\geqslant 0}$ converges to a real number, and we denote it by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\bullet})$.
-If $E_{n}=H^{0}(X,L^{\otimes n})$ equipped with the superior norms, it is denoted by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$ for simplicity.
+If $E_{n}=H^{0}(X,L^{\otimes n})$ equipped with the supremum norms, it is denoted by $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})$ for simplicity.
 The following result is an equivalent property of arithmetically bigness.
@@
 $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})=\lambda_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}).$
-Proof.
+Démonstration.
 The equivalence of $\widehat{\operatorname{vol}}(\overline{L})>0$ and $\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{L})>0$ is from [9, Proposition 6.4.18] directly.
@@
 $\lambda_{+}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet})=\frac{\widehat{\operatorname{vol}}(\overline{E}_{\bullet})}{\left(\dim(X)+1\right)\operatorname{vol}(E_{\bullet})},$
-where $\lambda_{+}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ is defined at (4) with respect to the canonical filtrated structure introduced in §3.4.3 over $S_{0}$.
+[…]
 3.4.5
@@
 where the adelic vector bundle $(H^{0}(X,L^{\otimes n}),n\varphi)$ over $S$ is defined in §3.1.1.
-We recall some results on the criteria of arithmetic positivities via $\widehat{\mu}_{\max}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ and $\widehat{\mu}_{\min}^{\operatorname{asy}}(\raisebox{1.72218pt}{.})$ in [11, §9] below, where we always suppose that $K$ is perfect.
+[…]
 […]
@@
 which is a graded linear series over $S_{0}$.
-4.1.2
 By the properties of $\overline{E}_{\operatorname{HN},\bullet}$ in §3.4.3, we have the following proposition.
@@
 The graded linear series (10) is a filtrated graded linear series over $S_{0}$.
-4.1.3
+4.1.2
 Next, we prove that $\overline{E}_{\operatorname{HN},D}^{\lambda}$ is approximable when $\lambda$ is of some particular values.
@@
 […]
-Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be an adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
+Let $S=(K,(\Omega,\mathcal{A},\nu),\phi)$ be a proper adelic curve, where either $\mathcal{A}$ is a discrete $\sigma$-algebra or $K$ has a countable sub-field which is dense in $K_{w}$ for all $w\in\Omega$.
 Let $\pi:X\rightarrow\operatorname{Spec}K$ be a projective morphism, and $\overline{L}$ be an adelic line bundle over $S$.
@@
 If $\overline{L}$ is big, then for all $\lambda\in\left(-\infty,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$, $\overline{E}^{\lambda}_{\operatorname{HN},\bullet}$ contains an ample divisor.
-Proof.
+Démonstration.
 By Proposition 3.4.4 and Proposition 3.4.4, from the fact that $\overline{L}$ is big, we have $\overline{E}_{\operatorname{HN},\bullet}$ is big.
@@
 factors through $\bigoplus\limits_{n\geqslant 0}E_{\operatorname{HN},np}^{\lambda}$. ∎
-[…]
-In [4, Proposition 3.11], a similar result of Proposition 4.1.3 is proved by an arithmetic analogue of Kodaira decomposition [25, Corollary 2.4].
-Until the author’s knowledge, we do not know whether it is still valid over a general adelic curve.
-To avoid this obstruction, we use a method quite similar to that of [2], and see also [9, §6.3].
 Same as the argument in [4, Corollary 3.13], we have the following result.
@@
 The graded linear series (10) over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ is approximable for any $\lambda\in\left(-\infty,\widehat{\mu}^{\operatorname{asy}}_{\max}(\overline{L})\right)$.
-Proof.
+Démonstration.
 By Proposition 3.2.1, we have the assertion from Proposition 4.1.2 directly. ∎
+4.1.3
+In [4, Proposition 3.11], a similar result of Proposition 4.1.2 is proved by an arithmetic analogue of Kodaira decomposition [25, Corollary 2.4] of X.
+Yuan.
+But it is not valid over a general adelic curve, and see Example 4.1.3 below for a counterexample.
+To avoid this obstruction, we use a method quite similar to that of [2] to prove Proposition 4.1.2, and see also [9, §6.3].
+[…]
+Let $K$ be a field, and $S_{3}=(K,\{1,2,3\},|\raisebox{1.72218pt}{.}|_{0})$, where the trivially valued absolute $|\raisebox{1.72218pt}{.}|_{0}$ (see Example 2.1.1 for the precise definition) is counted three times in a discrete $\sigma$-algebra.
+We consider the universal line bundle $\mathcal{O}(1)$ on $\mathbb{P}_{K}^{1}$, and we denote by $x_{0},x_{1}$ the basis of $H^{0}(\mathbb{P}^{1}_{K},\mathcal{O}(1))$ with respect to the homogeneous coordinate.
+Let $\overline{\mathcal{O}(1)}=(\mathcal{O}(1),(\|\raisebox{1.72218pt}{.}\|_{i})_{i=1}^{3})$ be the adelic line bundle on $\mathbb{P}^{1}_{K}$, where the induced supremum norms $\|\raisebox{1.72218pt}{.}\|_{i=1}^{3}$ on $H^{0}(\mathbb{P}^{1}_{K},\mathcal{O}(1))$ satisfies
+$\begin{cases}\|x_{i}\|_{1}=e^{-\frac{1}{2},},\text{ where }i=0,1;\\ \|x_{i}\|_{2}=e^{-\frac{1}{2},},\text{ where }i=0,1;\\ \|x_{i}\|_{3}=e^{\frac{3}{4},},\text{ where }i=0,1.\end{cases}$
+Then $\overline{\mathcal{O}(1)}$ is an arithmetically big line bundle on $\mathbb{P}^{1}_{K}$ over $S_{3}$, since
+$\widehat{\deg}\left(H^{0}(\mathbb{P}^{1}_{K},\mathcal{O}(D)),(\|\raisebox{1.72218pt}{.}\|_{i})_{i=1}^{3}\right)=\frac{1}{4}D(D+1)$
+for all $D\in\mathbb{N}^{+}$.
+If there exist a $D\in\mathbb{N}$, an adelic ample line bundle $\overline{A}$ and an line bundle $\overline{M}$ have a non-zero small section on $\mathbb{P}^{1}_{K}$, such that $\overline{\mathcal{O}(D)}=\overline{A}\otimes\overline{M}$, where $\overline{\mathcal{O}(D)}$ is equipped with the induced metrics from the one introduced above.
+Since $\operatorname{Pic}(\mathbb{P}^{1}_{K})=\mathbb{Z}$, then there exists a $D^{\prime}\in\mathbb{N}$ satisfying $D^{\prime}<D+$, such that
+$\overline{\mathcal{O}(D)}=\overline{\mathcal{O}(D^{\prime})}\otimes\overline{\mathcal{O}(D-D^{\prime})},$
+where $\overline{\mathcal{O}(D^{\prime})}$ is ample and $\overline{\mathcal{O}(D-D^{\prime})}$ has a non-zero small section.
+By the definition of arithmetically ample line bundle, all the global sections of $\overline{\mathcal{O}(D^{\prime})}$ should be of the norms smaller than $1$.
+But by the choice of $\|\raisebox{1.72218pt}{.}\|_{3}$ in $H^{0}(\mathbb{P}^{1}_{K},\mathcal{O}(1))$ over $|\raisebox{1.72218pt}{.}|_{0}$, all the global sections of $\overline{\mathcal{O}(D-D^{\prime})}$ should be equipped with a norm strictly larger than $1$ with respect to the third $\|\raisebox{1.72218pt}{.}\|_{0}$ in $S_{3}$.
+So there will be no small section in $\overline{\mathcal{O}(D-D^{\prime})}$, which shows there is no arithmetic Kodaira decomposition in this case.
 4.2 Fujita approximation of graded linear series
-Let $p\in\mathbb{N}$ such that $E_{\operatorname{HN},p}^{0}\neq 0$, where we follow the notation in (9).
-In this case, we define $\overline{E}_{\bullet}^{(p)}=\bigoplus\limits_{n\geqslant 0}\overline{E}^{(p)}_{\operatorname{HN},np}$ to be the graded $K$-algebra generated by $E_{\operatorname{HN},p}^{0}$, which is considered as a normed graded linear series over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ defined in Example 2.1.1.
+Let $p\in\mathbb{N}$ such that $\overline{E}_{\operatorname{HN},p}^{0}\neq 0$, where we follow the notation in (9).
+In this case, we define $\overline{E}_{\bullet}^{(p)}=\bigoplus\limits_{n\geqslant 0}\overline{E}^{(p)}_{\operatorname{HN},np}$ to be the graded $K$-algebra generated by $\overline{E}_{\operatorname{HN},p}^{0}$, which is considered as a normed graded linear series over $S_{0}=(K,\{0\},|\raisebox{1.72218pt}{.}|_{0})$ defined in Example 2.1.1.
 The following result is an analogue of [4, Theorem 4.1] over $S_{0}$, whose proof is quite similar to the ancient one.
@@
 $\widehat{\operatorname{vol}}(\overline{L})=\sup_{p\in\mathbb{N}}\widehat{\operatorname{vol}}(\overline{E}_{\bullet}^{(p)}).$
-Proof.
+Démonstration.
 For any $n\in\mathbb{N}^{+}$, let $\nu_{\mathrm{sp},n}=T_{\frac{1}{n}}\nu_{\left(E_{\operatorname{HN},n},\mathcal{F}_{\mathrm{sp}}\right)}$, where the notation $T_{\frac{1}{n}}\nu_{\left(E_{\operatorname{HN},n},\mathcal{F}_{\mathrm{sp}}\right)}$ is introduced in §2.3.2, and the $\mathbb{R}$-filtration $\mathcal{F}_{\mathrm{sp}}$ is defined in §3.4.2 with respect to the spectral norms.
-Then by Proposition 2.3.2, the sequence $(\nu_{\mathrm{sp},n})_{n\geqslant 1}$ converges weakly to a Borel probability measure, which is denoted by $\nu_{\mathrm{sp}}$.
+Then by Proposition 2.3.2 and Corollary 4.1.2, the sequence $(\nu_{\mathrm{sp},n})_{n\geqslant 1}$ converges weakly to a Borel probability measure, which is denoted by $\nu_{\mathrm{sp}}$.
 Similarly, let $\nu_{\mathrm{sp},n}^{(p)}=T_{\frac{1}{np}}\nu_{\left(E^{(p)}_{\operatorname{HN},np},\mathcal{F}_{\mathrm{sp}}\right)}$, which also converges weakly to a Borel probability measure noted by $\nu_{\mathrm{sp}}^{(p)}$ by Proposition 4.1.1.
-In order to prove the assertion, we will prove that the restriction of $\nu_{\mathrm{sp}}$ on $[0,+\infty)$ can be approximated well by $\nu_{\mathrm{sp}}^{(p)}$ for $p\in\mathbb{N}$.
+In order to prove the assertion, we will prove that the restriction of $\nu_{\mathrm{sp}}$ on $0,+\infty)$ can be approximated well by $\nu_{\mathrm{sp}}^{(p)}$ for $p\in\mathbb{N}$.
 Step I. – Simultaneous approximations.
 Let
-[…]
-be a division of the interval $[0,\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}))$ such that
+$D:\;0=t_{0}<t_{1}<\cdots<t_{m}<\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet})$
+be a division of the interval $0,\widehat{\mu}_{\max}^{\operatorname{asy}}(\overline{E}_{\operatorname{HN},\bullet}))$ such that
 $\nu_{\mathrm{sp}}(\{t_{1},\ldots,t_{m}\})=0.$ (11)
@@
   1. (1)
-    $\overline{M}$ is effective and $\overline{G}$ is arithmetically nef;
+    $\overline{M}$ has a non-zero small section and $\overline{G}$ is arithmetically nef ;
   2. (2)
     $p^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon$.
-Proof.
+Démonstration.
 We devide the proof into three steps.
@@
 We consider $\overline{E}^{(p)}_{\bullet}$ as a normed graded linear series over $S$, which is equipped with the subspace norms of $\overline{E}_{\bullet}=\bigoplus\limits_{D\geqslant 0}\overline{E}_{D}$.
-For any $n\geqslant 1$ and $w\in\Omega$, let .$\|\raisebox{1.72218pt}{.}\|_{n,w}$ be the quotient norm on $G_{p,w}=G_{p}\times_{\operatorname{Spec}K}\operatorname{Spec}K_{w}$ induced by the surjective homomorphism
+For any $n\geqslant 1$ and $w\in\Omega$, let $\|\raisebox{1.72218pt}{.}\|_{n,w}$ be the quotient norm on $G_{p,w}=G_{p}\times_{\operatorname{Spec}K}\operatorname{Spec}K_{w}$ induced by the surjective homomorphism
 $\phi_{p}^{*}\pi^{*}(E_{\operatorname{HN},np}^{(p)})\rightarrow G_{p}^{\otimes n},$
-where $E_{\operatorname{HN},np}^{(p)}$ is equipped with the superior norm .$\|\raisebox{1.72218pt}{.}\|_{\sup,w}$ with respect to $w\in\Omega$.
+where $E_{\operatorname{HN},np}^{(p)}$ is equipped with the superior norm $\|\raisebox{1.72218pt}{.}\|_{\sup,w}$ with respect to $w\in\Omega$.
 In fact, for any $x\in X_{p,w}(\overline{K}_{w})$ outside the exceptional divisor, it corresponds to a $1$-dimensional quotient of $E_{\operatorname{HN},p}^{0}$.
@@
 In other words, there exists a section $w\in E_{\operatorname{HN},np}^{(p)}$, whose image in $G_{p,w}^{\otimes n}(x)$ has norm $\|w\|_{\sup,w}$.
-As a section of $L_{w}^{\otimes np}$ on $X_{v}(\overline{K}_{w})$, we have $\|w(x)\|_{w}\leqslant\|w\|_{\sup,w}$, where .$\|\raisebox{1.72218pt}{.}\|_{w}$ is the metric on $\overline{L}$ indexed by $v\in\Omega$.
+As a section of $L_{w}^{\otimes np}$ on $X_{v}(\overline{K}_{w})$, we have $\|w(x)\|_{w}\leqslant\|w\|_{\sup,w}$, where $\|\raisebox{1.72218pt}{.}\|_{w}$ is the metric on $\overline{L}$ indexed by $v\in\Omega$.
 Then for any section $u\in G_{p,w}$ over a neighborhood of $x$, we may assume that $u(x)^{\otimes n}$ equals to the image of $w(x)$ in $G_{p,w}^{\otimes n}(x)$ by dilation, then we have
@@
 Let $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ be a family of adelic metrics on $G_{p}$, and we define $(M_{p},\beta_{n})=\phi_{p}^{*}(\overline{L})\otimes(G_{p},\alpha_{n})^{\vee}$.
-In this sense, $s$ is an effective section in $(M_{p},\beta_{n})$, then it is an effective adelic line bundle.
+In this sense, $s$ is an small section in $(M_{p},\beta_{n})$, then it is an adelic line bundle have a non-zero small section.
 Step III. – The relation among arithmetic volumes.
@@
 $\widehat{\operatorname{vol}}(E_{\bullet}^{(p)},\alpha_{n})=\lim_{m\rightarrow\infty}\frac{\widehat{\deg}_{+}(E_{\operatorname{HN},mp}^{(p)},(\|\raisebox{1.72218pt}{.}\|_{n,\sup,w})_{w\in\Omega})}{m^{d+1}/(d+1)!},$
-and .$\|\raisebox{1.72218pt}{.}\|_{n,\sup,w}$ is the superior norm induced by $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ on $G_{p}$. Then by [4, Corollary A.2], we have
+and $\|\raisebox{1.72218pt}{.}\|_{n,\sup,w}$ is the superior norm induced by $\alpha_{n}=(\|\raisebox{1.72218pt}{.}\|_{n,w})_{w\in\Omega}$ on $G_{p}$. Then by [4, Corollary A.2], we have
 $\limsup_{n\to\infty}\widehat{\operatorname{vol}}(E_{\bullet}^{(p)},\alpha_{n})\geqslant\widehat{\operatorname{vol}}(\overline{E}_{\bullet}^{(p)}).$
@@
 We keep all the notations and conditions in Theorem 5.1. Then we are able to require $\overline{G}$ to be arithmetically ample in the term (1) of Theorem 5.1.
-Proof.
+Démonstration.
 Let $\overline{B}$ be an arithmetically big line bundle on $X$.
@@
 $\phi^{*}(\overline{L}^{\otimes mp}\otimes\overline{B}^{\otimes(-p)})=\overline{G}\otimes\overline{M}$
-with $\overline{G}$ arithmetically nef and $\overline{M}$ effective, such that
+with $\overline{G}$ arithmetically nef and $\overline{M}$ having a non-zero small section, such that
 $p^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G})\geqslant\widehat{\operatorname{vol}}(\overline{L}^{\otimes m}\otimes\overline{B}^{\vee})-\epsilon.$
 Let $\overline{B}^{\prime}$ be an arithmetically ample line bundle on $X^{\prime}$.
-Since $\phi^{*}(\overline{B})$ is arithmetically big, there exists a $q\in\mathbb{N}$, such that $\phi^{*}(\overline{B}^{\otimes q})\otimes\overline{B}^{\prime\vee}$ is effective.
+Since $\phi^{*}(\overline{B})$ is arithmetically big, there exists a $q\in\mathbb{N}$, such that $\phi^{*}(\overline{B}^{\otimes q})\otimes\overline{B}^{\prime\vee}$ has a non-zero small section.
 We consider
 $\phi^{*}(\overline{L}^{\otimes mpq})=(\overline{G}^{\otimes q}\otimes\overline{B}^{\prime})\otimes(\overline{M}^{\otimes q}\otimes\phi^{*}(\overline{B}^{\otimes pq})\otimes\overline{B}^{\prime\vee}).$
-In the above decomposition, $\overline{G}^{\prime}:=\overline{G}^{\otimes q}\otimes\overline{B}^{\prime}$ is arithmetically ample, $\overline{M}^{\otimes q}\otimes\phi^{*}(\overline{B}^{\otimes pq})\otimes\overline{B}^{\prime\vee}$ is effective, and
+In the above decomposition, $\overline{G}^{\prime}:=\overline{G}^{\otimes q}\otimes\overline{B}^{\prime}$ is arithmetically ample, $\overline{M}^{\otimes q}\otimes\phi^{*}(\overline{B}^{\otimes pq})\otimes\overline{B}^{\prime\vee}$ has a non-zero small section, and
 $(mpq)^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G}^{\prime})\geqslant(mpq)^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G}^{\otimes q})\geqslant m^{-(d+1)}(\widehat{\operatorname{vol}}(\overline{L}^{\otimes m}\otimes\overline{B}^{\vee})-\epsilon).$
@@
   1. (1)
-    $\overline{M}$ and $\overline{G}$ are adelic line bundles, where $\overline{M}$ is effective and $\overline{G}$ is arithmetically ample;
+    $\overline{M}$ and $\overline{G}$ are adelic line bundles, where $\overline{M}$ has a non-zero small section and $\overline{G}$ is arithmetically ample ;
   2. (2)
     $p^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon$.
-Proof.
+Démonstration.
 By definition, there exists an $m\in\mathbb{N}$, such that $\overline{L}^{\otimes m}$ is a usual arithmetically big line bundle.
@@
 For this target, we prove an auxiliary lemma first, which allows us to approximate an arithmetically big $\mathbb{R}$-line bundle by an arithmetically big $\mathbb{Q}$-line bundle along an effective direction.
+Let $\overline{L}_{1},\ldots,\overline{L}_{n}$ be a family of adelic line bundle, and $a_{1},\ldots,a_{n}\in\mathbb{R}$.
+We denote by $\sum\limits_{i=1}^{n}a_{i}L_{i}$ and $\sum\limits_{i=1}^{n}a_{i}\overline{L}_{i}$ the tensor products of line bundles $\bigotimes\limits_{i=1}^{n}L_{i}^{\otimes a_{i}}$ and $\bigotimes\limits_{i=1}^{n}\overline{L}_{i}^{\otimes a_{i}}$ to simplify the notations in the proof below, which are a usual $\mathbb{R}$-line bundle and an adelic $\mathbb{R}$-line bundle respectively.
 […]
@@
 Let $\overline{L}$ be an arithmetically big $\mathbb{R}$-line bundle on $X$.
-Then for all $\epsilon>0$, there exists an arithmetically big $\mathbb{Q}$-line bundle $\overline{L}^{\prime}$, such that $\overline{L}-\overline{L}^{\prime}$ is effective and
+Then for all $\epsilon>0$, there exists an arithmetically big $\mathbb{Q}$-line bundle $\overline{L}^{\prime}$, such that $\overline{L}-\overline{L}^{\prime}$ has a non-zero small section and
 $0<\widehat{\operatorname{vol}}(\overline{L})-\widehat{\operatorname{vol}}(\overline{L}^{\prime})<\epsilon.$
-Proof.
+Démonstration.
 The adelic $\mathbb{R}$-line bundle $\overline{L}$ is arithmetically big, so $L$ is a big $\mathbb{R}$-line bundle.
@@
 By definition, $L_{1},\ldots,L_{r}$ are all effective in the geometric sense.
-We equip metrics on $L_{1},\ldots,L_{r}$ over $S$, such that $\overline{L}_{1},\ldots,\overline{L}_{r}$ are arithmetically effective.
-Then for any $b_{1},\ldots,b_{r}\in\mathbb{R}^{+}$, we have $b_{1}\overline{L}_{1}+\cdots+b_{r}\overline{L}_{r}$ is effective.
-By the continuity of $\widehat{\operatorname{vol}}(\raisebox{1.72218pt}{.})$ from [9, Corollary 6.4.24], for all $\epsilon>0$, there exist $\delta_{1},\ldots,\delta_{r}>0$, such that for all $t_{i}\in(0,\delta_{i})$, $i=1,\ldots,r$, we always have
+We equip metrics on $L_{1},\ldots,L_{r}$ over $S$, such that $\overline{L}_{1},\ldots,\overline{L}_{r}$ have non-zero small sections.
+Then for any $b_{1},\ldots,b_{r}\in\mathbb{R}^{+}$, we have $b_{1}\overline{L}_{1}+\cdots+b_{r}\overline{L}_{r}$ has a non-zero small section.
+[…]
 $0<\widehat{\operatorname{vol}}(\overline{L})-\widehat{\operatorname{vol}}\left(\overline{L}-\sum_{i=1}^{r}t_{i}\overline{L}_{i}\right)<\epsilon.$
@@
 By choosing the constant $\epsilon$ sufficiently small, we pick a family $t_{1},\ldots,t_{r}\in\mathbb{R}^{+}$ such that $\overline{L}-\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ is an arithmetically big $\mathbb{Q}$-line bundle.
-By the above construction, $\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ is effective.
+By the above construction, $\sum\limits_{i=1}^{r}t_{i}\overline{L}_{i}$ has a non-zero small section.
 Let
@@
   1. (1)
-    $\overline{M}$ is an arithmetically effective $\mathbb{R}$-line bundle, and $\overline{G}$ is an arithmetically ample line bundle;
+    $\overline{M}$ is an $\mathbb{R}$-line bundle having a non-zero small section, and $\overline{G}$ is an arithmetically ample line bundle ;
   2. (2)
     $p^{-(d+1)}\widehat{\operatorname{vol}}(\overline{G})\geqslant\widehat{\operatorname{vol}}(\overline{L})-\epsilon$.
-Proof.
-By Lemma 5.2.2, for all $\epsilon>0$, there exists an arithmetically big $\mathbb{Q}$-line bundle $\overline{L}^{\prime}$, such that $\overline{L}-\overline{L}^{\prime}$ is effective and
+Démonstration.
+By Lemma 5.2.2, for all $\epsilon>0$, there exists an arithmetically big $\mathbb{Q}$-line bundle $\overline{L}^{\prime}$, such that $\overline{L}-\overline{L}^{\prime}$ has a non-zero small section and
 $0<\widehat{\operatorname{vol}}(\overline{L})-\widehat{\operatorname{vol}}(\overline{L}^{\prime})<\epsilon.$
 Then we apply Corollary 5.2.1 to $\overline{L}^{\prime}$, and we obtain the assertion. ∎
+Declaration
+Conflict of interest statement
+The author declares that there is no potential conflict of interest.
+Ethical statement
+The author guarantees that the manuscript has neither been published elsewhere nor submitted for publication in another journal, and will not be submitted to other journals while being reviewed.
+The author certify that they have written entirely the manuscript and that its main results are original, and cite all necessary sources.
+The author will inform immediately to the editors if any error is found in the manuscript after the submission.
+The data availability statement
+The data that support the findings of this study are openly available.
+Références
+  • [1] J.-B. Bost & H. Chen – « Concerning the semistability of tensor products in Arakelov geometry », Journal de Mathématiques Pures et Appliquées. Neuvième Série 99 (2013), no. 4, p. 436–488.
+  • [2] S. Boucksom & H. Chen – « Okounkov bodies of filtered linear series », Compositio Mathematica 147 (2011), no. 4, p. 1205–1229.
+  • [3] N.
+BourbakiÉléments de mathématique.
+Fasc.
+XIII.
+Livre VI : Intégration.
+Chapitres 1, 2, 3 et 4 : Inégalités de convexité, Espaces de Riesz, Mesures sur les espaces localement compacts, Prolongement d’une mesure, Espaces $L\mathrm{sp}{p}$, Deuxième édition revue et augmentée.
+Actualités Scientifiques et Industrielles, No.
+1175, Hermann, Paris, 1965.
+  • [4] H. Chen – « Arithmetic Fujita approximation », Annales Scientifiques de l’École Normale Supérieure. Quatrième Série 43 (2010), no. 4, p. 555–578.
+  • [5] — , « Convergence des polygones de Harder-Narasimhan », Mémoires de la Société Mathématique de France 120 (2010), p. 1–120.
+  • [6] — , « Differentiability of the arithmetic volume function », Journal of the London Mathematical Society. Second Series 84 (2011), no. 2, p. 365–384.
+  • [7] — , « Newton-Okounkov bodies : an approach of function field arithmetic », Journal de Théorie des Nombres de Bordeaux 30 (2018), no. 3, p. 829–845.
+  • [8] H. Chen & H. Ikoma – « On subfiniteness of graded linear series », European Journal of Mathematics 6 (2020), no. 2, p. 367–399.
+  • [9] H. Chen & A. MoriwakiArakelov geometry over adelic curves, Lecture Notes in Mathematics, vol. 2258, Springer, Singapore, [2020] ©2020.
+  • [10] — , « Arithmetic intersection theory over adelic curves », (2021), to appear in Mémoires de la Société Mathématique de France. Nouvelle Série.
+  • [11] — , Positivity in Arakelov geometry over adelic curves—Hilbert-Samuel formula and equidistribution theorem, Progress in Mathematics, vol. 355, Birkhäuser/Springer, Cham, [2024] ©2024.
+  • [12] T. Fujita – « Approximating Zariski decomposition of big line bundles », Kodai Mathematical Journal 17 (1994), no. 1, p. 1–3.
+  • [13] G. Harder & M. S. Narasimhan – « On the cohomology groups of moduli spaces of vector bundles on curves », Mathematische Annalen 212 (1974/1975), p. 215–248.
+  • [14] H. Ikoma – « On the concavity of the arithmetic volumes », International Mathematics Research Notices. IMRN (2015), no. 16, p. 7063–7109.
+  • [15] — , « Adelic Cartier divisors with base conditions and the continuity of volumes », Kyoto Journal of Mathematics 61 (2021), no. 4, p. 905–947.
+  • [16] R.
+LazarsfeldPositivity in algebraic geometry.
+I, Ergebnisse der Mathematik und ihrer Grenzgebiete.
+3.
+Folge.
+A Series of Modern Surveys in Mathematics, vol.
+48, Springer-Verlag, Berlin, 2004, Classical setting : line bundles and linear series.
+  • [17] — , Positivity in algebraic geometry.
+II, Ergebnisse der Mathematik und ihrer Grenzgebiete.
+3.
+Folge.
+A Series of Modern Surveys in Mathematics, vol.
+49, Springer-Verlag, Berlin, 2004, Positivity for vector bundles, and multiplier ideals.
+  • [18] R. Lazarsfeld & M. Mustaţă – « Convex bodies associated to linear series », Annales Scientifiques de l’École Normale Supérieure. Quatrième Série 42 (2009), no. 5, p. 783–835.
+  • [19] A. Moriwaki – « Arithmetic height functions over finitely generated fields », Inventiones Mathematicae 140 (2000), no. 1, p. 101–142.
+  • [20] — , « Continuity of volumes on arithmetic varieties », Journal of algebraic geometry 18 (2009), no. 3, p. 407–457.
+  • [21] — , « Continuous extension of arithmetic volumes », International Mathematics Research Notices. IMRN (2009), no. 19, p. 3598–3638.
+  • [22] — , « Adelic divisors on arithmetic varieties », Memoirs of the American Mathematical Society 242 (2016), no. 1144, p. v+122.
+  • [23] A. Sédillot – « Differentiability of the $\chi$-volume function over an adelic curve », (2023), arXiv:2303.03377.
+  • [24] S. Takagi – « Fujita’s approximation theorem in positive characteristics », Journal of Mathematics of Kyoto University 47 (2007), no. 1, p. 179–202.
+  • [25] X. Yuan – « Big line bundles over arithmetic varieties », Inventiones Mathematicae 173 (2008), no. 3, p. 603–649.
+  • [26] — , « On volumes of arithmetic line bundles », Compositio Mathematica 145 (2009), no. 6, p. 1447–1464.
+  • [27] — , « Volumes of arithmetic Okounkov bodies », Mathematische Zeitschrift 280 (2015), no. 3-4, p. 1075–1084.
+  • [28] O. Zariski – « The theorem of Riemann-Roch for high multiples of an effective divisor on an algebraic surface », Annals of Mathematics, Second Series 76 (1962), p. 560–615.
+  • [29] S. Zhang – « Positive line bundles on arithmetic varieties », Journal of the American Mathematical Society 8 (1995), no. 1, p. 187–221.